### v2

# Automorphisms of free metabelian Lie algebras

Abstract.
We show that every automorphism of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ over an arbitrary field $K$ is almost tame, that is, it is a product of so-called Chein automorphisms (or one-row transformations). Moreover, we show that the group of all automorphisms $\mathrm{Aut}(M_{n})$ of $M_{n}$ of rank $n\geq 4$ over a field $K$ of characteristic $\neq 3$ is generated by all linear automorphisms, as well as one quadratic and one cubic automorphism. The same result holds for fields of any characteristic if $n\geq 5$.
We also show that all Chein automorphisms of lower degree $\geq 4$ and all exponential automorphisms of lower degree $\geq 5$ are tame, which contradicts the results of [6, 42].
Ualbai Umirbaev [^1] Department of Mathematics, Wayne State University, Detroit, MI 48202, USA; Department of Mathematics, Al-Farabi Kazakh National University, Almaty, 050040, Kazakhstan; and Institute of Mathematics and Mathematical Modeling, Almaty, 050010, Kazakhstan, e-mail: umirbaev@wayne.edu
Mathematics Subject Classification (2020): 17B40, 17B01, 17B30.
Key words: Automorphism, Lie algebra, free metabelian Lie algebra.
1. Introduction
The well known Jung–van der Kulk Theorem [19, 24] states that every automorphism of the polynomial algebra $K[x,y]$ in two variables $x,y$ over an arbitrary field $K$ is tame. An analogue of this result for free associative algebras in two variables was proven by Makar-Limanov and Czerniakiewicz [15, 28]. The same result is true for free Poisson algebras in two variables over a field of characteristic zero [30] (see also [29, 70, 31]). Moreover, the automorphism groups of polynomial algebras, free associative algebras, and free Poisson algebras in two variables are isomorphic. The tameness of automorphisms of free right-symmetric algebras in two variables over any field was proven in [22].
For any free algebra $A$ in the free variables $x_{1},\ldots,x_{n}$ denote by $\phi=(f_{1},\ldots,f_{n})$ the endomorphism of $A$ determined by $\phi(x_{i})=f_{i}$ for all $1\leq i\leq n$.
The automorphism groups of commutative and associative algebras generated by three elements are much more complicated. The well-known Nagata automorphism [35]
$(x+2y(zx-y^{2})+z(zx-y^{2})^{2},y+z(zx-y^{2}),z)$
of the polynomial algebra $K[x,y,z]$ over a field $K$ of characteristic 0 is proven to be wild [53, 66]. The well-known Anick automorphism (see [14, p. 398])
$(x+z(xz-zy),\,y+(xz-zy)z,\,z)$
of the free associative algebra $K\langle x,y,z\rangle$ over a field $K$ of characteristic 0 is also proven to be wild [67, 68]. The Nagata automorphism gives an example of a wild automorphism of free Poisson algebras in three variables. In 2006 I.P. Shestakov (unpublished) constructed an analogue of the Anick automorphism
$(x+\{xz-\{y,z\},z\},y+z(xz-\{y,z\}),z)$
for free Poisson algebras in three variables.
It is well known [60] that the Nagata and Anick automorphisms are stably tame, that is, they become tame after adding one more variable.
A classical theorem of combinatorial group theory asserts that every subgroup of a free group is free. Nielsen established this remarkable result in 1921 for finitely generated subgroups [40]; in 1927, it was extended to all subgroups by Schreier [57]. A variety of universal algebras is called Nielsen-Schreier, if any subalgebra of a free algebra of this variety is free. In 1953 Shirshiv and in 1956 Witt proved that [54, 72] the varieties of Lie algebras and Lie $p$-algebras over a field are Nielsen-Schreier. The varieties of all non-associative algebras [25], commutative and anti-commutative algebras [55], Lie superalgebras [32, 59], and Lie $p$-superalgebras [33] over a field are also Nielsen-Schreier. It was recently shown [16] that the varieties of pre-Lie (also known as right-symmetric) algebras and Lie-admissible algebras over a field of characteristic zero are Nielsen-Schreier. Some other examples of Nielsen-Schreier varieties can be found in [11, 34, 52, 63, 65].
In 1924, Nielsen also proved [41] that the automorphism group of a finitely generated free group is generated by Nielsen transformations, and represented this group as a finitely defined group. In 1964 P. Cohn proved [12] that all automorphisms of finitely generated free Lie algebras over a field are tame. Later this result was extended to free algebras of Nielsen-Schreier varieties [26]. Defining relations between the elementary automorphisms of free algebras of Nielsen-Schreier varieties are given in [69].
One of the classical and well-studied varieties of groups is the variety of metabelian groups. In 1965 Bachmuth [3] proved that all automorphisms of the free metabelian group of rank $2$ are tame. The first example of a non-tame automorphism of the free metabelian group of rank $3$ was constructed by Chein [10]. Bachmuth-Mochizuki proved [4] that the group of automorphisms of the free metabelian group of rank $3$ is not even finitely generated. Strengthening these results, Roman’kov proved [47] that the free metabelian group of rank $3$ contains non-tame primitive elements and gave a method of constructing of such elements. The well known Bachmuth-Mochizuki-Roman’kov Theorem [5, 46] states that every automorphism of the free metabelian group of rank $n\geq 4$ is tame. More information about automorphisms of free metabelian groups can be found in [50].
The variety of metabelian Lie algebras is also a classical and well-studied variety of Lie algebras. In this paper we denote by $M_{n}$ a free metabelian Lie algebra of rank $n$ over a field $K$. It is well known that any nontrivial exponential automorphism of $M_{2}$ is wild (see, for example [56, Proposition 4],[2, Theorem 3], [44, Theorem 3]). In 1992, Drensky [17] proved that the exponential automorphism $\mathrm{exp}(\mathrm{ad}[x_{1},x_{2}])$ of $M_{3}$ is wild. In 2008, Roman’kov [49] proved that the group of automorphisms of $M_{3}$ cannot be finitely generated modulo the subgroup generated by all exponential and tame automorphisms. Kabanov and Roman’kov [20] constructed strictly nontame primitive elements of $M_{3}$.
In 2015, Nauryzbaev [39] proved that the group of tame automorphisms of $M_{n}$ is generated by all linear automorphisms and the quadratic automorphism
(1) $(x_{1}+[x_{2},x_{3}],x_{2},\ldots,x_{n})$
if either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$, or $n\geq 5$ and $K$ is a field of arbitrary characteristic.
In 1992 V. Drensky [17], in 1993 Papistas [43], and in 1995 Bahturin and Shpilrain [7] proved that free algebras of rank $n\geq 2$ of any polynilpotent variety of Lie algebras that is not abelian or metabelian have wild automorphisms. This result also makes the case of free metabelian Lie algebras more important.
In 1992 Yu. Bahturin and S. Nabiyev [6] claimed that every nontrivial exponential automorphism of the free metabelian Lie algebra $M_{n}$ of any rank $n\geq 2$ over a field of characteristic zero is wild. More examples of wild automorphisms of $M_{n}$ of rank $n\geq 4$ were given in 2008 by Z. Özcurt and N. Ekici [42]. For a long time, the area specialists believed these results were true. Unfortunately, both articles have recently been shown to contain fatal errors [71].
Inner automorphisms of groups play a crucial role in the proofs of Bachmuth-Mochizuki-Roman’kov Theorem [5, 46]. The group automorphism
$(x_{2}^{-1}x_{1}x_{2},x_{2},\ldots,x_{n})$
is extreemly important in their proofs and there is no analogue of this automorphism in the case of free metabelian Lie algebras. This makes the case of algebras more complicated.
One property of the free metabelian Lie algebras is that they admit non-elementary automorphisms that change only one variable and fix the rest. All automorphisms of $M_{n}$ of this type (including elementary ones) are called Chein automorphisms or one-row transformations [18, 48](see definitions in Section 2). The first example of the Chein automorphism is
(2) $(x_{1}+[[x_{2},x_{3}],x_{1}],x_{2},\ldots,x_{n})$
and no one could answer whether it is tame or not. For this reason, the specialists in the area easily accepted the wrong results [6, 42]. This automorphism is wild if $n=3$ [36], and is the first candidate for wildness for all $n\geq 4$.
Chein automorphisms were accepted as almost elementary automorphisms. An automorphism of $M_{n}$ is called almost tame if it is a product of Chein automorphisms [36, 38].
In this paper we show that every automorphism of $M_{n}$ of rank $n\geq 4$ is almost tame. Moreover, $K$ is a field of characteristic $\neq 3$ then the group of all automorphisms $\mathrm{Aut}(M_{n})$ of $M_{n}$ of rank $n\geq 4$ is generated by all linear automorphisms, as well as a quadratic automorphism (1) and a cubic automorphism (2). If $n\geq 5$ then the same result holds for fields of any characteristic.
One of the most important points of the proof is that we notice that all Chein automorphisms of lower degree $\geq 4$ and all exponential automorphisms of lower degree $\geq 5$ are tame. These results contradict the results of [6, 42].
We follow the proof scheme given in [5], and in several places, we repeat the proofs for completeness. The real difficulties arise in those places where inner automorphisms were used in [5].
Thus the question about the existence of wild automorphisms for a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ is still open as well as in the case polynomial and free associative algebras. But in this case it is reduced to a concrete automorphism.
Problem 1.
Is the Chein automorphism (2) of the free metabelian algebra Lie algebra $M_{n}$ of rank $n\geq 4$ tame?
The paper is organized as follows. In Section 2 we define some important automorphisms and some subgroups of the automorphism group $\mathrm{Aut}(M_{n})$. In Section 3, using a representation of $M_{n}$ via Magnus embedding, we give the definition of the Fox derivatives and the Jacobian matrix of an endomorphism, and determine some of their properties. Using some representations of automorphisms from this section, in the following sections we make calculations in the language of the Jacobian matrices instead of automorphisms. For this, in Section 4 we clearly determine some matrix notations. In Section 5 we prove that a huge part of Chein and exponential automorphisms are tame. We begin the study of almost tame automorphisms in Section 6. Section 7 is devoted to the study of some special class of matrices that includes the Jacobian matrices of an important class of automorphisms. In Section 8 we prove a lemma on eliminating a variable in some special cases. Section 9 is devoted to the study of one exceptional automorphism
(3) $(x_{1}+[[x_{1},x_{2}],x_{1}],x_{2}+[[x_{1},x_{2}],x_{2}],x_{3},\ldots,x_{n})$
of $M_{n}$. In Section 10 we show how to eliminate a variable in the general case. In Section 11 the main result and some its corollaries are formulated.
2. Automorphisms of $M_{n}$
Let $L_{n}$ be a free Lie algebra over a field $K$ freely generated by $z_{1},z_{2},\ldots,z_{n}$. Then $Y_{n}=L_{n}/[L_{n},L_{n}]$ is an abelian Lie algebra with a linear basis $y_{1},y_{2},\ldots,y_{n}$ and $M_{n}=L_{n}/[[L_{n},L_{n}],[L_{n},L_{n}]]$ is a free metabelian Lie algebra with free generators $x_{1},\ldots,x_{n}$, where $y_{i}=z_{i}+[L_{n},L_{n}]$ and $x_{i}=z_{i}+[[L_{n},L_{n}],[L_{n},L_{n}]]$ for all $1\leq i\leq n$.
Thus $M_{n}^{\prime}=[M_{n},M_{n}]$ is abelian and the set of all right normed (nonassociative) words of the form
(4) $[\ldots[[x_{i},x_{i_{1}}],x_{i_{2}}],\ldots,x_{i_{k}}],\ i>i_{1}\leq i_{2}\leq\ldots\leq i_{k},$
is a linear basis of $M_{n}^{\prime}$ (see, for example [8, p. 139]).
Denote by $\phi=(f_{1},\ldots,f_{n})$ the endomorphism of $M_{n}$ such that $\phi(x_{i})=f_{i}$ for all $i$. In particular, $\mathrm{id}=(x_{1},\ldots,x_{n})$ is the identity automorphism. An automorphism of $M_{n}$ of the form
$(x_{1},\ldots,x_{i-1},\alpha x_{i}+f,x_{i+1},\ldots,x_{n}),$
where $0\neq\alpha\in K$ and $f$ does not depend on $x_{i}$, is called elementary. The subgroup $\mathrm{TAut}(M_{n})$ of $\mathrm{Aut}(M_{n})$ generated by all elementary automorphisms is called the tame automorphism group, and the elements of this subgroup are called tame automorphisms of $M_{n}$. Non-tame automorphisms of $M_{n}$ are called wild.
Free metabelian Lie algebras admit automorphisms of the form
$(x_{1},\ldots,x_{i-1},f_{i},x_{i+1},\ldots,x_{n}),\ 1\leq i\leq n,$
which include all elementary automorphisms but not necessarily elementary ones. They are called Chein automorphisms or one-row transformations [18]. These automorphisms move only one variable and fix the others. The subgroup $\mathrm{ATAut}(M_{n})$ of $\mathrm{Aut}(M_{n})$ generated by all Chein automorphisms will be called the almost tame automorphism group, and the elements of this subgroup will be called called almost tame automorphisms of $M_{n}$ [36]. Thus
$\mathrm{TAut}(M_{n})\subseteq\mathrm{ATAut}(M_{n})\subseteq\mathrm{Aut}(M_{n}).$
If $n=3$ then every non-elementary Chein automorphism is wild [36, 49] and the automorphism (3) is not almost tame [36]. The case $n=3$ is well studied and defining relations for elementary automorphisms [37] and for Chein automorphisms [38] are given.
As usual, the group of all linear automorphisms of $M_{n}$ will be identified with the group of all invertible matrices $GL_{n}(K)$. Recall that the symmetric group $S_{n}$ is embedded into $GL_{n}(K)$ and any permutation $\pi\in S_{n}$ represents the automorphism $\pi=(x_{\pi(1)},\ldots,x_{\pi(n)})$. For example, the transposition $(23)$ represents the automorphism $(23)=(x_{1},x_{3},x_{2},x_{4},\ldots,x_{n})$.
Obviously, the group of almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated by all linear automorphisms $GL_{n}(K)$ and all Chein automorphisms of the form
(5) $(x_{1}+f,x_{2},\ldots,x_{n}),\ f\in M_{n}^{2}.$
Chein automorphisms of this form are described in Section 3.
The group of tame automorphisms $\mathrm{TAut}(M_{n})$ is generated by all linear automorphisms $GL_{n}(K)$ and all automorphisms of the form (5) where $f\in M_{n}^{2}$ does not depend on $x_{1}$. Moreover, if $n\geq 4$ and $K$ is a field of characteristic $\neq 3$, then $\mathrm{TAut}(M_{n})$ is generated [39] by all linear automorphisms $GL_{n}(K)$ and the quadratic elementary automorphism (1). If $n\geq 5$ then the same result holds over fields of any characteristic.
Let $0\neq z\in[M_{n},M_{n}]$. Then $\mathrm{ad}\,z:M_{n}\to M_{n}(m\mapsto[z,m])$ is a locally nilpotent derivation and $(\mathrm{ad}\,z)^{2}=0$. The exponential automorphism $\exp(\mathrm{ad}\,z)$ of $M_{n}$ can be written as
$\exp(\mathrm{ad}\,z)=\mathrm{id}+\mathrm{ad}\,z=(y_{1}+[z,y_{1}],\ldots,y_{n}+[z,y_{n}]).$
It is known that the exponential automorphism $\exp(\mathrm{ad}\,[x_{1},x_{2}])$ is wild if $n=3$ [17, 36, 49].
Let
$M_{n}=T_{1}\oplus T_{2}\oplus\ldots\oplus T_{k}\oplus\ldots$
be the standard degree grading of $M_{n}$, i.e., $T_{k}$ is the linear span of (non-associative) monomials in $x_{1},x_{2},\ldots,x_{n}$ of degree $k$. If
$f=f_{s}+f_{s+1}+\ldots+f_{t}\in M_{n},f_{j}\in T_{j},f_{s}\neq 0,f_{t}\neq 0,$
then $\deg(f)=t$ is the degree and $\mathrm{ldeg}(f)=s$ is the lower degree of $f$ in the variables $x_{1},x_{2},\ldots,x_{n}$. Set also $\deg(0)=$-$\infty$ and $\mathrm{ldeg}(0)=\infty$.
Let
$\phi=(x_{1}+f_{1},x_{2}+f_{2},\ldots,x_{n}+f_{n}),$
where $f_{i}\in M_{n}^{2}$ for all $1\leq i\leq n$. Then we define $\deg(\phi)$ as the maximum of all $\deg(f_{i}),1\leq i\leq n$, and $\mathrm{ldeg}(\phi)$ as the minimum of all $\mathrm{ldeg}(f_{i}),1\leq i\leq n$. If $\deg(\phi)=\mathrm{ldeg}(\phi)=k\geq 1$ then we say that $\phi$ is homogeneous of degree $k$. In particular, if $k=2$ then $\phi$ is called quadratic and if $k=3$ then it is called cubic.
The group of automorphisms $\mathrm{IAut}(M_{n})$ of $M_{n}$ that induces the identity automorphism on $M_{n}/[M_{n},M_{n}]$ plays an important role in our study. Obviously, $\mathrm{IAut}(M_{n})$ is the set of all automorphisms of $M_{n}$ of lower degree $\geq 2$. The elements of $\mathrm{IAut}(M_{n})$ are called $\mathrm{IA}$-automorphisms.
Here we describe a set of generators of the group of tame $\mathrm{IA}$-automorphisms and almost tame $\mathrm{IA}$-automorphisms.
Lemma 1.
$(a)$ The group of tame $\mathrm{IA}$-automorphisms $\mathrm{TAut}(M_{n})\cap\mathrm{IAut}(M_{n})$ is generated by all linear conjugates of elementary automorphisms of the form (5) where $f\in[M_{n},M_{n}]$ does not depend on $x_{1}$.
$(b)$ The group almost tame $\mathrm{IA}$-automorphisms $\mathrm{TAut}(M_{n})\cap\mathrm{IAut}(M_{n})$ is generated by all linear conjugates of Chein automorphisms of the form (5).
Proof. Every tame automorphism $\phi$ can be represented as
$\phi=\alpha_{1}\phi_{1}\alpha_{2}\phi_{2}\ldots\phi_{k}\alpha_{k+1},$
where $\phi_{i}$ is an elementary automorphism of the form (5) and $\alpha_{i}$ is linear for all $i$. Since $\phi\in\mathrm{IAut}(M_{n})$ for all $i$ it follows that
$\alpha_{1}\alpha_{2}\ldots a_{k+1}=\mathrm{id}.$
since $\phi_{i}\in\mathrm{IAut}(M_{n})$ for all $i$. Then $\alpha_{k+1}=(\alpha_{1}\ldots\alpha_{k})^{-1}$ and, consequently,
$\phi=(\alpha_{1}\phi_{1}\alpha_{1}^{-1})(\alpha_{1}\alpha_{2}\phi_{2}(\alpha_{1}\alpha_{2})^{-1})\ldots(\alpha_{1}\ldots\alpha_{k}\phi_{k}(\alpha_{1}\ldots\alpha_{k})^{-1})$
The statement $(b)$ of the lemma can be proved similarly. $\Box$
3. Jacobian matrices
In this section, we define and state some necessary information about Fox derivatives and Jacobian matrices from [45, 58, 62, 64]. In fact, we define Fox derivatives more directly for free metabelian Lie algebras.
First we give another representation of the free metabelian algebra $M_{n}$ over $K$ with free generators $x_{1},\ldots,x_{n}$. Let $Y_{n}$ be the abelian Lie algebra with a linear basis $y_{1},y_{2},\ldots,y_{n}$. The universal enveloping algebra $U(Y_{n})$ of $Y_{n}$ is the polynomial algebra $U=U(Y_{n})=K[y_{1},\ldots,y_{n}]$ in the variables $y_{1},y_{2},\ldots,y_{n}$. Let
$U=U_{0}\oplus U_{1}\oplus\ldots\oplus U_{k}\oplus\ldots$
be the degree grading of $U$, i.e., $U_{k}$ is the linear span of all homogeneous of degree $k$ monomials in $y_{1},y_{2},\ldots,y_{n}$. If
$a=a_{s}+a_{s+1}+\ldots+a_{t}\in U,a_{j}\in U_{j},a_{s}\neq 0,a_{t}\neq 0,$
then $\deg(a)=t$ is the degree and $\mathrm{ldeg}(a)=s$ is the lower degree of $a$ in the variables $y_{1},y_{2},\ldots,y_{n}$. Set $\deg(0)=$-$\infty$ and $\mathrm{ldeg}(0)=\infty$ as usual.
Let $T_{n}$ be a free right $U$-module with basis $t_{1},\ldots,t_{n}$. Turn the direct sum
$M=Y_{n}\oplus T_{n}$
into a Lie algebra by
$[a+t,b+s]=tb-sa,$
where $a,b\in Y_{n}$ and $t,s\in T_{n}$. Then the subalgebra of $M$ generated by
$x_{i}=y_{i}+t_{i},\ \ 1\leq i\leq n,$
is the free metabelian Lie algebra $M_{n}$ with the free generators $x_{1},\ldots,x_{n}$. In the case of groups the corresponding embedding is called the Magnus embedding [23, 27] and in the case of Lie algebras it follows from Shmel’kin’s Theorem on wreath products [56].
Notice that $M_{n}^{\prime}=[M_{n},M_{n}]\subseteq T_{n}$ and for any $m\in M_{n}^{\prime}$ and $y_{i}$ we get
$my_{i}=[m,x_{i}].$
Every element $m\in M_{n}^{\prime}$ can be written in the form
$m=\sum_{1\leq i<j\leq n}[x_{i},x_{j}]a_{ij},\ \ a_{ij}\in U.$
Let $f\in M_{n}$ be an arbitrary element $M_{n}$. Then $f$ is uniquely represented as
$f=y+t,y\in Y_{n},t\in T_{n}.$
Moreover, $t$ is uniquely represented as
$t=t_{1}d_{1}+\ldots+t_{n}d_{n},\ \ \ \ d_{1},\ldots,d_{n}\in K[y_{1},\ldots,y_{n}]=U.$
Set $d_{i}=\frac{\partial f}{\partial x_{i}}$ for all $1\leq i\leq n$ and
$\partial(f)=(d_{1},\ldots,d_{n})^{t}=\big{(}\frac{\partial f}{\partial x_{1}},\ldots,\frac{\partial f}{\partial x_{n}}\big{)}^{t},$
where $t$ means the transpose. We call $\frac{\partial f}{\partial x_{1}},\ldots,\frac{\partial f}{\partial x_{n}}$ the Fox derivatives of $f$ (see [58, 62]).
Let $e_{1},e_{2},\ldots,e_{n}$ be the columns of the identity matrix $I=I_{n}$ of order $n$.
The statement of the next lemma is well known (see [62, Lemma 5],[64]).
Lemma 2.
Let $a=(a_{1},a_{2},\ldots,a_{n})^{t}\in(U^{n})^{t}$ be an arbitrary column of dimension $n$ over $U$. Then the following conditions are equivalent.
$(1)$ $a=\partial(f)$ for some $f\in[M_{n},M_{n}]$;
$(2)$ $Ya=y_{1}a_{1}+\ldots+y_{n}a_{n}=0$ where $Y=(y_{1},\ldots,y_{n})$;
$(3)$ $a=\sum_{i<j}(e_{i}y_{j}-e_{j}y_{i})a_{ij},\ a_{ij}\in U$.
Proof. If $i<j$ and $a\in U$, then
$\partial([x_{i},x_{j}]a)=e_{i}y_{j}a-e_{j}y_{i}a=(e_{i}y_{j}-e_{j}y_{i})a.$
Consequently, $(1)$ implies $(3)$. Obviously, $(3)$ implies $(2)$ since $Y(e_{i}y_{j}-e_{j}y_{i})=0$ for all $i<j$.
Suppose that $(2)$ holds. Then $a_{n}=y_{1}b_{1}+\ldots+y_{n-1}b_{n-1}$ for some $b_{j}\in U$. Set $g=[x_{n},x_{1}]b_{1}+\ldots+[x_{n},x_{n-1}]b_{n-1}$. Then $\frac{\partial g}{\partial x_{n}}=a_{n}$ and the last component of $a^{\prime}=a-\partial(g)$ is zero. Moreover, if $a^{\prime}=(a_{1}^{\prime},\ldots,a_{n-1}^{\prime},0)$ then $y_{1}a_{1}^{\prime}+\ldots+y_{n-1}a_{n-1}^{\prime}=0$. Leading an induction on the number of nonzero components of $a$, we may assume that there exists $h\in[M_{n},M_{n}]$ such that $a^{\prime}=\partial(h)$. Then $a=\partial(g+h)$. $\Box$
Every endomorphism $\phi$ of $M_{n}$ induces an endomorphism of $Y_{n}=M_{n}/[M_{n},M_{n}]$, which can be uniquely extended to an endomorphism of $U=U(Y_{n})$. This endomorphism will be denoted again by $\phi$ for simplicity of notation. If $f\in U$, then the image $\phi(f)$ will be denoted by $f^{\phi}$.
We define the Jacobian matrix of $\phi=(f_{1},\ldots,f_{n})$ by
$J(\phi)=\left[\frac{\partial f_{j}}{\partial x_{i}}\right]_{1\leq i,j\leq n}=\left[\partial(f_{1})\ \ldots\ \partial(f_{n})\right].$
If $\phi=(f_{1},\ldots,f_{n})$ and $\psi=(g_{1},\ldots,g_{n})$ then
$\phi\circ\psi(x_{i})=g_{i}(f_{1},\ldots,f_{n})$
for all $1\leq i\leq n$. The corresponding Chain Rule gives [58, 62, 64] that
(6) $J(\phi\circ\psi)=J(\phi)J(\psi)^{\phi}.$
The statement of the next lemma is also well known (see [58],[62, Theorem 2]).
Lemma 3.
An endomorphism $\phi$ of $M_{n}$ is an automorphism if and only if the Jacobian matrix $J(\phi)$ is invertible.
Here is a description of Chein automorphisms.
Corollary 1.
An endomorphism of $M_{n}$ of the form (5) is an automorphism if and only if $f$ belongs to the ideal generated by all $x_{s},x_{t}]$, where $s,t\neq 1$.
Proof. Let $\phi=(x_{1}+f,x_{2},\ldots,x_{n})$ where $f\in M_{n}^{2}$. The Jacobian matrix of $\phi$ is $J(\phi)=I+[\partial(f)\ 0\ldots 0]$. This matrix is invertible if and only if $1+\frac{\partial f}{\partial x_{1}}$ is invertible in $U$ or, equivalently, $\frac{\partial f}{\partial x_{1}}=0$. We show that $\frac{\partial f}{\partial x_{1}}=0$ if and only if $f$ belongs to the ideal generated by all $x_{s},x_{t}]$, where $s,t\neq 1$.
The basis (4) of $M_{n}^{2}$ is convenient to work with $\frac{\partial f}{\partial x_{n}}=0$. Any $f\in M_{n}^{2}$ can be written in the form
$f=f^{\prime}+[x_{n},x_{1}]a_{1}+[x_{n},x_{2}]a_{2}+\ldots+[x_{n},x_{n-1}]a_{n-1},$
where $f^{\prime}$ belongs to the ideal generated by all $x_{s},x_{t}]$ with $s,t\neq n$ and $a_{i}\in K[y_{i},\ldots,y_{n}]$ for all $i$. Since $\frac{\partial f^{\prime}}{\partial x_{n}}=0$ it follows that
$\frac{\partial f}{\partial x_{n}}=y_{1}a_{1}+y_{2}a_{2}+\ldots+y_{n-1}a_{n-1}.$
If
$y_{1}a_{1}+y_{2}a_{2}+\ldots+y_{n-1}a_{n-1}=0$
then the substitution $y_{1}=\ldots=y_{n-2}=0$ gives that $a_{n-1}=0$. Repeated substitution $y_{1}=\ldots=y_{n-3}=0$ gives $a_{n-2}=0$. Continuing the same discussions, we get $a_{n-1}=a_{n-2}=\ldots=a_{1}=0$. Threfore $\frac{\partial f}{\partial x_{n}}=0$ if and only if $f=f^{\prime}$. $\Box$
Consider the set of Chein automorphisms of the form
(7) $(x_{1}+[x_{2},x_{3}]a,x_{2},\ldots,x_{n}),\ a\in U.$
These automorphisms are elementary if $a$ does not depend on $y_{1}$.
In this paper we denote by $\mathrm{Mat}_{n}(U)$ the set of all square matrices of order $n$ over $U$ (since the more standard notation $M_{n}(U)$ can be confused with the notation of the free metabelian Lie algebra $M_{n}$). As usual, denote by $GL_{n}(U)$ and $SL_{n}(U)$ the general linear group and the special linear group of matrices of order $n$ over $U$, respectively. The identity matrix of order $n$ is already denoted by $I$.
In Section 2 we mentioned that the group $GL_{n}(K)$ will be identified with the group of all linear automorphisms of $M_{n}$. We meant that each linear automorphism $\alpha$ will be identified with its Jacobian matrix $J(\alpha)$. This means that $\alpha$ is a matrix and instead of $J(\alpha)$ in our calculations we will write $\alpha$.
Lemma 4.
$(a)$ If $\phi,\psi\in\mathrm{IAut}(M_{n})$ then
$J(\phi\psi)=J(\phi)J(\psi);$
$(b)$ If $\phi\in\mathrm{IAut}(M_{n})$ and $\alpha\in GL_{n}(K)$ then
$J(\alpha\phi\alpha^{-1})=\alpha J(\phi)^{\alpha}\alpha^{-1};$
$(c)$ The map
$J:\mathrm{IAut}(M_{n})\to SL_{n}(U)$
is an injective homomorphism of groups and $G=J(\mathrm{IAut}(M_{n}))$ is the set of all matrics from $SL_{n}(U)$ of the form
$I+A,$
where $A\in\mathrm{Mat}(U)$ is a matrix such that $YA=0$.
Proof. Since $\phi$ induces an identity automorphism on $U$, $(a)$ is a direct consequence of (6). Using (6) we also get
$J(\alpha\phi\alpha^{-1})=J(\alpha(\phi\alpha^{-1}))=J(\alpha)J(\phi\alpha^{-1})^{\alpha}=J(\alpha)(J(\phi)J(\alpha^{-1})^{\phi})^{\alpha}=J(\alpha)(J(\phi)J(\alpha^{-1}))^{\alpha}$
$=J(\alpha)J(\phi)^{\alpha}J(\alpha^{-1})^{\alpha}=J(\alpha)J(\phi)^{\alpha}J(\alpha^{-1})=J(\alpha)J(\phi)^{\alpha}J(\alpha)^{-1}=\alpha J(\phi)^{\alpha}\alpha^{-1}.$
In this calculation we again used that $\phi$ induces an identity automorphism on $U$, $J(\alpha^{-1})\in GL_{n}(K)$ is a constant matrix, and that $J(\alpha^{-1})=J(\alpha)^{-1}$ by (6).
If $\phi\in\mathrm{IAut}(M_{n})$, then
$\phi=(x_{1}+f_{1},x_{2}+f_{2},\ldots,x_{n}+f_{n}),$
where $f_{i}\in[M_{n},M_{n}]$ for all $i$, and
$J(\phi)=I+A,\ A=\left[\partial(f_{1})\ \ldots\ \partial(f_{n})\right].$
We get $YA=0$ by Lemma 2. If $J(\phi)=I$ then $\partial(f_{i})=0$ for all $i$. Notice that for any $f\in[M_{n},M_{n}]$ we get $f=(t_{1},\ldots,t_{n})\partial(f)$ by the definition of $\partial(f)$. Consequently, $f_{i}=(t_{1},\ldots,t_{n})\partial(f_{i})=0$. This means $\phi=\mathrm{id}$, i.e., $J$ is injective.
Conversely, let $A\in\mathrm{Mat}_{n}(U)$ such that $YA=0$. By Lemma 2 there exist $g_{1},g_{2},\ldots,g_{n}\in[M_{n},M_{n}]$ such that $A=\left[\partial(g_{1})\ \ldots\ \partial(g_{n})\right]$. Set $\psi=(x_{1}+g_{1},x_{2}+g_{2},\ldots,x_{n}+g_{n})$. Then $J(\psi)=I+A$. If $I+A\in SL_{n}(U)$ then $\psi$ is an automorphism by Lemma 3 and $\psi\in\mathrm{IAut}(M_{n})$. $\Box$
4. Matrix notations
Using Lemma 4, we will represent automorphisms from $\mathrm{IAut}(M_{n})$ by their Jacobian matrices and make calculations in the language of Jacobian matrices. Results formulated in the language of matrices will be also reformulated in the language of automorphisms.
Thus $G$ represents the set of all Jacobian matrices of automorphisms from $\mathrm{IAut}(M_{n})$. Set $T=J(\mathrm{TAut}(M_{n})\cap\mathrm{IAut}(M_{n}))$ and $S=J(\mathrm{ATAut}(M_{n})\cap\mathrm{IAut}(M_{n}))$. We have
$T\subseteq S\subseteq G.$
In order to prove that $\mathrm{ATAut}(M_{n})=\mathrm{Aut}(M_{n})$ it is sufficient to prove that $\mathrm{ATAut}(M_{n})\cap\mathrm{IAut}(M_{n})=\mathrm{IAut}(M_{n})$ since both groups include all linear automorphisms. Thus, it is sufficient to prove that $S=G$.
Let $E_{ij}$ a square matrix with $1$ in the $(i,j)$ position and zeros elsewhere. The order of $E_{ij}$ is usually clear from the context. Recall that
$E_{ij}E_{kl}=\delta_{jk}E_{il},$
where $\delta_{jk}$ is the Kronecker delta. A matrix of the form
$I+aE_{ij},\ i\neq j,\ a\in U,$
is called elementary.
If $i<j$, then for any $k$ denote by $F_{ijk}$ the $n\times n$ matrix with $y_{j}$ in position $(i,k)$, $y_{i}$ in position $(j,k)$, and zeros in other places, that is,
$F_{ijk}=y_{j}E_{ik}-y_{i}E_{jk}.$
In other words, the $k$th column of $F_{ijk}$ is equal to $\partial([x_{i},x_{j}])=e_{i}y_{j}-e_{j}y_{i}$, and the rest are zeros. It is easy to check that
$F_{ijk}F_{stm}=\left\{\begin{array}[]{ccc}y_{t}F_{ijm}&{\rm if}&k=s,\\ -y_{s}F_{ijm}&{\rm if}&k=t,\\ 0&{\rm if}&k\neq s,t\\ \end{array}\right$.,
$E_{ij}F_{kst}=y_{s}E_{ij}E_{kt}-y_{k}E_{ij}E_{st}=\left\{\begin{array}[]{ccc}y_{s}E_{it}&{\rm if}&j=k,\\ -y_{k}E_{it}&{\rm if}&j=s,\\ 0&{\rm if}&j\neq k,s\\ \end{array}\right$.,
and
$F_{kst}E_{ij}=y_{s}E_{kt}E_{ij}-y_{k}E_{st}E_{ij}=\left\{\begin{array}[]{ccc}F_{ksj}&{\rm if}&t=i,\\ 0&{\rm if}&t\neq i\\ \end{array}\right.$.
We refer to these three formulas as the multiplication tables but often use them without reference. Properties of the product of matrices of types
$I+A+B=(I+A)(I+B),$
which is true if $AB=0$, and
$I+(a+b)C=(I+aC)(I+bC),$
which is valid if $C^{2}=0$, to simplify terminology, will be called additivity.
For example, if $f=\sum_{i<j}[x_{i},x_{j}]a_{ij}$, then the Jacobian matrix of the endomorphism (5) is $I+\sum_{i<j}a_{ij}F_{ij1}$, and, by additivity,
$I+\sum_{i<j}a_{ij}F_{ij1}=\prod_{i<j}(I+a_{ij}F_{ij1}).$
Moreover,
$I+(a+b)F_{ij1}=(I+aF_{ij1})(I+bF_{ij1}).$
Consider the set of matrices of the form
(11) $I+aF_{231},\ a\in U.$
These matrices are Jacobian matrices of Chein automorphisms (7).
Here are two remarks:
$(a)$ If $A\in G$ and $\alpha\in GL_{n}(K)$ then $\alpha A^{\alpha}\alpha^{-1}\in G$ by Lemma 4(b)(b). This matrix will be called the linear conjugate of $A$ by $\alpha$. In many places, conjugation by linear automorphisms helps to “linearize” automorphisms [73].
$(b)$ If $J(\phi)=I+A\in S$ and $A$ is a homogeneous matrix, then, usually, it is easy to show that $I+\mu A\in S$ for any $\mu\in K$. This can be done by adding some coefficients in the proofs. For example, this can be done when conjugating matrices (11) using $(x_{1},\mu x_{2},x_{3},\ldots,x_{n})$. For this reason we ignore the difference between the coefficients of this type.
Corollary 2.
$(a)$ The group $T$ is generated by all linear conjugates of matrices of the form (11) where $a$ does not depend on $y_{1}$.
$(b)$ The group $S$ is generated by all linear conjugates of matrices of the form (11).
Proof. By Lemma 1, every automorphism $\mathrm{TAut}(M_{n})\cap\mathrm{IAut}(M_{n})$ is a product of linear conjugates of automorphisms of the form (5). Notice that every automorphism of the form (5) is itself a product of linear conjugates of automorphisms of the form (7) with $a\in K[y_{2},\ldots,y_{n}]$. Then $J(\tau)$ is a matrix of the form (11). If $\alpha\in GL_{n}(K)$ then
$J(\alpha\tau\alpha^{-1})=\alpha J(\tau)^{\alpha}\alpha^{-1}$
by Lemma 4, that is, $J(\alpha\tau\alpha^{-1})$ is the linear conjugate of $J(\tau)$ by $\alpha$. If $\phi,\psi\in\mathrm{IAut}(M_{n})$ then $J(\phi\psi)=J(\phi)J(\psi)$ by the same lemma. This proves $(a)$.
The statement $(b)$ can be checked similarly. $\Box$
5. Tame automorphisms
In this section we assume that either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$ or $n\geq 5$.
Lemma 5.
Every matrix of the form (11) belongs to $T$ if $\mathrm{ldeg}(a)\geq 2$.
Proof. The proof practically repeats the proof of Theorem 1 from [39].
Due to the additivity of (11) with respect to $a$, we can assume that $a=y_{1}^{i_{1}}y_{2}^{i_{2}}\ldots y_{n}^{i_{n}}$. If $i_{1}=0$ then $I+aF_{231}\in T$ by the definition of $T$. Assume that $i_{1}\geq 1$.
Case 1. Suppose that $i_{j}\geq 1$ for some $2\leq j\leq n-1$. Set
$A=I+y_{1}^{i_{1}}\ldots y_{j}^{i_{j}-1}F_{23n},\ \ B=I-y_{j+1}^{i_{j+1}}\ldots y_{n}^{i_{n}}F_{jn1}.$
A direct calculation using the multiplication tables gives that
$[A,B]=ABA^{-1}B^{-1}=I+aF_{231}\in T$
since $A,B\in T$.
Case 2. Let $i_{2}=i_{3}=\ldots=i_{n-1}=0$ and $i_{n}\neq 0$. Then $a=y_{1}^{i_{1}}y_{n}^{i_{n}}$ and we get
$aF_{231}=y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1}F_{2n1}-y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1}F_{3n1}$
by the Jacobi identity. If $i_{n}\geq 2$ then
$I+y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1}F_{2n1},\ I-y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1}F_{3n1}\in T$
by Case 1. Consequently,
$I+aF_{231}=(I+y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1}F_{2n1})(I-y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1}F_{3n1})\in T.$
If $i_{n}=1$, then set $\alpha=I+E_{42}$. Notice that $I+y_{1}^{i_{1}}y_{2}F_{231}\in T$ by Case 1. Hence
$(I+E_{42})(I+y_{1}^{i_{1}}y_{2}F_{231})^{\alpha}(I-E_{42})=I+y_{1}^{i_{1}}y_{2}F_{231}+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341})-y_{1}^{i_{1}}y_{4}F_{341}$
$=(I+y_{1}^{i_{1}}y_{2}F_{231})(I+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341}))(I-y_{1}^{i_{1}}y_{4}F_{341})\in T.$
Since the first and the last factors of this product belong to $T$ it follows that
$C=I+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341})\in T.$
The Jacobi identity implies that
$y_{4}F_{231}=y_{3}F_{241}-y_{2}F_{341}.$
Consequently,
$3aF_{231}=y_{1}^{i_{1}}[(y_{4}F_{231}-y_{2}F_{341})+(y_{4}F_{231}+y_{3}F_{241})]$
and, by additivity,
$I+3aF_{231}=[I+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341})][I+y_{1}^{i_{1}}(y_{4}F_{231}+y_{3}F_{241})].$
The first factor is $C$ and belongs to $T$. The second factor also belongs to $T$ since
$(23)(C^{-1})^{(23)}(23)=(23)((I+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341}))^{-1})^{(23)}(23)$
$=(23)(I-y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341}))^{(23)}(23)=I+(y_{4}F_{231}+y_{3}F_{241}),$
where $(23)=(x_{1},x_{3},x_{2},x_{4},\ldots,x_{n})$. Consequently, $I+3aF_{231}\in T$. If the characteristic of $K$ is not $3$, then its conjugation by $(x_{1},3x_{2},x_{3},\ldots,x_{n})$ gives that $I+aF_{231}\in T$.
If $n\geq 5$, then conjugating (11) by $(4n)$, we can make $i_{4}>0$. Case 1 gives that $I+aF_{231}\in T$.
Case 3. Let $i_{2}=i_{3}=\ldots=i_{n-1}=i_{n}=0$. In this case $a=y_{1}^{i_{1}}$ and $s=i_{1}\geq 2$ since $\mathrm{ldeg}(a)\geq 2$. Case 2 gives that $I+y_{4}y_{1}^{s-1}F_{231}\in T$. Set $\beta=I+E_{14}\in GL_{n}(K)$. Using the multiplication tables we get
$\beta(I+y_{4}y_{1}^{s-1}F_{231})^{\beta}\beta^{-1}=I+(y_{4}+y_{1})y_{1}^{s-1}F_{231}-(y_{4}+y_{1})y_{1}^{s-1}F_{234}$
$=(I+(y_{4}+y_{1})y_{1}^{s-1}F_{231})(I-(y_{4}+y_{1})y_{1}^{s-1}F_{234})$
$=(I+y_{4}y_{1}^{s-1}F_{231})(I+y_{1}^{s}F_{231})(I-y_{4}y_{1}^{s-1}F_{234})(I-y_{1}^{s}F_{234})\in T.$
Since all factors of this product except $I+y_{1}^{s}F_{231}$ belong to $T$, this implies that $I+y_{1}^{s}F_{231}$ also belongs to $T$. $\Box$
Here is a reformulation of this lemma in the language of automorphisms.
Corollary 3.
Every Chein automorphism of the form (5) with $\mathrm{ldeg}(f)\geq 4$ is tame.
Proof. By Corollary 1, any such Chein automorphism is a product of linear conjugates of Chein automorphisms of the form (7) with $\mathrm{ldeg}(a)\geq 2$. Then $J(\tau)$ is a matrix of the form (11) and belongs to $T$ by Lemma 5. $\Box$
This corollary contradicts the results of [42].
Corollary 4.
Assume that either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$ or $n\geq 5$. Then the group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated by all linear automorphisms, as well as a quadratic automorphism (1) and a cubic automorphism (2).
Proof. The group of all tame automorphisms $\mathrm{TAut}(M_{n})$ is generated by all linear automorphisms, as well as a quadratic automorphism (1) [39]. All Chein automorphisms of the form (7) with $\mathrm{ldeg}(a)\geq 2$ belong to $\mathrm{TAut}(M_{n})$. By the additivity of automorphisms of the form (7) with respect to $a$, we may assume that $a$ is homogeneous of degree $\leq 1$. If $a$ does not depend on $y_{1}$ then (7). Again, by additivity, we may assume that $a=y_{1}$. In this case (7) is the cubic automorphism (2). $\Box$
Lemma 6.
Every matrix of the form
$A(a)=I+ay_{1}F_{121}+ay_{2}F_{122}$
belongs to $T$ if $\mathrm{ldeg}(a)\geq 2$.
Proof. By Lemma 5,
$B=I-by_{1}F_{341}-by_{2}F_{342}=(I-by_{1}F_{341})(I-by_{2}F_{342})\in T$
if $\mathrm{ldeg}(b)\geq 1$. Set $C=I+F_{123}\in T$. Then
$[B,C]=BCB^{-1}C^{-1}=I+by_{1}y_{4}F_{121}+by_{2}y_{4}F_{122}=A(y_{4}b)\in T.$
Consequently, the statement of the lemma is true if $y_{4}$ divides $a$. Obviously, the same is true if $y_{i}$ divides $a$ for all $3\leq i\leq n$. By the additivity of $A(a)$ with respect to $a$, we may assume that $a\in K[y_{1},y_{2}]$.
Set $\alpha=I+E_{14}\in T$. If $b\in K[y_{1},y_{2}]$ then, using the multiplication tables, we get
$\alpha A(y_{4}b)^{\alpha}\alpha^{-1}=I+y_{1}(y_{4}+y_{1})bF_{121}+y_{2}(y_{4}+y_{1})bF_{122}-y_{2}(y_{4}+y_{1})bF_{124}$
$=A(y_{4}b)A(y_{1}b)(I-y_{2}(y_{4}+y_{1})bF_{124})\in T.$
Since $A(y_{4}b)\in T$ and $I-y_{2}(y_{4}+y_{1})bF_{124}\in T$ by Lemma 5, this implies that $A(y_{1}b)\in T$. Therefore the statement of the lemma is true if $a\in K[y_{1},y_{2}]$ and $y_{1}$ divides $a$. Obviously, the same is true for $y_{2}$. $\Box$
Corollary 5.
Every automorphism of the form
$(x_{1}+[x_{1},x_{2}]y_{1}a,x_{2}+[x_{1},x_{2}]y_{2}a,x_{3},\ldots,x_{n})$
is tame if $\mathrm{ldeg}(a)\geq 2$.
Proof. By Lemma 6 the Jacobian matrices of these automorphisms belong to $T$. $\Box$
Corollary 6.
Every exponential automorphism of lower degree $\geq 5$ is tame.
Proof. Let
$\exp(\mathrm{ad}(m))=(x_{1}+[m,x_{1}],x_{2}+[m,x_{2}],\ldots,x_{n}+[m,x_{n}])$
be an exponential automorphism with $\mathrm{ldeg}(m)\geq 4$. By the additivity with respect to $m$ we may assume that $m=[x_{1},x_{2}]a$ with $\mathrm{ldeg}(a)\geq 2$. Then the Jacobian matrix of $\exp(\mathrm{ad}(m))$ is
$I+\sum_{i}y_{i}aF_{12i}=(I+y_{1}aF_{121}+y_{2}aF_{122})\prod_{i\geq 3}(I+y_{i}aF_{12i}).$
This matrix belongs to $T$ by Lemmas 5 and 6. $\Box$
This corollary contradicts the main result of [6].
6. Almost tame automorphisms
For the remainder of this article, we always assume that $n\geq 4$ and $K$ is an arbitrary field. By Corollary 2, every matrix of the form (11) and its conjugates belong to $S$. For this reason we do not need the proof of Lemma 5, where the condition that either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$ or $n\geq 5$ was used. The other proofs of Section 5 did not use this condition.
Lemma 7.
Every matrix of the form
$C(a)=I+ay_{1}F_{121}+ay_{2}F_{122}$
with $\mathrm{ldeg}(a)\geq 1$ belongs to $S$.
Proof. For any $b\in U$ we have
$B=I-by_{1}F_{341}-by_{2}F_{342}=(I-by_{1}F_{341})(I-by_{2}F_{342})\in S.$
Repeating the proof of Lemma 6 with this data, we get that $C(a)\in S$ for all $a$ with $\mathrm{ldeg}(a)\geq 1$. $\Box$
Notice that Lemma 7 does not include only the matrix
(12) $I+y_{1}F_{121}+y_{2}F_{122}$
of the same type, which is the Jacobian matrix of (3).
Corollary 7.
Every exponential automorphism of lower degree $\geq 4$ is almost tame.
Proof. Using the decomposition of the Jacobian matrix of an exponential automorphism from the proof of Corollary 6, we conclude that it belongs to $S$ by Lemma 7 if its lower degree $\geq 3$. Consequently, every exponential automorphism of lower degree $\geq 4$ belongs to $\mathrm{ATAut}(M_{n})$. $\Box$
7. Special matrices
We use matrices instead of matrix notation in many places because it makes the calculations easier to understand. To some extent this is due to the application of Suslin’s theorem [61].
The main aim of this section is to prove that every matrix of the form
(13) $I+\begin{bmatrix}y_{n}C&\begin{matrix}0\\ .\\ .\\ .\\ 0\end{matrix}\\ \begin{matrix}*&*&\ldots&*\end{matrix}&0\end{bmatrix}\in G$
belongs to $S$. We will provide several lemmas to the proof of this result. A matrix of the form (13) will be called special.
Lemma 8.
Let $B$ be any matrix from $SL_{n-1}(U,y_{n}U)$. Then $B=I_{n-1}+y_{n}C$ for some matrix $C\in\mathrm{Mat}_{n-1}(U)$ and there exists a unique special matrix $A$ such that
$A=I+\left[\begin{array}[]{cc}y_{n}C&0\\ &0\\ \end{array}\right].$
Proof. Obviously, $B=I_{n-1}+y_{n}C$ for some matrix $C\in\mathrm{Mat}_{n-1}(U)$. If $C=[c_{ij}]$ then
$A=I+\begin{bmatrix}y_{n}D&\begin{matrix}0\\ .\\ .\\ .\\ 0\end{matrix}\\ \begin{matrix}-\sum_{i=1}^{n-1}y_{i}d_{i1}&-\sum_{i=1}^{n-1}y_{i}d_{i2}&\ldots&-\sum_{i=1}^{n-1}y_{i}d_{i(n-1)}\end{matrix}&0\end{bmatrix}$
belongs to $G$ by Lemma 2. If $C=0$ then $A=I$ by Lemma 2. This implies the uniqueness of $A$. $\Box$
If $B\in SL_{n}(U)$ is a matrix of the form $B=\begin{bmatrix}C&0\\ 0&I_{n-k}\end{bmatrix}$, where $C\in SL_{k}(U)$ for some $1\leq k\leq n$, then we say that $B$ is $k$-reducible and $B_{k}=C$ will be called its $k$-reduction. Calculations with $k$-reducible matrices from $G$ can be done by using their $k$-reductions.
Set $A_{n-1}=B$ for matrices $A,B$ satisfying the conditions of Lemma 8. We use this notation for special matrices despite the fact that $A$ is not n1n-1-reducible in general. But Lemma 16 allows us to use the matrix $B$ instead of $C$ in some calculations. Moreover, if $B$ is $k$-reducible for some $1\leq k\leq n-1$, then $B_{k}$ will also be called the $k$-reduction of the special matrix $A$ and the special matrix $A$ will be called special $k$-reducible. Let us specify how can we use $k$-reductions of special matrices. Using (6) it is easy to check the following statements:
$(1)$ If $A$ and $B$ are two special $k$-reducible matrices, then $AB$ is a special $k$-reducible and $(AB)_{k}=A_{k}B_{k}$;
$(2)$ If $A$ is special $k$-reducible and $B\in SL_{n}(U)$ (not necessarily $B\in G$) is $k$-reducible, then $A^{B}=BAB^{-1}$ is special $k$-reducible and $(A^{B})_{k}=B_{k}A_{k}B_{k}^{-1}$;
$(3)$ If $A$ is special $k$-reducible and $\alpha\in SL_{n}(K)$ is $k$-reducible, then the linear conjugation $\alpha A^{\alpha}\alpha^{-1}$ of $A$ by $\alpha$ is special $k$-reducible and $(\alpha A^{\alpha}\alpha^{-1})_{k}=\alpha_{k}A_{k}^{\alpha_{k}}\alpha_{k}^{-1}$.
Lemma 9.
Every special matrix of the form
$A(h,g)=I+hgF_{1n1}+hg^{2}F_{2n1}-hF_{1n2}-hgF_{2n2}$
belongs to $S$.
Proof. We have
$A(h,g)_{2}=I_{2}+y_{n}h\begin{bmatrix}g&-1\\ g^{2}&-g\end{bmatrix}$.
$(a)$ Let $\alpha=I-\lambda E_{21},\lambda\in K$. Then
$A(h,g)_{2}^{\alpha}=I_{2}+y_{n}h^{\alpha}\begin{bmatrix}g^{\alpha}&-1\\ (g^{\alpha})^{2}&-g^{\alpha}\end{bmatrix}$
and
$\alpha_{2}A(h,g)_{2}^{\alpha}\alpha_{2}^{-1}=\begin{bmatrix}1&0\\ -\lambda&1\end{bmatrix}\left(I_{2}+y_{n}h^{\alpha}\begin{bmatrix}g^{\alpha}&-1\\ (g^{\alpha})^{2}&-g^{\alpha}\end{bmatrix}\right)\begin{bmatrix}1&0\\ \lambda&1\end{bmatrix}$
$=I_{2}+y_{n}h^{\alpha}\begin{bmatrix}(g^{\alpha}-\lambda)&-1\\ (g^{\alpha}-\lambda)^{2}&-(g^{\alpha}-\lambda)\end{bmatrix}=A(h^{\alpha},(g^{\alpha}-\lambda))_{2}.$
Consequently, $\alpha A(h,g)^{\alpha}\alpha^{-1}=A(h^{\alpha},(g^{\alpha}-\lambda))$. This means that $A(h,g)\in S$ if and only if $A(h^{\alpha},(g^{\alpha}-\lambda))\in S$. Using this we may assume that $g(0,\ldots,0)=0$.
$(b)$ Set $B=I-fF_{231}\in S$. Then
$B_{3}=\begin{bmatrix}1&0&0\\ -y_{3}f&1&0\\ y_{2}f&0&1\end{bmatrix}$.
Direct calculations give that
$B_{3}A(h,g)_{3}B_{3}^{-1}=I_{3}+y_{n}h\begin{bmatrix}(g-y_{3}f)&-1&0\\ (g-y_{3}f)^{2}&-(g-y_{3}f)&0\\ y_{2}(g-y_{3}f)f&-y_{2}f&0\end{bmatrix}$
[…]
$=A(h,g-y_{3}f)_{3}C_{3},$
where
$C=I+y_{2}h(g-y_{3}f)fF_{3n1}-y_{2}hfF_{3n2}=(I+y_{2}h(g-y_{3}f)fF_{3n1})(I-y_{2}hfF_{3n2}).$
Notice that $C\in S$. Consequently, $A(h,g)\in S$ if and only if $A(h,g-y_{3}f)\in S$. Of course, the same trick can be done with any variable $y_{i}$ if $3\leq i\leq n-1$. Therefore, replacing $g$ by $g-y_{i}f$ for all $3\leq i\leq n-1$ we may assume that
$g=G(y_{1},y_{2})-y_{n}a.$
$(c)$ Now we show how to eliminate $y_{n}a$. First we replace $g$ to
$g+y_{3}a=G(y_{1},y_{2})+y_{3}a-y_{n}a$
using $(b)$. If $\gamma=I+E_{3n}$, then it is easy to check that
$\gamma A(h,g+y_{3}a)^{\gamma}\gamma^{-1}=A(h^{\gamma},(g+y_{3}a)^{\gamma}).$
We have
$(g+y_{3}a)^{\gamma}=g^{\gamma}+(y_{3}+y_{n})a^{\gamma}$
$=G(y_{1},y_{2})-y_{n}a^{\gamma}+(y_{3}+y_{n})a^{\gamma}=G(y_{1},y_{2})+y_{3}a^{\gamma}.$
Using $(b)$ again, we may assume that
$g=G(y_{1},y_{2}).$
(d) Let $f$ be a polynomial in $y_{1},y_{2}$. Then $g$ can be replaced by $g+y_{i}f$, where $i=1,2$. For this we first replace $g$ by $g+y_{3}f$ using $(b)$, i.e., we can assume that $g=G(y_{1},y_{2})+y_{3}f$. If $\delta=I+E_{i3}$, then
$\delta A(h,g)^{\delta}\delta^{-1}=(I-h^{\delta}g^{\delta}F_{1n1}-h^{\delta}(g^{\delta})^{2}F_{2n1})A(h^{\delta},g^{\delta}).$
Since $I-h^{\delta}g^{\delta}F_{1n1}-h^{\delta}(g^{\delta})^{2}F_{2n1}\in S$ it follows that $A(h,g)\in S$ if and only if $A(h^{\delta},g^{\delta})\in S$.
Notice that
$g^{\gamma}=G(y_{1},y_{2})+y_{3}f^{\gamma}+y_{i}f^{\gamma}.$
Using $(b)$, we can replace $g^{\gamma}$ by
$G(y_{1},y_{2})+y_{i}f^{\gamma}.$
This means that we can replace $g=G(y_{1},y_{2})$ by $G(y_{1},y_{2})+y_{i}f$ for any $f$. Then we can make $g=0$. Notice that $A(h,0)=I-hF_{1n2}\in S$. $\Box$
Corollary 8.
Every automorphism of the form
$(x_{1}+[x_{1},x_{n}]hg+[x_{2},x_{n}]hg^{2},x_{2}+[x_{1},x_{n}]h+[x_{2},x_{n}]hg,x_{3},\ldots,x_{n})$
belongs to $W$.
Lemma 10.
Every matrix of the form
$A(h,f,g)=I+hfgF_{1n1}+hg^{2}F_{2n1}-hf^{2}F_{1n2}-hfgF_{2n2}$
belongs to $S$.
Proof. The proof of the lemma repeats the proof given on pages 90–91 of [5]. Notice that $A(h,f,g)$ is an induced $2$-reducible matrix and
$A(h,f,g)_{2}=I_{2}+y_{n}h\begin{bmatrix}fg&-f^{2}\\ g^{2}&-fg\end{bmatrix}$
We have
$A(h,f,g)_{3}=\begin{bmatrix}1&0&0\\ 0&1&0\\ g&-f&1\end{bmatrix}\begin{bmatrix}1&0&-y_{n}hf\\ 0&1&-y_{n}hg\\ 0&0&1\end{bmatrix}\begin{bmatrix}1&0&0\\ 0&1&0\\ -g&f&1\end{bmatrix}\begin{bmatrix}1&0&y_{n}hf\\ 0&1&y_{n}hg\\ 0&0&1\end{bmatrix}$.
In matrix notation, this can be written as
$A(h,f,g)=[(I+gE_{31}-fE_{32}),(I-hfF_{1n3}-hgF_{2n3})]$
$=(I-hfF_{1n3}-hgF_{2n3})^{(I+gE_{31}-fE_{32})}(I+hfF_{1n3}+hgF_{2n3})$
The last factor belongs to $S$. Therefore it is sufficient to prove that
$(I-hfF_{1n3}-hgF_{2n3})^{(I+gE_{31}-fE_{32})}=(I-hfF_{1n3})^{(I+gE_{31}-fE_{32})}(I-hgF_{2n3})^{(I+gE_{31}-fE_{32})}$
belongs to $S$. Moreover, it is sufficient to prove that the first factor of the last product belongs to $S$ since its conjugation by the linear automorphism $(12)$ gives a matrix that has the form of the second factor. We have
$\begin{bmatrix}1&0&0\\ 0&1&0\\ g&-f&1\end{bmatrix}\begin{bmatrix}1&0&-y_{n}hf\\ 0&1&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}1&0&0\\ 0&1&0\\ -g&f&1\end{bmatrix}=I_{3}+y_{n}h\begin{bmatrix}fg&-f^{2}&-f\\ 0&0&0\\ fg^{2}&-f^{2}g&-fg\end{bmatrix}$
[…]
which means that
$(I-hfF_{1n3})^{(I+gE_{31}-fE_{32})}=(I-hf^{2}F_{1n2}-hf^{2}gF_{3n2})$
$\times(I+(hf)gF_{1n1}+(hf)g^{2}F_{3n1}-(hf)F_{1n3}-(hf)gF_{3n3}).$
The first factor of this product belongs to $S$ and the second factor belongs to $S$ by Lemma 9. $\Box$
Corollary 9.
Every automorphism of the form
$(x_{1}+[x_{1},x_{n}]hfg+[x_{2},x_{n}]hg^{2},x_{2}+[x_{1},x_{n}]hf^{2}+[x_{2},x_{n}]hfg,x_{3},\ldots,x_{n})$
belongs to $W$.
First, let us recall some results of Cohn [13] and Suslin [61] on generators of special linear groups over polynomial algebras.
In 1966 P. Cohn [13] proved that the matrix
$\begin{bmatrix}1+x_{1}x_{2}&x_{2}^{2}\\ -x_{1}^{2}&1-x_{1}x_{2}\\ \end{bmatrix}$
is not a product of elementary matrices over the polynomial algebra $K[x_{1},x_{2}]$.
Denote by $E_{k}(U)$ the subgroup of $SL_{k}(U)$ generated by all elementary matrices, that is, by all matrices of the form
$I_{k}+aE_{ij},\ a\in U.$
The well-known Suslin’s Theorem [61] states that if $k\geq 3$ then $SL_{k}(U)=E_{k}(U)$, that is, every matrix from $SL_{k}(U)$ is a product of elementary matrices. If $I$ is an ideal of $U$ then the natural map $U\to U/I$ induces the natural maps $GL_{k}(U)\to GL_{k}(U/I)$ and $GL_{k}(U)\to GL_{k}(U/I)$. Denote by $GL_{k}(U,I)$ the kernel of $GL_{k}(U)\to GL_{k}(U/I)$ and by $SL_{k}(U,I)$ the kernel of $SL_{k}(U)\to SL_{k}(U/I)$. Denote by $E_{k}(U,I)$ the normal closure in $E_{k}(U)$ of the set of all elementary matrices of the form $I+aE_{ij}$, where $i\neq j$ and $a\in I$.
The next corollary of Suslin’s results is formulated in [5].
Proposition 1.
Let $k\geq 3$ and $I=y_{n}U$. Then
(1) $SL_{k}(U,I)=E_{k}(U,I)$;
(2) $E_{k}(U,I)$ is generated by the set of all matrices of the form
$I+h(fe_{i}+ge_{j})(ge_{i}-fe_{j})^{t}$
where $i<j$, $h\in I$, and $f,g\in U$.
Proof. We repeat a short proof given in [5]. By [61], every matrix $\Phi\in SL_{k}(U,I)$ is a product of elementary matrices $\Phi=E_{1}\ldots E_{s}$. Any $a\in U=K[y_{1},\ldots,y_{n}]$ can be written in the form
$a=y_{n}b+c,\ \ b\in U,c\in K[y_{1},\ldots,y_{n-1}],$
and, consequently,
$I+aE_{ij}=(I+y_{n}bE_{ij})(I+cE_{ij}).$
Thus, $E_{i}=H_{i}G_{i}$ where $H_{i}\in E_{k}(U,I)$ and $G_{i}\in E_{k}(K[y_{1},\ldots,y_{n-1}])$. Then
$\Phi=H_{1}(G_{1}H_{2}G_{1}^{-1})(G_{1}G_{2}H_{3}G_{2}^{-1}G_{1}^{-1})\ldots(G_{1}\ldots G_{s-1}H_{s}G_{s-1}^{-1}\ldots G_{1}^{-1})(G_{1}\ldots G_{s}).$
Substituting $y_{n}=0$ we get $G_{1}\ldots G_{s}=1$. Since
$G_{1}\ldots G_{i-1}H_{i}G_{i-1}^{-1}\ldots G_{1}^{-1}\in E_{k}(U,I)$
it follows that $\Phi\in E_{k}(U,I)$.
By [61, Corollary 1.4], the group $E_{k}(U,I)$ is generated by the set of all matrices of the form
$I_{k}+hu(u_{j}e_{i}-u_{i}e_{j})^{t}$
where $h\in I$, $u=(u_{1},\ldots,u_{k})$ is a unimodular column over $U$, and $1\leq i<j\leq k$. Recall that a column $u$ is called unimodular if the ideal generated by its components contains the identity element. We have
$I_{k}+hu(u_{j}e_{i}-u_{i}e_{j})^{t}=(I_{k}+h(u_{i}e_{i}+u_{j}e_{j})(u_{j}e_{i}-u_{i}e_{j})^{t})$
$\prod_{l\neq i,j}(I_{k}+hu_{l}e_{l}(u_{j}e_{i})^{t})\prod_{l\neq i,j}(I_{k}-h(u_{l}e_{l})^{t}u_{i}e_{j}).$
This proves $(b)$. $\Box$
Notice that if $i=1,j=2$, then the matrix from Proposition 1(2)(2) is $2$-reducible and its $2$-reduction is
$I_{2}+h\left[\begin{array}[]{cc}fg&-f^{2}\\ g^{2}&-fg\\ \end{array}\right],$
which is a generalization of the Cohn matrix mentioned above.
Proposition 2.
Every special matrix belongs to $S$.
Proof. Since $n\geq 4$ it follows by Proposition 1 that $SL_{n-1}(U,y_{n}U)$ is generated by the set of all matrices
$B=I_{n-1}+y_{n}h(fe_{i}+ge_{j})(ge_{i}-fe_{j})^{t},\ 1\leq i<j\leq n-1.$
It is sufficient to prove that for any such matrix $B$ there exists a special matrix $A$ from $S$ such that its special (n1)(n-1)-reduction is $B$. Without loss of generality we may assume that $i=1$ and $j=2$, that is,
$B=I_{n-1}+y_{n}h(fe_{1}+ge_{2})(ge_{1}-fe_{2})^{t}.$
Then for
$A=I+hfgF_{1n1}+hg^{2}F_{2n1}-hf^{2}F_{1n2}-hfgF_{2n2}$
we get $A_{n-1}=B$ and $A$ belongs to $S$ by Lemma 10. $\Box$
Corollary 10.
Every automorphism of the form
$(x_{1}+f_{1},x_{2}+f_{2},\ldots,x_{n-1}+f_{n-1},x_{n}),$
where $f_{1},f_{2},\ldots,f_{n-1}$ belong to the ideal generated by $x_{1},x_{n}],\ldots,[x_{n-1},x_{n}]$, belongs to $W$.
8. Inductive lemma
The following lemma is part of the excellent proof given in [5, p. 99–100], and we will repeat this for completeness. Since we have no exponential automorphisms of degree $3$, we first present this statement with a slight restriction. Using this lemma, in the next section we will deal with exponential automorphisms of degree $3$. Then we can prove this lemma in full form. Fortunately, the proof given here works in the general case.
Lemma 11.
Let $A=I+[a_{ij}]\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)\cap G$, $\mathrm{ldeg}(a_{ij})\geq 3$, and $a_{1n},a_{2n},\ldots,a_{(n-1)n}\in\sum_{p=s}^{n-1}y_{p}U+y_{n}^{2}U$. Then there exists $B\in S$ such that $B^{-1}A\in SL_{n}(U,\sum_{p=s}^{n-1}y_{p}U)$.
Proof. For any $a\in U$ and any matrix $R$ over $U$ let $\overline{a}$ and $\overline{R}$ be the result of setting $y_{s}=y_{s+1}=\ldots=y_{n-1}=0$ in $a$ and $R$ respectively.
Let
$M=\begin{bmatrix}I_{n-1}&\begin{matrix}0\\ .\\ .\\ .\\ 0\end{matrix}\\ \begin{matrix}y_{1}&y_{2}&\ldots&y_{n-1}\end{matrix}&y_{n}\end{bmatrix}$.
Then
$\overline{M}=\begin{bmatrix}I_{n-1}&\begin{matrix}0\\ .\\ .\\ .\\ 0\end{matrix}\\ \begin{matrix}y_{1}&\ldots&y_{s-1}&0&\ldots&0\end{matrix}&y_{n}\end{bmatrix}$.
We have
$MAM^{-1}=I_{n}+\begin{bmatrix}[a_{ij}-y_{n}^{-1}y_{j}a_{in}]&\begin{matrix}y_{n}^{-1}a_{1n}\\ .\\ .\\ .\\ y_{n}^{-1}a_{(n-1)n}\end{matrix}\\ \begin{matrix}0&\ldots&0\end{matrix}&0\end{bmatrix}$
[…]
and
[…]
where $\overline{d}_{ij}\in K[y_{1},\ldots,y_{s-1},y_{n}]$. The last equality is possible since $\overline{a}_{ij}\in y_{n}K[y_{1},\ldots,y_{s-1},y_{n}]$ for all $i,j$ and $\overline{a}_{in}\in y_{n}^{2}K[y_{1},\ldots,y_{s-1},y_{n}]$ for all $1\leq i\leq n-1$. Moreover, notice that $\mathrm{ldeg}(\overline{d}_{ij})\geq 2$ for all $1\leq i,j\leq n-1$ and $\mathrm{ldeg}(\overline{d}_{in})\geq 1$ for all $1\leq i\leq n-1$ since $\mathrm{ldeg}(\overline{a}_{ij})\geq 3$ for all $i,j$.
For any $1\leq k\leq n-1$, let
$\delta_{k}=I_{n}+\begin{bmatrix}0&0&\ldots&0\\ .&.&\ldots&.\\ .&.&\ldots&.\\ .&.&\ldots&.\\ 0&0&\ldots&0\\ y_{1}y_{n}\overline{d}_{kn}&y_{2}y_{n}\overline{d}_{kn}&\ldots&y_{n}^{2}\overline{d}_{kn}\\ 0&0&\ldots&0\\ .&.&\ldots&.\\ .&.&\ldots&.\\ .&.&\ldots&.\\ 0&0&\ldots&0\\ -y_{1}y_{k}\overline{d}_{kn}&-y_{2}y_{k}\overline{d}_{kn}&\ldots&-y_{n}y_{k}\overline{d}_{kn}\end{bmatrix}$
$=\prod_{l\neq k,n}(I_{n}+y_{l}\overline{d}_{kn}F_{knl})\cdot(I_{n}+y_{k}\overline{d}_{kn}F_{knk}-y_{n}\overline{d}_{kn}F_{knn}).$
Notice that all factors of the product $\prod_{l\neq k,n}(I_{n}+y_{l}\overline{d}_{kn}F_{knl})$ belong to $S$ and
$I_{n}+y_{k}\overline{d}_{kn}F_{knk}-y_{n}\overline{d}_{kn}F_{knn}\in S$
by Lemma 7 since $\mathrm{ldeg}(\overline{d}_{kn})\geq 1$ (this is the only place where the restriction $\mathrm{ldeg}(a_{ij})\geq 3$ was used). Consequently, $\delta_{k}\in S$.
Notice that
$\overline{M}\overline{\delta}_{1}\ldots\overline{\delta}_{n-1}\overline{M}^{-1}=\begin{bmatrix}I_{n-1}&\begin{matrix}y_{n}\overline{d}_{1n}\\ .\\ .\\ .\\ y_{n}\overline{d}_{(n-1)n}\end{matrix}\\ \begin{matrix}0&\ldots&0\end{matrix}&0\end{bmatrix}$
and
$\overline{M}(\overline{\delta}_{1}\ldots\overline{\delta}_{n-1})^{-1}\Phi\overline{M}^{-1}=\begin{bmatrix}[y_{n}\overline{d}_{ij}]&\begin{matrix}0\\ .\\ .\\ .\end{matrix}\\ \begin{matrix}0&\ldots\end{matrix}&0\end{bmatrix}$.
Let
$\epsilon=I_{n}+\begin{bmatrix}y_{n}\overline{d}_{11}&\ldots&y_{n}\overline{d}_{1(n-1)}&0\\ .&\ldots&.&.\\ .&\ldots&.&.\\ .&\ldots&.&.\\ y_{n}\overline{d}_{(n-1)1}&\ldots&y_{n}\overline{d}_{(n-1)(n-1)}&0\\ -\sum_{j=1}^{n-1}y_{j}\overline{d}_{j1}&\ldots&-\sum_{j=1}^{n-1}y_{j}\overline{d}_{j(n-1)}&0\end{bmatrix}$.
By Proposition 2, $\epsilon\in S\cap SL_{n}(U,\sum_{j=s}^{n}y_{j}U)$, and
$M\epsilon M^{-1}=\overline{M}\overline{\epsilon}\overline{M}^{-1}=\begin{bmatrix}[y_{n}\overline{d}_{ij}]&\begin{matrix}0\\ .\\ .\\ .\end{matrix}\\ \begin{matrix}0&\ldots\end{matrix}&0\end{bmatrix}$.
Consequently,
$\overline{M}\ \overline{\epsilon}^{-1}\ (\overline{\delta}_{1}\ldots\overline{\delta}_{n-1})^{-1}\overline{A}\ \overline{M}^{-1}=\overline{M}\ \overline{B}^{-1}\ \overline{A}\ \overline{M}^{-1}=I_{n},$
where $B=\delta_{1}\ldots\delta_{n-1}\epsilon$. This means that $B^{-1}A\in S\cap SL_{n}(U,\sum_{p=s}^{n-1}y_{p}U)$. $\Box$
9. Exceptional automorphism
In 1981 Shafarevich [51] and in 1983 Anick [1] independently proved that every automorphism of the polynomial algebra $K[x_{1},x_{2},\ldots,x_{n}]$ over a field $K$ of characteristic zero can be approximated by tame automorphisms with respect to two different but similar topologies. This can be considered as a weaker generalization of the Jung-van der Kulk Theorem for polynomial algebras of rank $n\geq 3$. In particular, the Nagata automorphism is wild but can be approximated by tame automorphisms.
An analogue of the Anick-Shafarevich Theorem [1, 51] for free metabelian Lie algebras was proven in 1993 by Bryant and Drensky [9]. They proved that the group of tame automorphisms is dense in the group of all automorphisms of $M_{n}$ if $n\geq 4$. A sharpened version of this result can be found in [21].
Now we prove that the automorphism (3) is almost tame using the ideas of approximation [9] and the results of the preceding sections. Recall that the Jacobian matrix of (3) is the matrix (12).
Lemma 12.
The matrix (12) belongs to $S$.
Proof. We have
$A=I+y_{2}F_{231}-\lambda y_{1}F_{231}\in S.$
Let $\alpha=I+\lambda E_{12}$. Then $B=\alpha A^{\alpha}\alpha^{-1}\in S$. Since $A$ and $\alpha$ are $3$-reducible it follows that $B$ is $3$-reducible and
$B_{3}=\alpha_{3}A_{3}^{\alpha}\alpha_{3}^{-1}=\begin{bmatrix}1&\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}1&0&0\\ (y_{2}-\lambda y_{1})y_{3}&1&0\\ -(y_{2}-\lambda y_{1})y_{2}&0&1\end{bmatrix}^{\alpha}\begin{bmatrix}1&-\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$
$=\begin{bmatrix}1&\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}1&0&0\\ y_{2}y_{3}&1&0\\ -y_{2}(y_{2}+\lambda y_{1})&0&1\end{bmatrix}\begin{bmatrix}1&-\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$
$=\begin{bmatrix}1+\lambda y_{2}y_{3}&-\lambda^{2}y_{2}y_{3}&0\\ y_{2}y_{3}&1-\lambda y_{2}y_{3}&0\\ -y_{2}(y_{2}+\lambda y_{1})&\lambda y_{2}(y_{2}+\lambda y_{1})&1\end{bmatrix}$
Set $C=(I-y_{2}F_{231})(I+\lambda^{2}y_{2}F_{132})B$. Then
$C_{3}=\begin{bmatrix}1&0&0\\ -y_{2}y_{3}&1&0\\ y_{2}^{2}&0&1\end{bmatrix}\begin{bmatrix}1&\lambda^{2}y_{2}y_{3}&0\\ 0&1&0\\ 0&-\lambda^{2}y_{2}y_{1}&1\end{bmatrix}\begin{bmatrix}1+\lambda y_{2}y_{3}&-\lambda^{2}y_{2}y_{3}&0\\ y_{2}y_{3}&1-\lambda y_{2}y_{3}&0\\ -y_{2}(y_{2}+\lambda y_{1})&\lambda y_{2}(y_{2}+\lambda y_{1})&1\end{bmatrix}$
$=\begin{bmatrix}1+\lambda y_{2}y_{3}+\lambda^{2}y_{2}^{2}y_{3}^{2}&-\lambda^{3}y_{2}^{2}y_{3}^{2}&0\\ -\lambda y_{2}^{2}y_{3}^{2}(1+\lambda y_{2}y_{3})&1-\lambda y_{2}y_{3}+\lambda^{3}y_{2}^{3}y_{3}^{3}&0\\ \lambda(-y_{1}y_{2}+y_{2}^{3}y_{3})(1+\lambda y_{2}y_{3})&\lambda y_{2}^{2}+\lambda^{3}y_{1}y_{2}^{2}y_{3}-\lambda^{3}y_{2}^{4}y_{3}^{2}&1\end{bmatrix}$
$=I_{3}+\lambda\begin{bmatrix}y_{2}y_{3}&0&0\\ 0&-y_{2}y_{3}&0\\ -y_{1}y_{2}&y_{2}^{2}&0\end{bmatrix}+y_{2}^{2}y_{3}C^{\prime},$
where
$C^{\prime}=\begin{bmatrix}\lambda^{2}y_{3}&-\lambda^{3}y_{3}&0\\ -\lambda y_{3}(1+\lambda y_{2}y_{3})&\lambda^{3}y_{2}y_{3}^{2}&0\\ \lambda(y_{2}-\lambda y_{1}+\lambda y_{2}^{2}y_{3})&\lambda^{3}y_{1}-\lambda^{3}y_{2}^{2}y_{3}&0\end{bmatrix}$.
Set also $D=(I+y_{2}F_{231})(I-\lambda^{2}y_{2}F_{132})B^{-1}$. Then
$D_{3}=\begin{bmatrix}1-\lambda y_{2}y_{3}+\lambda^{2}y_{2}^{2}y_{3}^{2}&-\lambda^{3}y_{2}^{2}y_{3}^{2}&0\\ -\lambda y_{2}^{2}y_{3}^{2}(1-\lambda y_{2}y_{3})&1+\lambda y_{2}y_{3}-\lambda^{3}y_{2}^{3}y_{3}^{3}&0\\ \lambda(y_{1}y_{2}+y_{2}^{3}y_{3})(1-\lambda y_{2}y_{3})&-\lambda y_{2}^{2}+\lambda^{3}y_{1}y_{2}^{2}y_{3}+\lambda^{3}y_{2}^{4}y_{3}^{2}&1\end{bmatrix}$
and
$(23)_{3}D_{3}^{(23)}(23)_{3}=\begin{bmatrix}1-\lambda y_{2}y_{3}+\lambda^{2}y_{2}^{2}y_{3}^{2}&0&-\lambda^{3}y_{2}^{2}y_{3}^{2}\\ \lambda(y_{1}y_{3}+y_{2}y_{3}^{3})(1-\lambda y_{2}y_{3})&1&-\lambda y_{3}^{2}+\lambda^{3}y_{1}y_{2}y_{3}^{2}+\lambda^{3}y_{2}^{2}y_{3}^{4}\\ -\lambda y_{2}^{2}y_{3}^{2}(1-\lambda y_{2}y_{3})&0&1+\lambda y_{2}y_{3}-\lambda^{3}y_{2}^{3}y_{3}^{3}\end{bmatrix}$
$=I_{3}+\lambda\begin{bmatrix}-y_{2}y_{3}&0&0\\ y_{1}y_{3}&0&-y_{3}^{2}\\ 0&0&y_{2}y_{3}\end{bmatrix}+y_{2}y_{3}^{2}D^{\prime},$
where
$D^{\prime}=\begin{bmatrix}\lambda^{2}y_{2}&0&-\lambda^{3}y_{2}\\ \lambda(y_{3}-\lambda y_{1}-\lambda y_{2}y_{3}^{2})&0&\lambda^{3}y_{1}+\lambda^{3}y_{2}y_{3}^{2}\\ -\lambda y_{2}(1-\lambda y_{2}y_{3})&0&-\lambda^{3}y_{2}^{2}y_{3}\end{bmatrix}$.
Let $E=C(23)D^{(23)}(23)\in S$. A direct calculation gives that
$E_{3}=C_{3}(23)_{3}D_{3}^{(23)}(23)_{3}=(I_{3}+\lambda\begin{bmatrix}y_{2}y_{3}&0&0\\ 0&-y_{2}y_{3}&0\\ -y_{1}y_{2}&y_{2}^{2}&0\end{bmatrix}+y_{2}^{2}y_{3}C^{\prime})$
$\times(I_{3}+\lambda\begin{bmatrix}-y_{2}y_{3}&0&0\\ y_{1}y_{3}&0&-y_{3}^{2}\\ 0&0&y_{2}y_{3}\end{bmatrix}+y_{2}y_{3}^{2}D^{\prime})=I_{3}+\lambda\begin{bmatrix}0&0&0\\ y_{1}y_{3}&-y_{2}y_{3}&-y_{3}^{2}\\ -y_{1}y_{2}&y_{2}^{2}&y_{2}y_{3}\end{bmatrix}$
$+\lambda^{2}y_{2}y_{3}\begin{bmatrix}-y_{2}y_{3}&0&0\\ -y_{1}y_{3}&0&y_{3}^{2}\\ 2y_{1}y_{2}&0&-y_{2}y_{3}\end{bmatrix}+(I_{3}+\lambda\begin{bmatrix}y_{2}y_{3}&0&0\\ 0&-y_{2}y_{3}&0\\ -y_{1}y_{2}&y_{2}^{2}&0\end{bmatrix}+y_{2}^{2}y_{3}C^{\prime})y_{2}y_{3}^{2}D^{\prime}$
$+y_{2}^{2}y_{3}C^{\prime}+\lambda y_{2}^{2}y_{3}C^{\prime}\begin{bmatrix}-y_{2}y_{3}&0&0\\ y_{1}y_{3}&0&-y_{3}^{2}\\ 0&0&y_{2}y_{3}\end{bmatrix}=I_{3}+\lambda\begin{bmatrix}0&0&0\\ y_{1}y_{3}&-y_{2}y_{3}&-y_{3}^{2}\\ -y_{1}y_{2}&y_{2}^{2}&y_{2}y_{3}\end{bmatrix}+E^{\prime},$
where all elements of $E^{\prime}$ are divisible by $y_{3}$ and all elements of the last column of $E^{\prime}$ are divisible by $y_{3}^{2}$. Let
$P=(I+\lambda y_{1}F_{231})(I-\lambda y_{2}F_{232}-\lambda y_{3}F_{233})=I+\lambda y_{1}F_{231}-\lambda y_{2}F_{232}-\lambda y_{3}F_{233}.$
Then
$P_{3}=I_{3}+\lambda\begin{bmatrix}0&0&0\\ y_{1}y_{3}&-y_{2}y_{3}&-y_{3}^{2}\\ -y_{1}y_{2}&y_{2}^{2}&y_{2}y_{3}\end{bmatrix}$.
Then
$P_{3}^{-1}E_{3}=(I_{3}-\lambda\begin{bmatrix}0&0&0\\ y_{1}y_{3}&-y_{2}y_{3}&-y_{3}^{2}\\ -y_{1}y_{2}&y_{2}^{2}&y_{2}y_{3}\end{bmatrix})(I_{3}+\lambda\begin{bmatrix}0&0&0\\ y_{1}y_{3}&-y_{2}y_{3}&-y_{3}^{2}\\ -y_{1}y_{2}&y_{2}^{2}&y_{2}y_{3}\end{bmatrix}+E^{\prime})$
$=I_{3}+(I_{3}-\lambda\begin{bmatrix}0&0&0\\ y_{1}y_{3}&-y_{2}y_{3}&-y_{3}^{2}\\ -y_{1}y_{2}&y_{2}^{2}&y_{2}y_{3}\end{bmatrix})E^{\prime}=I_{3}+P^{\prime}.$
Notice that all elements of $P^{\prime}$ are of lower degree $\geq 4$, divisible by $y_{3}$, and all elements of the last column are divisible by $y_{3}^{2}$. Thus the matrix $(3n)(P^{-1}E)^{(3n)}(3n)\in SL_{n}(U,y_{n}U)$ and satisfies the conditions of Lemma 11 for $s=n$. By Lemma 11, there exists a matrix $Q$ such that $Q^{-1}(3n)(P^{-1}E)^{(3n)}(3n)\in SL_{n}(U,0)$, that is, $Q^{-1}(3n)(P^{-1}E)^{(3n)}(3n)=I$. Consequently, $(3n)(P^{-1}E)^{(3n)}(3n)=Q\in S$ and $P^{-1}E\in S$. Since $E\in S$ it follows that $P\in S$. Notice that the first factor $I+\lambda y_{1}F_{231}$ of $P$ belongs to $S$. Consequently,
$(I-\lambda y_{1}F_{231})P=I-\lambda y_{2}F_{232}-\lambda y_{3}F_{233}\in S.$
Conjugating the inverse of this matrix we can get the matrix (12). $\Box$
Corollary 11.
The automorphism (3) is almost tame.
10. Elimination of a variable
In this section, we prove the main proposition to reduce the number of variables in matrices from $G$ modulo $S$, repeating the proofs from [5, pages 99–100], with some minimal necessary changes. We present two lemmas to avoid yet another use of inner automorphisms in [5].
Lemma 13.
There exists a $3$-reducible matrix $C$ in $T$ such that $C_{3}$ after setting $y_{2}=y_{3}=0$ becomes equal to
$I_{3}+\lambda y_{1}E_{31}=I_{3}+\begin{bmatrix}0&0&0\\ 0&0&0\\ \lambda y_{1}&0&0\end{bmatrix}$,
where $\lambda\in K$.
Proof. Set
$A=I+F_{231},\ \alpha=I+\lambda E_{12},\ B=I-\lambda^{2}F_{132}.$
Then
$A_{3}\alpha_{3}(A_{3}^{-1})^{\alpha}\alpha_{3}^{-1}=\begin{bmatrix}1&0&0\\ y_{3}&1&0\\ -y_{2}&0&1\end{bmatrix}\begin{bmatrix}1&\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}1&0&0\\ -y_{3}&1&0\\ y_{2}&0&1\end{bmatrix}^{\alpha}\begin{bmatrix}1&-\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$
$=\begin{bmatrix}1&0&0\\ y_{3}&1&0\\ -y_{2}&0&1\end{bmatrix}\begin{bmatrix}1&\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}1&0&0\\ -y_{3}&1&0\\ y_{2}+\lambda y_{1}&0&1\end{bmatrix}\begin{bmatrix}1&-\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$
$=\begin{bmatrix}1-\lambda y_{3}&\lambda^{2}y_{3}&0\\ -\lambda y_{3}^{2}&1+\lambda y_{3}+\lambda^{2}y_{3}^{2}&0\\ \lambda y_{1}+\lambda y_{2}y_{3}&-\lambda y_{2}-\lambda^{2}y_{1}-\lambda^{2}y_{2}y_{3}&1\end{bmatrix}$
and
$B_{3}A_{3}\alpha_{3}(A_{3}^{-1})^{\alpha}\alpha_{3}^{-1}=\begin{bmatrix}1&-\lambda^{2}y_{3}&0\\ 0&1&0\\ 0&\lambda^{2}y_{1}&1\end{bmatrix}\begin{bmatrix}1-\lambda y_{3}&\lambda^{2}y_{3}&0\\ -\lambda y_{3}^{2}&1+\lambda y_{3}+\lambda^{2}y_{3}^{2}&0\\ \lambda y_{1}+\lambda y_{2}y_{3}&-\lambda y_{2}-\lambda^{2}y_{1}-\lambda^{2}y_{2}y_{3}&1\end{bmatrix}$
$=\begin{bmatrix}1-\lambda y_{3}+\lambda^{3}y_{3}^{3}&-\lambda^{3}y_{3}^{2}-\lambda^{4}y_{3}^{3}&0\\ -\lambda y_{3}^{2}&1+\lambda y_{3}+\lambda^{2}y_{3}^{2}&0\\ \lambda y_{1}+\lambda y_{2}y_{3}-\lambda^{3}y_{1}y_{3}^{2}&-\lambda y_{2}-\lambda^{2}y_{2}y_{3}+\lambda^{3}y_{1}y_{3}+\lambda^{4}y_{1}y_{3}^{2}&1\end{bmatrix}$.
If $C=BA\alpha(A^{-1})^{\alpha}\alpha^{-1}$ then $C\in T$ and $C_{3}=B_{3}A_{3}\alpha_{3}(A_{3}^{-1})^{\alpha}\alpha_{3}^{-1}$ after setting $y_{2}=y_{3}=0$ gives $I_{3}+\lambda y_{1}E_{31}$. $\Box$
Lemma 14.
Let
$I_{n}+[a_{ij}]\in SL_{n}(U,y_{1}U+y_{2}U)\cap G.$
Then $\mathrm{ldeg}(a_{ij})\geq 2$ for all $1\leq i,j\leq 2$.
Proof. Let
$A=[a_{ij}]=A_{1}+\ldots+A_{k},$
where all elements of $A_{i}$ are homogeneous of degree $i$ for all $1\leq i\leq k$ and let $A_{1}=[b_{ij}]$. Since $a_{ij}\in y_{1}U+y_{2}U$ then $b_{ij}\in Ky_{1}+Ky_{2}$. Moreover, since $YA=0$ by Lemma 2, we get $YA_{1}=0$. By Lemma 2 there exist quadratic elements $f_{1},\ldots,f_{n}\in M_{n}$ such that
$A_{1}=[b_{ij}]=[\partial(f_{1})\ldots\partial(f_{n})].$
Therefore all components of $\partial(f_{i})$ belong to $Ky_{1}+Ky_{2}$.
Let
$f_{i}=\sum_{k=1}^{n-1}\alpha_{k}[x_{n},x_{k}]+f_{i}^{\prime},$
where $f_{i}^{\prime}$ does not depend on $x_{n}$ and $\alpha_{1},\ldots,\alpha_{n-1}\in K$. If $\alpha_{j}\neq 0$ for some $1\leq j\leq n-1$ then
$\frac{\partial f_{i}}{\partial x_{j}}=-\alpha_{j}x_{n}+\frac{\partial f_{i}^{\prime}}{\partial x_{j}}\notin Ky_{1}+Ky_{2}.$
Consequently, $\alpha_{1}=\ldots=\alpha_{n-1}=0$ and $f_{i}$ does not depend on $x_{n}$. The same discussions give that $f_{i}$ does not depend on $x_{3},\ldots,x_{n}$. Hence $f_{i}=\lambda_{i}[x_{1},x_{2}]$ for all $i$. Then
$A_{1}=\lambda_{1}F_{121}+\lambda_{2}F_{122}+\ldots+\lambda_{n}F_{12n}$
and the trace $\mathrm{tr}(A_{1})$ of $A_{1}$ is $\lambda_{1}y_{2}-\lambda y_{1}$.
Recall that the determinat of $I+A$ is $1$. The homogeneous part of degree $1$ of the determinant of $I+A_{1}+\ldots+A_{k}$ is $\mathrm{tr}(A_{1})$. Consequently, $\mathrm{tr}(A_{1})=0$. The equation $\lambda_{1}y_{2}-\lambda y_{1}=0$ gives that $\lambda_{1}=\lambda_{2}=0$. This means that $b_{ij}=0$ and $\mathrm{ldeg}(a_{ij})\geq 2$ for all $1\leq i,j\leq 2$. $\Box$
Lemma 15.
Let
$A=I+[a_{ij}]\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)\cap G.$
Then there exists $R\in S$ such that $R^{-1}A=I+[e_{ij}]\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)\cap G$ and $e_{1n},e_{2n},\ldots,e_{n-1n}\in\sum_{p=s}^{n-1}y_{p}U+y_{n}^{2}U$.
Proof. For any $a\in U$ and any matrix $R$ over $U$ let $\overline{a}$ and $\overline{R}$ be the result of setting $y_{s}=y_{s+1}=\ldots=y_{n-1}=0$ in $a$ and $R$, respectively.
By Lemma 2,
$y_{1}a_{1j}+y_{2}a_{2j}+\ldots y_{n}a_{nj}=0,\ 1\leq j\leq n.$
Consequently,
$y_{1}\overline{a}_{1j}+\ldots y_{s-1}\overline{a}_{(s-1)j}+y_{n}\overline{a}_{nj}=0,\ 1\leq j\leq n.$
Since $A=I+[a_{ij}]\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)$ it follows that $a_{ij}\in\sum_{p=s}^{n}y_{p}U$ and, consequently, $\overline{a}_{ij}\in y_{n}K[y_{1},\ldots,y_{s-1},y_{n}]$, that is, $\overline{a}_{ij}=y_{n}\overline{b}_{ij}$ for some $\overline{b}_{ij}\in K[y_{1},\ldots,y_{s-1},y_{n}]$. Therefore $\overline{A}=I+[\overline{a}_{ij}]=I+[y_{n}\overline{b}_{ij}]$ and the preceding equality gives that
$y_{1}\overline{b}_{1j}+\ldots y_{s-1}\overline{b}_{(s-1)j}+y_{n}\overline{b}_{nj}=0,\ 1\leq j\leq n.$
If $(\overline{c}_{1n},\ldots,\overline{c}_{(n-1)n})$ is the result of setting $y_{n}=0$ in $(\overline{b}_{1n},\ldots,\overline{b}_{(n-1)n})$ then from this we get
(16) $y_{1}\overline{c}_{1n}+\ldots y_{s-1}\overline{c}_{(s-1)n}=0.$
First we show that there exists a matrix $R\in S\cap SL_{n}(U,\sum_{p=s}^{n}y_{p}U)$ such that
$\overline{R}=C=\begin{bmatrix}I_{n-1}&\begin{matrix}y_{n}\overline{c}_{1n}\\ .\\ .\\ .\\ y_{n}\overline{c}_{(n-1)n}\end{matrix}\\ \begin{matrix}0&\ldots&0\end{matrix}&1\end{bmatrix}$.
belongs to $S$. In fact,
$C=\begin{bmatrix}I_{n-1}&\begin{matrix}y_{n}\overline{c}_{1n}\\ .\\ .\\ .\\ y_{n}\overline{c}_{(s-1)n}\\ 0\\ .\\ .\\ .\\ 0\end{matrix}\\ \begin{matrix}0&\ldots&0\end{matrix}&1\end{bmatrix}\begin{bmatrix}I_{n-1}&\begin{matrix}0\\ .\\ .\\ .\\ 0\\ y_{n}\overline{c}_{sn}\\ .\\ .\\ .\\ y_{n}\overline{c}_{(n-1)n}\end{matrix}\\ \begin{matrix}0&\ldots&0\end{matrix}&1\end{bmatrix}$.
The first factor belongs to $G$ by (16) and Lemma 2, and, consequently, belongs to $S$. The second factor is equal to the product
$D=\prod_{k=s}^{n-1}(I+y_{n}\overline{c}_{kn}E_{kn})$
Let
$\overline{c}_{kn}=y_{1}f_{1}+\ldots+y_{s-1}f_{s-1}+\lambda,$
where $f_{i}\in K[y_{1},\ldots,y_{s-1}]$ and $\lambda\in K$. Then
$y_{n}\overline{c}_{kn}=y_{1}y_{n}f_{1}+\ldots+y_{s-1}y_{n}f_{s-1}+y_{n}\lambda.$
Let
$P_{k}=\begin{bmatrix}I_{n-1}&\begin{matrix}-y_{k}y_{n}f_{1}\\ .\\ .\\ .\\ -y_{k}y_{n}f_{s-1}\\ 0\\ y_{1}y_{n}f_{1}+\ldots+y_{s-1}y_{n}f_{s-1}\\ 0\end{matrix}\\ \begin{matrix}0&\ldots&0\end{matrix}&1\end{bmatrix}\in S.$
Let s=n1s=n-1. Then k=n1k=n-1 and $a_{ij}\in y_{n-1}U+y_{n}U$. By Lemma 14, $\mathrm{ldeg}(a_{ij})\geq 2$ for all $n-1\leq i,j\leq n$. By construction the linear part of $y_{n}\overline{c}_{kn}$ is obtained from the linear part of $a_{(n-1)n}$. Consequently, $\lambda=0$ and $\overline{P}_{k}=I+y_{n}\overline{c}_{kn}E_{kn}$.
Let s<n1s<n-1, i.e., the set $\{s,\ldots,n\}$ contains at least $3$ elements. Then, by Lemma 13, there exists a matrix $Q\in T$ such that
$\overline{Q}=I_{n}+\lambda y_{n}E_{kn}.$
Consequently, $QP\in S$ and $\overline{QP}=I+y_{n}\overline{c}_{kn}E_{kn}$.
Therefore in both cases there exists a matrix $\gamma_{i}\in S$ such that $\overline{\gamma_{i}}=I+y_{n}\overline{c}_{kn}E_{kn}$. Then $\overline{\gamma_{s}\ldots\gamma_{n-1}}=D$. Consequently, there exists a matrix $R\in S\cap SL_{n}(U,\sum_{p=s}^{n}y_{p}U)$ such that $\overline{R}=C$. Then $R^{-1}A\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)\cap G$. Let $R^{-1}A=I_{n}+[e_{ij}]$. Recall that, by construction, $\overline{a}_{kn}=y_{n}\overline{c}_{kn}+y_{n}^{2}\overline{d}_{kn}$ for all $1\leq k\leq n-1$, where $\overline{d}_{kn}\in K[y_{1},\ldots,y_{s-1},y_{n}]$. Then
[…]
$=I+\begin{bmatrix}[\overline{a}_{ij}]_{i,j\leq n-1}&\begin{matrix}\overline{a}_{1n}-y_{n}\overline{c}_{1n}\\ .\\ .\\ .\\ \overline{a}_{(n-1)n}-y_{n}\overline{c}_{(n-1)n}\\ \overline{a}_{nn}\end{matrix}\end{bmatrix}=I+\begin{bmatrix}[\overline{a}_{ij}]_{i,j\leq n-1}&\begin{matrix}y_{n}^{2}\overline{d}_{1n}\\ .\\ .\\ .\\ y_{n}^{2}\overline{d}_{(n-1)n}\\ \overline{a}_{nn}\end{matrix}\end{bmatrix}$
Then $\overline{e}_{kn}=y_{n}^{2}\overline{d}_{kn}$. This means that $e_{kn}\in\sum_{p=s}^{n-1}y_{p}U+y_{n}^{2}U$ for all $1\leq k\leq n-1$. The matrix $R^{-1}A$ satisfies the statement of the lemma. $\Box$
For convenience, we formulate the following complete version of Lemma 11.
Lemma 16.
Let $A=I+[a_{ij}]\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)\cap G$ and $a_{1n},a_{2n},\ldots,a_{(n-1)n}\in\sum_{p=s}^{n-1}y_{p}U+y_{n}^{2}U$. Then there exists $B\in S$ such that $B^{-1}A\in SL_{n}(U,\sum_{p=s}^{n-1}y_{p}U)$.
Proof. In the proof of Lemma 11, the restriction $\mathrm{ldeg}(a_{ij})\geq 3$ was used only once to apply that the matrix
$I_{n}+y_{k}\overline{d}_{kn}F_{knk}-y_{n}\overline{d}_{kn}F_{knn}$
belong to $S$ by Lemma 7 since $\mathrm{ldeg}(\overline{d}_{kn})\geq 1$. By Lemma 12, it belongs to $S$ anyway. Repeating the proof of Lemma 11, we get the statement of the lemma. $\Box$
Proposition 3.
Let
$A=I+[a_{ij}]\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)\cap G.$
Then there exists $B\in S$ such that $B^{-1}A\in SL_{n}(U,\sum_{p=s}^{n-1}y_{p}U)$.
Proof. By Lemma 15, there exists a matrix $C\in S$ such that $C^{-1}A=I+[c_{ij}]\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)$ and $c_{1n},c_{2n},\ldots,c_{n-1n}\in\sum_{p=s}^{n-1}y_{p}U+y_{n}^{2}U$. By Lemma 16, there exists a matrix $D\in S$ such that $D^{-1}C^{-1}A\in SL_{n}(U,\sum_{p=s}^{n-1}y_{p}U)$. $\Box$
11. Main results
Theorem 1.
Every automorphism of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ over an arbitrary field $K$ is almost tame.
Proof. First we show that $S=G$. Let $A\in G$. Then
$A\in SL_{n}(U,\sum_{j=1}^{n}y_{j}U).$
By Proposition 3, there exist matrices $B_{1},\ldots,B_{n}\in S$ such that
$B_{s}^{-1}\ldots B_{1}^{-1}A\in SL_{n}(U,\sum_{j=s+1}^{n}y_{j}U)$
for all $s<n$ and
$B_{n}^{-1}\ldots B_{1}^{-1}A\in SL_{n}(U,0)=\{I\}.$
Consequently,
$B_{n}^{-1}\ldots B_{1}^{-1}A=I$
and
$A=B_{1}\ldots B_{n}\in S.$
This proves that $S=G$. Therefore $\mathrm{ATAut}(M_{n})\cap\mathrm{IAut}(M_{n})=\mathrm{IAut}(M_{n})$. It follows that $\mathrm{ATAut}(M_{n})=\mathrm{Aut}(M_{n})$ since both groups contain all linear automorphisms. $\Box$
Theorem 2.
The group of all automorphisms $\mathrm{Aut}(M_{n})$ of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ over a field $K$ of characteristic $\neq 3$ is generated by all linear automorphisms, as well as a quadratic automorphism (1) and a cubic automorphism (2).
Corollary 12.
The group of all automorphisms $\mathrm{Aut}(M_{n})$ of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 5$ over any field $K$ is generated by all linear automorphisms, as well as a quadratic automorphism (1) and a cubic automorphism (2).
Theorem 2 and Corollary 12 are direct corollaries of Corollary 4 and Theorem 1.
Acknowledgments
The author would like to thank the Max Planck Institute für Mathematik for its hospitality and excellent working conditions, where some part of this work has been done. The author thanks Professors V. Dotsenko, M. Jibladze, L. Makar-Limanov, and I. Shestakov for useful discussions.
The project is supported by the grant AP14872073 of the Ministry of Education and Science of the Republic of Kazakhstan.

### v3

@@
-Automorphisms of free metabelian Lie algebras
+Automorphisms of free metabelian Lie algebras, I
 Abstract.
-We show that every automorphism of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ over an arbitrary field $K$ is almost tame, that is, it is a product of so-called Chein automorphisms (or one-row transformations).
-Moreover, we show that the group of all automorphisms $\mathrm{Aut}(M_{n})$ of $M_{n}$ of rank $n\geq 4$ over a field $K$ of characteristic $\neq 3$ is generated by all linear automorphisms, as well as one quadratic and one cubic automorphism.
-The same result holds for fields of any characteristic if $n\geq 5$.
-We also show that all Chein automorphisms of lower degree $\geq 4$ and all exponential automorphisms of lower degree $\geq 5$ are tame, which contradicts the results of [6, 42].
+We show that all Chein automorphisms (or one-row transformations) of lower degree $\geq 4$ of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ over an arbitrary field $K$ of characteristic $\neq 3$ are tame.
+We then show that all exponential automorphisms of $M_{n}$ of lower degree $\geq 5$ are also tame under the same conditions.
+The same results hold for fields of any characteristic when $n\geq 5$.
+These results contradict some long-standing results in the area.
+We also prove that a large class of automorphisms of $M_{n}$ of rank $n\geq 4$ that move only two variables are almost tame, that is, they can be expressed as a product of Chein automorphisms.
 Ualbai Umirbaev [^1] Department of Mathematics, Wayne State University, Detroit, MI 48202, USA; Department of Mathematics, Al-Farabi Kazakh National University, Almaty, 050040, Kazakhstan; and Institute of Mathematics and Mathematical Modeling, Almaty, 050010, Kazakhstan, e-mail: umirbaev@wayne.edu
-Mathematics Subject Classification (2020): 17B40, 17B01, 17B30.
-Key words: Automorphism, Lie algebra, free metabelian Lie algebra.
+Mathematics Subject Classification (2020): 17B40, 17B01, 17B30, 16W20.
+Key words: Automorphism, derivation, free metabelian Lie algebra.
 1. Introduction
 The well known Jung–van der Kulk Theorem [16, 20] states that every automorphism of the polynomial algebra $K[x,y]$ in two variables $x,y$ over an arbitrary field $K$ is tame.
-An analogue of this result for free associative algebras in two variables was proven by Makar-Limanov and Czerniakiewicz [15, 28].
+An analogue of this result for free associative algebras in two variables was proven by Czerniakiewicz and Makar-Limanov [12, 24].
 The same result is true for free Poisson algebras in two variables over a field of characteristic zero [26] (see also [25, 27, 65]).
 Moreover, the automorphism groups of polynomial algebras, free associative algebras, and free Poisson algebras in two variables are isomorphic.
-The tameness of automorphisms of free right-symmetric algebras in two variables over any field was proven in [22].
+The tameness of automorphisms of free right-symmetric algebras in two variables was proven in [18].
 For any free algebra $A$ in the free variables $x_{1},\ldots,x_{n}$ denote by $\phi=(f_{1},\ldots,f_{n})$ the endomorphism of $A$ determined by $\phi(x_{i})=f_{i}$ for all $1\leq i\leq n$.
@@
 $(x+2y(zx-y^{2})+z(zx-y^{2})^{2},y+z(zx-y^{2}),z)$
-of the polynomial algebra $K[x,y,z]$ over a field $K$ of characteristic 0 is proven to be wild [53, 66]. The well-known Anick automorphism (see [14, p. 398])
+of the polynomial algebra $K[x,y,z]$ over a field $K$ of characteristic 0 is proven to be wild [46, 60]. The well-known Anick automorphism (see [11, p. 398])
 $(x+z(xz-zy),\,y+(xz-zy)z,\,z)$
@@
 The Nagata automorphism gives an example of a wild automorphism of free Poisson algebras in three variables.
-In 2006 I.P.
-Shestakov (unpublished) constructed an analogue of the Anick automorphism
-$(x+\{xz-\{y,z\},z\},y+z(xz-\{y,z\}),z)$
-for free Poisson algebras in three variables.
+Recently Shestakov and Zhang [48] constructed an analogue of the Anick automorphism for free Poisson algebras in three variables.
 It is well known [55] that the Nagata and Anick automorphisms are stably tame, that is, they become tame after adding one more variable.
@@
 The variety of metabelian Lie algebras is also a classical and well-studied variety of Lie algebras.
-In this paper we denote by $M_{n}$ a free metabelian Lie algebra of rank $n$ over a field $K$.
-It is well known that any nontrivial exponential automorphism of $M_{2}$ is wild (see, for example [56, Proposition 4],[2, Theorem 3], [44, Theorem 3]).
+Throughout this paper, $M_{n}$ denotes the free metabelian Lie algebra of rank $n$ over a field $K$ generated by $x_{1},\ldots,x_{n}$.
+It is well known that any nontrivial exponential automorphism of $M_{2}$ is wild (see, for example [1, 51]).
 In 1992, Drensky [14] proved that the exponential automorphism $\mathrm{exp}(\mathrm{ad}[x_{1},x_{2}])$ of $M_{3}$ is wild.
-In 2008, Roman’kov [49] proved that the group of automorphisms of $M_{3}$ cannot be finitely generated modulo the subgroup generated by all exponential and tame automorphisms.
-Kabanov and Roman’kov [20] constructed strictly nontame primitive elements of $M_{3}$.
-In 2015, Nauryzbaev [39] proved that the group of tame automorphisms of $M_{n}$ is generated by all linear automorphisms and the quadratic automorphism
+More examples of wild automorphisms of $M_{3}$ are given in [17, 32, 43].
+In 1992 Bahturin and Nabiyev [5] claimed that every nontrivial exponential automorphism of the free metabelian Lie algebra $M_{n}$ of any rank $n\geq 2$ over a field of characteristic zero is wild.
+More examples of wild automorphisms of $M_{n}$ of rank $n\geq 4$ were given in 2008 by Özcurt and Ekici [39].
+For a long time, the area specialists believed these results were true.
+Unfortunately, both articles have fatal errors [66].
+We show that all Chein automorphisms (or one-row transformations) of lower degree $\geq 4$ of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ over an arbitrary field $K$ of characteristic $\neq 3$ are tame.
+We then show that all exponential automorphisms of $M_{n}$ of lower degree $\geq 5$ are also tame under the same conditions.
+The same results hold for fields of any characteristic when $n\geq 5$.
+These results contradict those of [5, 39].
+We also prove that a large important class of automorphisms of $M_{n}$ of rank $n\geq 4$ that move only two variables are almost tame, that is, they can be expressed as a product of Chein automorphisms.
+More precise formulations of the results are given in the following sections.
+The paper is organized as follows.
+In Section 2 we define some important automorphisms of $M_{n}$ and give precise formulations some of our results.
+In Section 3, using a representation of $M_{n}$ via Magnus embedding and Fox derivatives, we describe Chein autoimorphisms of $M_{n}$.
+In Section 4 we prove that a huge part of Chein and exponential automorphisms are tame.
+In Section 5 we prove that a huge class of automorphisms that move only two variables are almost tame.
+2. Automorphisms of $M_{n}$
+Let $M_{n}$ be the free metabelian Lie algebra over a field $K$ with free generators $x_{1},\ldots,x_{n}$. The set of all right normed Lie monomials of the form
+(1) $[\ldots[[x_{i},x_{i_{1}}],x_{i_{2}}],\ldots,x_{i_{k}}],\ i>i_{1}\leq i_{2}\leq\ldots\leq i_{k},$
+is a linear basis of $M_{n}^{\prime}=[M_{n},M_{n}]$ (see, for example [6, p. 139]).
+Let $\mathrm{Aut}(M_{n})$ be the group of all automorphisms of $M_{n}$. An automorphism of $M_{n}$ of the form
+$(x_{1},x_{2},\ldots,x_{i-1},\alpha x_{i}+f,x_{i+1},\ldots x_{n}),$
+where $0\neq\alpha\in K$ and $f$ does not contain $x_{i}$, is called elementary.
+The subgroup $\mathrm{TAut}(M_{n})$ of $\mathrm{Aut}(M_{n})$ generated by all elementary automorphisms is called the tame automorphism group, and the elements of this subgroup are called tame automorphisms of $M_{n}$.
+Nontame automorphisms of $M_{n}$ are called wild.
+Defining relations among elementary automorphisms have been described for polynomial algebras in three variables [61, 62], for finitely generated free Lie algebras [64], and for free metabelian Lie algebras in three variables [33].
+As usual, the group of all linear automorphisms of $M_{n}$ will be identified with the group of all invertible matrices $GL_{n}(K)$.
+For every pair of integers $s,t$ with $1\leq s\neq t\leq n$, let $(st)$ denote the automorphism that simply permutes the variables $x_{s}$ and $x_{t}$ and leaves all other variables unchanged.
+The group generated by all $(st)$ is isomorphic to the symmetric group $S_{n}$.
+In 2015, Nauryzbaev [35] proved that the group of tame automorphisms of $M_{n}$ is generated by all linear automorphisms $GL_{n}(K)$ and the quadratic automorphism
 (2) $(x_{1}+[x_{2},x_{3}],x_{2},\ldots,x_{n})$
 if either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$, or $n\geq 5$ and $K$ is a field of arbitrary characteristic.
-In 1992 V.
-Drensky [17], in 1993 Papistas [43], and in 1995 Bahturin and Shpilrain [7] proved that free algebras of rank $n\geq 2$ of any polynilpotent variety of Lie algebras that is not abelian or metabelian have wild automorphisms.
-This result also makes the case of free metabelian Lie algebras more important.
-In 1992 Yu.
-Bahturin and S.
-Nabiyev [6] claimed that every nontrivial exponential automorphism of the free metabelian Lie algebra $M_{n}$ of any rank $n\geq 2$ over a field of characteristic zero is wild.
-More examples of wild automorphisms of $M_{n}$ of rank $n\geq 4$ were given in 2008 by Z.
-Özcurt and N.
-Ekici [42].
-For a long time, the area specialists believed these results were true.
-Unfortunately, both articles have recently been shown to contain fatal errors [71].
-Inner automorphisms of groups play a crucial role in the proofs of Bachmuth-Mochizuki-Roman’kov Theorem [5, 46]. The group automorphism
-$(x_{2}^{-1}x_{1}x_{2},x_{2},\ldots,x_{n})$
-is extreemly important in their proofs and there is no analogue of this automorphism in the case of free metabelian Lie algebras. This makes the case of algebras more complicated.
-One property of the free metabelian Lie algebras is that they admit non-elementary automorphisms that change only one variable and fix the rest.
-All automorphisms of $M_{n}$ of this type (including elementary ones) are called Chein automorphisms or one-row transformations [18, 48](see definitions in Section 2).
-The first example of the Chein automorphism is
-(2) $(x_{1}+[[x_{2},x_{3}],x_{1}],x_{2},\ldots,x_{n})$
-and no one could answer whether it is tame or not.
-For this reason, the specialists in the area easily accepted the wrong results [6, 42].
-This automorphism is wild if $n=3$ [36], and is the first candidate for wildness for all $n\geq 4$.
-Chein automorphisms were accepted as almost elementary automorphisms. An automorphism of $M_{n}$ is called almost tame if it is a product of Chein automorphisms [36, 38].
-In this paper we show that every automorphism of $M_{n}$ of rank $n\geq 4$ is almost tame.
-Moreover, $K$ is a field of characteristic $\neq 3$ then the group of all automorphisms $\mathrm{Aut}(M_{n})$ of $M_{n}$ of rank $n\geq 4$ is generated by all linear automorphisms, as well as a quadratic automorphism (1) and a cubic automorphism (2).
-If $n\geq 5$ then the same result holds for fields of any characteristic.
-One of the most important points of the proof is that we notice that all Chein automorphisms of lower degree $\geq 4$ and all exponential automorphisms of lower degree $\geq 5$ are tame.
-These results contradict the results of [6, 42].
-We follow the proof scheme given in [5], and in several places, we repeat the proofs for completeness. The real difficulties arise in those places where inner automorphisms were used in [5].
-Thus the question about the existence of wild automorphisms for a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ is still open as well as in the case polynomial and free associative algebras.
-But in this case it is reduced to a concrete automorphism.
-Problem 1.
-Is the Chein automorphism (2) of the free metabelian algebra Lie algebra $M_{n}$ of rank $n\geq 4$ tame?
-The paper is organized as follows.
-In Section 2 we define some important automorphisms and some subgroups of the automorphism group $\mathrm{Aut}(M_{n})$.
-In Section 3, using a representation of $M_{n}$ via Magnus embedding, we give the definition of the Fox derivatives and the Jacobian matrix of an endomorphism, and determine some of their properties.
-Using some representations of automorphisms from this section, in the following sections we make calculations in the language of the Jacobian matrices instead of automorphisms.
-For this, in Section 4 we clearly determine some matrix notations.
-In Section 5 we prove that a huge part of Chein and exponential automorphisms are tame.
-We begin the study of almost tame automorphisms in Section 6.
-Section 7 is devoted to the study of some special class of matrices that includes the Jacobian matrices of an important class of automorphisms.
-In Section 8 we prove a lemma on eliminating a variable in some special cases.
-Section 9 is devoted to the study of one exceptional automorphism
-(3) $(x_{1}+[[x_{1},x_{2}],x_{1}],x_{2}+[[x_{1},x_{2}],x_{2}],x_{3},\ldots,x_{n})$
-of $M_{n}$. In Section 10 we show how to eliminate a variable in the general case. In Section 11 the main result and some its corollaries are formulated.
-2. Automorphisms of $M_{n}$
-Let $L_{n}$ be a free Lie algebra over a field $K$ freely generated by $z_{1},z_{2},\ldots,z_{n}$.
-Then $Y_{n}=L_{n}/[L_{n},L_{n}]$ is an abelian Lie algebra with a linear basis $y_{1},y_{2},\ldots,y_{n}$ and $M_{n}=L_{n}/[[L_{n},L_{n}],[L_{n},L_{n}]]$ is a free metabelian Lie algebra with free generators $x_{1},\ldots,x_{n}$, where $y_{i}=z_{i}+[L_{n},L_{n}]$ and $x_{i}=z_{i}+[[L_{n},L_{n}],[L_{n},L_{n}]]$ for all $1\leq i\leq n$.
-Thus $M_{n}^{\prime}=[M_{n},M_{n}]$ is abelian and the set of all right normed (nonassociative) words of the form
-(4) $[\ldots[[x_{i},x_{i_{1}}],x_{i_{2}}],\ldots,x_{i_{k}}],\ i>i_{1}\leq i_{2}\leq\ldots\leq i_{k},$
-is a linear basis of $M_{n}^{\prime}$ (see, for example [8, p. 139]).
-Denote by $\phi=(f_{1},\ldots,f_{n})$ the endomorphism of $M_{n}$ such that $\phi(x_{i})=f_{i}$ for all $i$.
-In particular, $\mathrm{id}=(x_{1},\ldots,x_{n})$ is the identity automorphism.
-An automorphism of $M_{n}$ of the form
-$(x_{1},\ldots,x_{i-1},\alpha x_{i}+f,x_{i+1},\ldots,x_{n}),$
-where $0\neq\alpha\in K$ and $f$ does not depend on $x_{i}$, is called elementary.
-The subgroup $\mathrm{TAut}(M_{n})$ of $\mathrm{Aut}(M_{n})$ generated by all elementary automorphisms is called the tame automorphism group, and the elements of this subgroup are called tame automorphisms of $M_{n}$.
-Non-tame automorphisms of $M_{n}$ are called wild.
+If $\phi=(f_{1},\ldots,f_{n}),\psi=(g_{1},\ldots,g_{n})\in\mathrm{Aut}(M_{n})$, then the composition $\phi\psi$ is defined as
+$\phi\psi=(g_{1}(f_{1},\ldots,f_{n}),\ldots,g_{n}(f_{1},\ldots,f_{n})).$
+We also adopt the standard notation $\phi,\psi]=\phi\psi\phi^{-1}\psi^{-1}$ for the commutator, and $\phi^{\psi}=\psi\phi\psi^{-1}$ for conjugation.
+Let $0\neq z\in[M_{n},M_{n}]$.
+Then $\mathrm{ad}\,z:M_{n}\to M_{n}(m\mapsto[z,m])$ is a nilpotent derivation and $(\mathrm{ad}\,z)^{2}=0$.
+The exponential automorphism $\exp(\mathrm{ad}\,z)$ of $M_{n}$ can be written as
+$\exp(\mathrm{ad}\,z)=\mathrm{id}+\mathrm{ad}\,z=(x_{1}+[z,x_{1}],\ldots,x_{n}+[z,x_{n}]).$
+It is known that the exponential automorphism $\exp(\mathrm{ad}\,[x_{1},x_{2}])$ of $M_{3}$ is wild [14].
+Recall that in 1993 Bryant and Drensky proved [7] that every automorphism of $M_{n}$ of rank $n\geq 4$ over a field $K$ of characteristic zero is approximately tame ( see the definitions in [47]); that is, it can be approximated by tame automorphisms with respect to power series topology [7, 47].
+In contrast, all wild automorphisms of $M_{3}$ are absolutely wild [47], meaning that they cannot be approximated by tame automorphisms.
 Free metabelian Lie algebras admit automorphisms of the form
@@
 which include all elementary automorphisms but not necessarily elementary ones.
-They are called Chein automorphisms or one-row transformations [18].
+They are called Chein automorphisms or one-row transformations [15, 42].
 These automorphisms move only one variable and fix the others.
-The subgroup $\mathrm{ATAut}(M_{n})$ of $\mathrm{Aut}(M_{n})$ generated by all Chein automorphisms will be called the almost tame automorphism group, and the elements of this subgroup will be called called almost tame automorphisms of $M_{n}$ [36].
+The subgroup $\mathrm{ATAut}(M_{n})$ of $\mathrm{Aut}(M_{n})$ generated by all Chein automorphisms is called the almost tame automorphism group, and the elements of this subgroup is called almost tame automorphisms of $M_{n}$ [32, 34].
 Thus
 $\mathrm{TAut}(M_{n})\subseteq\mathrm{ATAut}(M_{n})\subseteq\mathrm{Aut}(M_{n}).$
-If $n=3$ then every non-elementary Chein automorphism is wild [36, 49] and the automorphism (3) is not almost tame [36].
-The case $n=3$ is well studied and defining relations for elementary automorphisms [37] and for Chein automorphisms [38] are given.
-As usual, the group of all linear automorphisms of $M_{n}$ will be identified with the group of all invertible matrices $GL_{n}(K)$.
-Recall that the symmetric group $S_{n}$ is embedded into $GL_{n}(K)$ and any permutation $\pi\in S_{n}$ represents the automorphism $\pi=(x_{\pi(1)},\ldots,x_{\pi(n)})$.
-For example, the transposition $(23)$ represents the automorphism $(23)=(x_{1},x_{3},x_{2},x_{4},\ldots,x_{n})$.
 Obviously, the group of almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated by all linear automorphisms $GL_{n}(K)$ and all Chein automorphisms of the form
-(5) $(x_{1}+f,x_{2},\ldots,x_{n}),\ f\in M_{n}^{2}.$
+(3) $(x_{1}+f,x_{2},\ldots,x_{n}),\ f\in[M_{n},M_{n}].$
 Chein automorphisms of this form are described in Section 3.
-The group of tame automorphisms $\mathrm{TAut}(M_{n})$ is generated by all linear automorphisms $GL_{n}(K)$ and all automorphisms of the form (5) where $f\in M_{n}^{2}$ does not depend on $x_{1}$.
-Moreover, if $n\geq 4$ and $K$ is a field of characteristic $\neq 3$, then $\mathrm{TAut}(M_{n})$ is generated [39] by all linear automorphisms $GL_{n}(K)$ and the quadratic elementary automorphism (1).
-If $n\geq 5$ then the same result holds over fields of any characteristic.
-Let $0\neq z\in[M_{n},M_{n}]$.
-Then $\mathrm{ad}\,z:M_{n}\to M_{n}(m\mapsto[z,m])$ is a locally nilpotent derivation and $(\mathrm{ad}\,z)^{2}=0$.
-The exponential automorphism $\exp(\mathrm{ad}\,z)$ of $M_{n}$ can be written as
-$\exp(\mathrm{ad}\,z)=\mathrm{id}+\mathrm{ad}\,z=(y_{1}+[z,y_{1}],\ldots,y_{n}+[z,y_{n}]).$
-It is known that the exponential automorphism $\exp(\mathrm{ad}\,[x_{1},x_{2}])$ is wild if $n=3$ [17, 36, 49].
+The lowest-degree example of a non-elementary Chein automorphism is
+(4) $(x_{1}+[[x_{2},x_{3}],x_{1}],x_{2},\ldots,x_{n}).$
+This automorphism is wild if $n=3$ [32, 43], and is the first candidate for wildness for all $n\geq 4$, but it is still unknown whether it is tame or not.
+It was recently shown [36] that an analogue of this automorphism for free metabelian anticommutative algebras is absolutely wild for all $n\geq 3$.
+In 2008, Roman’kov [43] proved that the group of automorphisms of $M_{3}$ cannot be finitely generated modulo the subgroup generated by all exponential and Chein automorphisms.
+Kabanov and Roman’kov [17] constructed strictly nontame primitive elements of $M_{3}$.
+Defining relations for Chein automorphisms of $M_{3}$ are given in [34].
+We introduce the notion of the lower degree of an automorphism. Consider the standard degree grading of
+$M_{n}=L_{1}\oplus L_{2}\oplus\ldots\oplus L_{k}\oplus\ldots,$
+where $L_{k}$ is the linear span of all Lie monomials of degree $k$ in $x_{1},\ldots,x_{n}$. Every nonzero $m\in M_{n}$ can be uniquely written in the form
+$m=l_{s}+l_{s+1}+\ldots+l_{t},$
+where $l_{j}\in L_{j}$ for all $s\leq j\leq t$, $l_{s}\neq 0$, and $l_{t}\neq 0$.
+We call $\mathrm{ldeg}(m)=s$ the lower degree of $m$, while $\deg(m)=t$ denotes the degree of $m$.
+If $\mathrm{ldeg}(m)=\deg(m)=t$, then $m$ is homogeneous of degree $t$.
+Set also $\mathrm{ldeg}(0)=\infty$ and $\deg(0)=$-$\infty$.
 Let
-$M_{n}=T_{1}\oplus T_{2}\oplus\ldots\oplus T_{k}\oplus\ldots$
-be the standard degree grading of $M_{n}$, i.e., $T_{k}$ is the linear span of (non-associative) monomials in $x_{1},x_{2},\ldots,x_{n}$ of degree $k$. If
-$f=f_{s}+f_{s+1}+\ldots+f_{t}\in M_{n},f_{j}\in T_{j},f_{s}\neq 0,f_{t}\neq 0,$
-then $\deg(f)=t$ is the degree and $\mathrm{ldeg}(f)=s$ is the lower degree of $f$ in the variables $x_{1},x_{2},\ldots,x_{n}$. Set also $\deg(0)=$-$\infty$ and $\mathrm{ldeg}(0)=\infty$.
-Let
-$\phi=(x_{1}+f_{1},x_{2}+f_{2},\ldots,x_{n}+f_{n}),$
-where $f_{i}\in M_{n}^{2}$ for all $1\leq i\leq n$.
-Then we define $\deg(\phi)$ as the maximum of all $\deg(f_{i}),1\leq i\leq n$, and $\mathrm{ldeg}(\phi)$ as the minimum of all $\mathrm{ldeg}(f_{i}),1\leq i\leq n$.
-If $\deg(\phi)=\mathrm{ldeg}(\phi)=k\geq 1$ then we say that $\phi$ is homogeneous of degree $k$.
-In particular, if $k=2$ then $\phi$ is called quadratic and if $k=3$ then it is called cubic.
-The group of automorphisms $\mathrm{IAut}(M_{n})$ of $M_{n}$ that induces the identity automorphism on $M_{n}/[M_{n},M_{n}]$ plays an important role in our study.
-Obviously, $\mathrm{IAut}(M_{n})$ is the set of all automorphisms of $M_{n}$ of lower degree $\geq 2$.
-The elements of $\mathrm{IAut}(M_{n})$ are called $\mathrm{IA}$-automorphisms.
-Here we describe a set of generators of the group of tame $\mathrm{IA}$-automorphisms and almost tame $\mathrm{IA}$-automorphisms.
-Lemma 1.
-$(a)$ The group of tame $\mathrm{IA}$-automorphisms $\mathrm{TAut}(M_{n})\cap\mathrm{IAut}(M_{n})$ is generated by all linear conjugates of elementary automorphisms of the form (5) where $f\in[M_{n},M_{n}]$ does not depend on $x_{1}$.
-$(b)$ The group almost tame $\mathrm{IA}$-automorphisms $\mathrm{TAut}(M_{n})\cap\mathrm{IAut}(M_{n})$ is generated by all linear conjugates of Chein automorphisms of the form (5).
-Proof. Every tame automorphism $\phi$ can be represented as
-$\phi=\alpha_{1}\phi_{1}\alpha_{2}\phi_{2}\ldots\phi_{k}\alpha_{k+1},$
-where $\phi_{i}$ is an elementary automorphism of the form (5) and $\alpha_{i}$ is linear for all $i$. Since $\phi\in\mathrm{IAut}(M_{n})$ for all $i$ it follows that
-$\alpha_{1}\alpha_{2}\ldots a_{k+1}=\mathrm{id}.$
-since $\phi_{i}\in\mathrm{IAut}(M_{n})$ for all $i$. Then $\alpha_{k+1}=(\alpha_{1}\ldots\alpha_{k})^{-1}$ and, consequently,
-$\phi=(\alpha_{1}\phi_{1}\alpha_{1}^{-1})(\alpha_{1}\alpha_{2}\phi_{2}(\alpha_{1}\alpha_{2})^{-1})\ldots(\alpha_{1}\ldots\alpha_{k}\phi_{k}(\alpha_{1}\ldots\alpha_{k})^{-1})$
-The statement $(b)$ of the lemma can be proved similarly. $\Box$
-3. Jacobian matrices
-In this section, we define and state some necessary information about Fox derivatives and Jacobian matrices from [45, 58, 62, 64].
-In fact, we define Fox derivatives more directly for free metabelian Lie algebras.
-First we give another representation of the free metabelian algebra $M_{n}$ over $K$ with free generators $x_{1},\ldots,x_{n}$.
-Let $Y_{n}$ be the abelian Lie algebra with a linear basis $y_{1},y_{2},\ldots,y_{n}$.
+$\phi=(l_{1}+m_{1},\ldots,l_{n}+m_{n})$
+be an automorphism of $M_{n}$ such that $l_{i}\in L_{1}$ and $m_{i}\in[M_{n},M_{n}]$ for all $i$. Set
+$\mathrm{ldeg}(\phi)=\mathrm{min}\{\mathrm{ldeg}(m_{i})|1\leq i\leq n\}$
+and
+$\mathrm{deg}(\phi)=\mathrm{max}\{\mathrm{deg}(m_{i})|1\leq i\leq n\}.$
+If $\mathrm{ldeg}(\phi)=\mathrm{deg}(\phi)=t$, then we say that $\phi$ is homogeneous of degree $t$.
+In particular, (2) is quadratic and (4) is cubic.
+Clearly, every nonlinear automorphism has a finite lower degree and degree.
+We show that all Chein automorphisms of lower degree $\geq 4$ are tame.
+This means that all Chein automorphisms ”except” (4) are tame.
+We also show that all exponential automorphisms of lower degree $\geq 5$ are tame.
+These results contradict the results of [5, 39].
+3. Chein automorphisms
+In this section, using an analogue of the well known Magnus embedding [19, 23], we give another representation of the free metabelian algebra $M_{n}$, and give direct definition of the Fox derivatives for $M_{n}$.
+This construction for Lie algebras is a particular case of Shmel’kin’s wreath products [51].
+Let $Y_{n}=M_{n}/[M_{n},M_{n}]$ be the abelian Lie algebra with a linear basis $y_{1},y_{2},\ldots,y_{n}$, where $y_{i}=x_{i}+[M_{n},M_{n}]$.
 The universal enveloping algebra $U(Y_{n})$ of $Y_{n}$ is the polynomial algebra $U=U(Y_{n})=K[y_{1},\ldots,y_{n}]$ in the variables $y_{1},y_{2},\ldots,y_{n}$.
@@
 $U=U_{0}\oplus U_{1}\oplus\ldots\oplus U_{k}\oplus\ldots$
-be the degree grading of $U$, i.e., $U_{k}$ is the linear span of all homogeneous of degree $k$ monomials in $y_{1},y_{2},\ldots,y_{n}$. If
+be the degree grading of $U$, i.e., $U_{k}$ is the linear span of all homogeneous monomials in $y_{1},y_{2},\ldots,y_{n}$ of degree $k$. If
 $a=a_{s}+a_{s+1}+\ldots+a_{t}\in U,a_{j}\in U_{j},a_{s}\neq 0,a_{t}\neq 0,$
 then $\deg(a)=t$ is the degree and $\mathrm{ldeg}(a)=s$ is the lower degree of $a$ in the variables $y_{1},y_{2},\ldots,y_{n}$. Set $\deg(0)=$-$\infty$ and $\mathrm{ldeg}(0)=\infty$ as usual.
-Let $T_{n}$ be a free right $U$-module with basis $t_{1},\ldots,t_{n}$. Turn the direct sum
+Let $T_{n}$ be the free right $U$-module with basis $t_{1},\ldots,t_{n}$. Turn the direct sum
 $M=Y_{n}\oplus T_{n}$
@@
 $x_{i}=y_{i}+t_{i},\ \ 1\leq i\leq n,$
-is the free metabelian Lie algebra $M_{n}$ with the free generators $x_{1},\ldots,x_{n}$.
-In the case of groups the corresponding embedding is called the Magnus embedding [23, 27] and in the case of Lie algebras it follows from Shmel’kin’s Theorem on wreath products [56].
+is the free metabelian Lie algebra with free generators $x_{1},\ldots,x_{n}$ [19, 23, 51]. We identify this algebra with $M_{n}$.
 Notice that $M_{n}^{\prime}=[M_{n},M_{n}]\subseteq T_{n}$ and for any $m\in M_{n}^{\prime}$ and $y_{i}$ we get
@@
 Let $e_{1},e_{2},\ldots,e_{n}$ be the columns of the identity matrix $I=I_{n}$ of order $n$.
-The statement of the next lemma is well known (see [62, Lemma 5],[64]).
+The statement of the next lemma is well known [56, 58].
 Lemma 1.
 Let $a=(a_{1},a_{2},\ldots,a_{n})^{t}\in(U^{n})^{t}$ be an arbitrary column of dimension $n$ over $U$. Then the following conditions are equivalent.
-$(1)$ $a=\partial(f)$ for some $f\in[M_{n},M_{n}]$;
-$(2)$ $Ya=y_{1}a_{1}+\ldots+y_{n}a_{n}=0$ where $Y=(y_{1},\ldots,y_{n})$;
-$(3)$ $a=\sum_{i<j}(e_{i}y_{j}-e_{j}y_{i})a_{ij},\ a_{ij}\in U$.
+$(a)$ $a=\partial(f)$ for some $f\in[M_{n},M_{n}]$;
+$(b)$ $Ya=y_{1}a_{1}+\ldots+y_{n}a_{n}=0$ where $Y=(y_{1},\ldots,y_{n})$;
+$(c)$ $a=\sum_{i<j}(e_{i}y_{j}-e_{j}y_{i})a_{ij},\ a_{ij}\in U$.
 Proof. If $i<j$ and $a\in U$, then
 $\partial([x_{i},x_{j}]a)=e_{i}y_{j}a-e_{j}y_{i}a=(e_{i}y_{j}-e_{j}y_{i})a.$
-Consequently, $(1)$ implies $(3)$. Obviously, $(3)$ implies $(2)$ since $Y(e_{i}y_{j}-e_{j}y_{i})=0$ for all $i<j$.
-Suppose that $(2)$ holds.
-Then $a_{n}=y_{1}b_{1}+\ldots+y_{n-1}b_{n-1}$ for some $b_{j}\in U$.
+Consequently, $(a)$ implies $(c)$. Obviously, $(c)$ implies $(b)$ since $Y(e_{i}y_{j}-e_{j}y_{i})=0$ for all $i<j$.
+Suppose that $(b)$ holds.
+Let $a_{i}=-y_{n}b_{i}+a_{i}^{0}$, where $a_{i}^{0}\in K[y_{1},\ldots,y_{n-1}]$ for all $1\leq i\leq n-1$.
+Then $(b)$ implies that $a_{n}=y_{1}b_{1}+\ldots+y_{n-1}b_{n-1}$.
 Set $g=[x_{n},x_{1}]b_{1}+\ldots+[x_{n},x_{n-1}]b_{n-1}$.
-Then $\frac{\partial g}{\partial x_{n}}=a_{n}$ and the last component of $a^{\prime}=a-\partial(g)$ is zero.
+We get $\frac{\partial g}{\partial x_{n}}=a_{n}$ and the last component of $a^{\prime}=a-\partial(g)$ is zero.
 Moreover, if $a^{\prime}=(a_{1}^{\prime},\ldots,a_{n-1}^{\prime},0)$ then $y_{1}a_{1}^{\prime}+\ldots+y_{n-1}a_{n-1}^{\prime}=0$.
@@
 This endomorphism will be denoted again by $\phi$ for simplicity of notation.
-If $f\in U$, then the image $\phi(f)$ will be denoted by $f^{\phi}$.
+If $f\in U$, then the image $\phi(f)$ will be denoted by $f^{\phi}$ for convenience.
 We define the Jacobian matrix of $\phi=(f_{1},\ldots,f_{n})$ by
@@
 If $\phi=(f_{1},\ldots,f_{n})$ and $\psi=(g_{1},\ldots,g_{n})$ then
-$\phi\circ\psi(x_{i})=g_{i}(f_{1},\ldots,f_{n})$
+$\phi\psi(x_{i})=g_{i}(f_{1},\ldots,f_{n})$
 for all $1\leq i\leq n$. The corresponding Chain Rule gives [53, 56, 58] that
-(6) $J(\phi\circ\psi)=J(\phi)J(\psi)^{\phi}.$
-The statement of the next lemma is also well known (see [58],[62, Theorem 2]).
+$J(\phi\psi)=J(\phi)J(\psi)^{\phi}.$
+Consequently, if $\phi$ is an automorphism, then $J(\phi)$ is invertible. The converse is also true [53, 56].
 Lemma 2.
 An endomorphism $\phi$ of $M_{n}$ is an automorphism if and only if the Jacobian matrix $J(\phi)$ is invertible.
-Here is a description of Chein automorphisms.
-Corollary 1.
+Proof. Suppose that $J(\phi)$ is invertible. Without loss of generality, we may assume that
+$\phi=(x_{1}+f_{1},\ldots,x_{n}+f_{n}),\ \ f_{1},\ldots,f_{n}\in[M_{n},M_{n}].$
+Then
+$J(\phi)=I+A,\ \ A=[\partial(f_{1}),\ldots,\partial(f_{n})].$
+We get
+$J(\phi)^{-1}=I+B,B=\sum_{i\geq 1}(-A)^{i}.$
+By Lemma 1, we have $YA=0$. Consequently, $YB=0$. Again, by Lemma 1, there exist $g_{1},\ldots,g_{n}\in[M_{n},M_{n}]$ such that
+$B=[\partial(g_{1}),\ldots,\partial(g_{n})].$
+Let $\psi=(x_{1}+g_{1},\ldots,x_{n}+g_{n})$. Then
+$J(\psi)=I+B=J(\phi)^{-1}$
+and
+$J(\phi\psi)=J(\phi)J(\psi)^{\phi}=J(\phi)J(\psi)=I.$
+Consequently, $\phi\psi=\mathrm{Id}$ and $\phi$ is an automorphism. $\Box$
+We now give a description of Chein automorphisms of the form (3).
+Lemma 3.
 An endomorphism of $M_{n}$ of the form (3) is an automorphism if and only if $f$ belongs to the ideal generated by all $x_{s},x_{t}]$, where $s,t\neq 1$.
-Proof.
-Let $\phi=(x_{1}+f,x_{2},\ldots,x_{n})$ where $f\in M_{n}^{2}$.
-The Jacobian matrix of $\phi$ is $J(\phi)=I+[\partial(f)\ 0\ldots 0]$.
+Proof. Let $\phi=(x_{1}+f,x_{2},\ldots,x_{n})$ where $f\in[M_{n},M_{n}]$. The Jacobian matrix of $\phi$ is
+$J(\phi)=I+[\partial(f)\ 0\ldots 0].$
 This matrix is invertible if and only if $1+\frac{\partial f}{\partial x_{1}}$ is invertible in $U$ or, equivalently, $\frac{\partial f}{\partial x_{1}}=0$.
-We show that $\frac{\partial f}{\partial x_{1}}=0$ if and only if $f$ belongs to the ideal generated by all $x_{s},x_{t}]$, where $s,t\neq 1$.
-The basis (4) of $M_{n}^{2}$ is convenient to work with $\frac{\partial f}{\partial x_{n}}=0$. Any $f\in M_{n}^{2}$ can be written in the form
+We want to show that $\frac{\partial f}{\partial x_{1}}=0$ if and only if $f$ belongs to the ideal generated by all $x_{s},x_{t}]$, where $s,t\neq 1$.
+The basis (1) of $M_{n},M_{n}]$ is convenient to work with $\frac{\partial f}{\partial x_{n}}=0$. Any $f\in[M_{n},M_{n}]$ can be written in the form
 $f=f^{\prime}+[x_{n},x_{1}]a_{1}+[x_{n},x_{2}]a_{2}+\ldots+[x_{n},x_{n-1}]a_{n-1},$
@@
 Continuing the same discussions, we get $a_{n-1}=a_{n-2}=\ldots=a_{1}=0$.
-Threfore $\frac{\partial f}{\partial x_{n}}=0$ if and only if $f=f^{\prime}$.
+Therefore $\frac{\partial f}{\partial x_{n}}=0$ if and only if $f=f^{\prime}$.
 $\Box$
-Consider the set of Chein automorphisms of the form
-(7) $(x_{1}+[x_{2},x_{3}]a,x_{2},\ldots,x_{n}),\ a\in U.$
-These automorphisms are elementary if $a$ does not depend on $y_{1}$.
-In this paper we denote by $\mathrm{Mat}_{n}(U)$ the set of all square matrices of order $n$ over $U$ (since the more standard notation $M_{n}(U)$ can be confused with the notation of the free metabelian Lie algebra $M_{n}$).
-As usual, denote by $GL_{n}(U)$ and $SL_{n}(U)$ the general linear group and the special linear group of matrices of order $n$ over $U$, respectively.
-The identity matrix of order $n$ is already denoted by $I$.
-In Section 2 we mentioned that the group $GL_{n}(K)$ will be identified with the group of all linear automorphisms of $M_{n}$.
-We meant that each linear automorphism $\alpha$ will be identified with its Jacobian matrix $J(\alpha)$.
-This means that $\alpha$ is a matrix and instead of $J(\alpha)$ in our calculations we will write $\alpha$.
-Lemma 4.
-$(a)$ If $\phi,\psi\in\mathrm{IAut}(M_{n})$ then
-$J(\phi\psi)=J(\phi)J(\psi);$
-$(b)$ If $\phi\in\mathrm{IAut}(M_{n})$ and $\alpha\in GL_{n}(K)$ then
-$J(\alpha\phi\alpha^{-1})=\alpha J(\phi)^{\alpha}\alpha^{-1};$
-$(c)$ The map
-$J:\mathrm{IAut}(M_{n})\to SL_{n}(U)$
-is an injective homomorphism of groups and $G=J(\mathrm{IAut}(M_{n}))$ is the set of all matrics from $SL_{n}(U)$ of the form
-$I+A,$
-where $A\in\mathrm{Mat}(U)$ is a matrix such that $YA=0$.
-Proof. Since $\phi$ induces an identity automorphism on $U$, $(a)$ is a direct consequence of (6). Using (6) we also get
-$J(\alpha\phi\alpha^{-1})=J(\alpha(\phi\alpha^{-1}))=J(\alpha)J(\phi\alpha^{-1})^{\alpha}=J(\alpha)(J(\phi)J(\alpha^{-1})^{\phi})^{\alpha}=J(\alpha)(J(\phi)J(\alpha^{-1}))^{\alpha}$
-$=J(\alpha)J(\phi)^{\alpha}J(\alpha^{-1})^{\alpha}=J(\alpha)J(\phi)^{\alpha}J(\alpha^{-1})=J(\alpha)J(\phi)^{\alpha}J(\alpha)^{-1}=\alpha J(\phi)^{\alpha}\alpha^{-1}.$
-In this calculation we again used that $\phi$ induces an identity automorphism on $U$, $J(\alpha^{-1})\in GL_{n}(K)$ is a constant matrix, and that $J(\alpha^{-1})=J(\alpha)^{-1}$ by (6).
-If $\phi\in\mathrm{IAut}(M_{n})$, then
-$\phi=(x_{1}+f_{1},x_{2}+f_{2},\ldots,x_{n}+f_{n}),$
-where $f_{i}\in[M_{n},M_{n}]$ for all $i$, and
-$J(\phi)=I+A,\ A=\left[\partial(f_{1})\ \ldots\ \partial(f_{n})\right].$
-We get $YA=0$ by Lemma 2.
-If $J(\phi)=I$ then $\partial(f_{i})=0$ for all $i$.
-Notice that for any $f\in[M_{n},M_{n}]$ we get $f=(t_{1},\ldots,t_{n})\partial(f)$ by the definition of $\partial(f)$.
-Consequently, $f_{i}=(t_{1},\ldots,t_{n})\partial(f_{i})=0$.
-This means $\phi=\mathrm{id}$, i.e., $J$ is injective.
-Conversely, let $A\in\mathrm{Mat}_{n}(U)$ such that $YA=0$.
-By Lemma 2 there exist $g_{1},g_{2},\ldots,g_{n}\in[M_{n},M_{n}]$ such that $A=\left[\partial(g_{1})\ \ldots\ \partial(g_{n})\right]$.
-Set $\psi=(x_{1}+g_{1},x_{2}+g_{2},\ldots,x_{n}+g_{n})$.
-Then $J(\psi)=I+A$.
-If $I+A\in SL_{n}(U)$ then $\psi$ is an automorphism by Lemma 3 and $\psi\in\mathrm{IAut}(M_{n})$.
-$\Box$
-4. Matrix notations
-Using Lemma 4, we will represent automorphisms from $\mathrm{IAut}(M_{n})$ by their Jacobian matrices and make calculations in the language of Jacobian matrices.
-Results formulated in the language of matrices will be also reformulated in the language of automorphisms.
-Thus $G$ represents the set of all Jacobian matrices of automorphisms from $\mathrm{IAut}(M_{n})$.
-Set $T=J(\mathrm{TAut}(M_{n})\cap\mathrm{IAut}(M_{n}))$ and $S=J(\mathrm{ATAut}(M_{n})\cap\mathrm{IAut}(M_{n}))$.
+4. Tame automorphisms
+In this section we always assume that either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$ or $n\geq 5$.
+Theorem 1.
+Every Chein automorphism of $M_{n}$ of lower degree $\geq 4$ is tame.
+Proof. It suffices to prove that every Chein automorphism of the form
+$C(a)=(x_{1}+[x_{2},x_{3}]a,x_{2},\ldots,x_{n}),\ a\in U,$
+with $\mathrm{ldeg}(a)\geq 2$ is tame. Since
+$C(a+b)=C(a)C(b),\ \ a,b\in U,$
+we may assume that $a=\gamma y_{1}^{i_{1}}y_{2}^{i_{2}}\ldots y_{n}^{i_{n}}$, $0\neq\gamma\in K$. If $i_{1}=0$ then $C(a)$ is elementary. Assume that $i_{1}\geq 1$.
+Case 1. Suppose that $i_{j}\geq 1$ for some $2\leq j\leq n-1$. Consider the elementary automorphisms
+$\phi=(x_{1},\ldots,x_{n-1},x_{n}+[x_{2},x_{3}]y_{1}^{i_{1}}\ldots y_{j}^{i_{j}-1}),$
+$\psi=(x_{1}-\gamma[x_{j},x_{n}]y_{j+1}^{i_{j+1}}\ldots y_{n}^{i_{n}},x_{2},\ldots,x_{n}).$
+Direct calculations give that $\phi,\psi]=C(a)$ is tame.
+Case 2.
+Let $i_{2}=i_{3}=\ldots=i_{n-1}=0$ and $i_{n}\neq 0$.
+Then $a=\gamma y_{1}^{i_{1}}y_{n}^{i_{n}}$.
+We can make $\gamma=1$ by conjugating $C(a)$ with $(x_{1},\gamma^{-1}x_{2},x_{3}\ldots,x_{n})$.
+Then assume that $a=y_{1}^{i_{1}}y_{n}^{i_{n}}$.
 We have
-$T\subseteq S\subseteq G.$
-In order to prove that $\mathrm{ATAut}(M_{n})=\mathrm{Aut}(M_{n})$ it is sufficient to prove that $\mathrm{ATAut}(M_{n})\cap\mathrm{IAut}(M_{n})=\mathrm{IAut}(M_{n})$ since both groups include all linear automorphisms.
-Thus, it is sufficient to prove that $S=G$.
-Let $E_{ij}$ a square matrix with $1$ in the $(i,j)$ position and zeros elsewhere. The order of $E_{ij}$ is usually clear from the context. Recall that
-$E_{ij}E_{kl}=\delta_{jk}E_{il},$
-where $\delta_{jk}$ is the Kronecker delta. A matrix of the form
-$I+aE_{ij},\ i\neq j,\ a\in U,$
-is called elementary.
-If $i<j$, then for any $k$ denote by $F_{ijk}$ the $n\times n$ matrix with $y_{j}$ in position $(i,k)$, $y_{i}$ in position $(j,k)$, and zeros in other places, that is,
-$F_{ijk}=y_{j}E_{ik}-y_{i}E_{jk}.$
-In other words, the $k$th column of $F_{ijk}$ is equal to $\partial([x_{i},x_{j}])=e_{i}y_{j}-e_{j}y_{i}$, and the rest are zeros. It is easy to check that
-$F_{ijk}F_{stm}=\left\{\begin{array}[]{ccc}y_{t}F_{ijm}&{\rm if}&k=s,\\ -y_{s}F_{ijm}&{\rm if}&k=t,\\ 0&{\rm if}&k\neq s,t\\ \end{array}\right$.,
-$E_{ij}F_{kst}=y_{s}E_{ij}E_{kt}-y_{k}E_{ij}E_{st}=\left\{\begin{array}[]{ccc}y_{s}E_{it}&{\rm if}&j=k,\\ -y_{k}E_{it}&{\rm if}&j=s,\\ 0&{\rm if}&j\neq k,s\\ \end{array}\right$.,
+(5) $[x_{2},x_{3}]a=[x_{2},x_{n}]y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1}+[x_{n},x_{3}]y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1}$
+by the Jacobi identity. If $i_{n}\geq 2$ then
+$\phi=(x_{1}+[x_{2},x_{n}]y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1},x_{2},\ldots,x_{n})=(x_{1}+[x_{2},x_{3}]y_{1}^{i_{1}}y_{3}^{i_{n}-1}y_{n},x_{2},\ldots,x_{n})^{(3n)}$
 and
-$F_{kst}E_{ij}=y_{s}E_{kt}E_{ij}-y_{k}E_{st}E_{ij}=\left\{\begin{array}[]{ccc}F_{ksj}&{\rm if}&t=i,\\ 0&{\rm if}&t\neq i\\ \end{array}\right.$.
-We refer to these three formulas as the multiplication tables but often use them without reference. Properties of the product of matrices of types
-$I+A+B=(I+A)(I+B),$
-which is true if $AB=0$, and
-$I+(a+b)C=(I+aC)(I+bC),$
-which is valid if $C^{2}=0$, to simplify terminology, will be called additivity.
-For example, if $f=\sum_{i<j}[x_{i},x_{j}]a_{ij}$, then the Jacobian matrix of the endomorphism (5) is $I+\sum_{i<j}a_{ij}F_{ij1}$, and, by additivity,
-$I+\sum_{i<j}a_{ij}F_{ij1}=\prod_{i<j}(I+a_{ij}F_{ij1}).$
-Moreover,
-$I+(a+b)F_{ij1}=(I+aF_{ij1})(I+bF_{ij1}).$
-Consider the set of matrices of the form
-(11) $I+aF_{231},\ a\in U.$
-These matrices are Jacobian matrices of Chein automorphisms (7).
-Here are two remarks:
-$(a)$ If $A\in G$ and $\alpha\in GL_{n}(K)$ then $\alpha A^{\alpha}\alpha^{-1}\in G$ by Lemma 4(b)(b).
-This matrix will be called the linear conjugate of $A$ by $\alpha$.
-In many places, conjugation by linear automorphisms helps to “linearize” automorphisms [73].
-$(b)$ If $J(\phi)=I+A\in S$ and $A$ is a homogeneous matrix, then, usually, it is easy to show that $I+\mu A\in S$ for any $\mu\in K$.
-This can be done by adding some coefficients in the proofs.
-For example, this can be done when conjugating matrices (11) using $(x_{1},\mu x_{2},x_{3},\ldots,x_{n})$.
-For this reason we ignore the difference between the coefficients of this type.
+$\psi=(x_{1}+[x_{n},x_{3}]y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1},x_{2},\ldots,x_{n})=(x_{1}+[x_{2},x_{3}]y_{1}^{i_{1}}y_{2}^{i_{n}-1}y_{n},x_{2},\ldots,x_{n})^{(2n)}$
+are tame by Case 1. Therefore $C(a)=\phi\psi$ is tame.
+If $i_{n}=1$, then set $\alpha=(x_{1},x_{2}+x_{n},x_{3},\ldots,x_{n})$. We have
+$\tau=(x_{1}+[x_{2},x_{3}]y_{1}^{i_{1}}y_{2},x_{2},\ldots,x_{n})$
+is tame by Case 1. Direct calculations give that
+$[\alpha,\tau]=(x_{1}+([x_{2},x_{3}]y_{n}+[x_{n},x_{3}]y_{2})y_{1}^{i_{1}},x_{2},\ldots,x_{n})(x_{1}+[x_{n},x_{3}]y_{1}^{i_{1}}y_{n},x_{2},\ldots,x_{n})$
+is tame. Since the second factor
+$(x_{1}+[x_{n},x_{3}]y_{1}^{i_{1}}y_{n},x_{2},\ldots,x_{n})=(x_{1}+[x_{2},x_{3}]y_{1}^{i_{1}}y_{2},x_{2},\ldots,x_{n})^{(2n)}=\tau^{(2n)}$
+of this product is also tame it follows that its first factor
+$\sigma=(x_{1}+([x_{2},x_{3}]y_{n}+[x_{n},x_{3}]y_{2})y_{1}^{i_{1}},x_{2},\ldots,x_{n})$
+is tame. By (5), we get
+$3[x_{2},x_{3}]a=([x_{2},x_{3}]y_{n}+[x_{n},x_{3}]y_{2})y_{1}^{i_{1}}+([x_{2},x_{3}]y_{n}+[x_{2},x_{n}]y_{3})y_{1}^{i_{1}}.$
+Consequently,
+$C(3a)=(x_{1}+([x_{2},x_{3}]y_{n}+[x_{n},x_{3}]y_{2})y_{1}^{i_{1}},x_{2},\ldots,x_{n})$
+$\times(x_{1}+([x_{2},x_{3}]y_{n}+[x_{2},x_{n}]y_{3})y_{1}^{i_{1}},x_{2},\ldots,x_{n})=\sigma(\sigma^{-1})^{(23)}$
+is tame. If the characteristic of $K$ is not $3$, then its conjugation by $(x_{1},3x_{2},x_{3},\ldots,x_{n})$ gives $C(a)$.
+If $n\geq 5$, then conjugating $C(a)$ by $(4n)$, we can make $i_{4}>0$. Case 1 gives that $C(a)$ is tame again.
+Case 3. Let $i_{2}=i_{3}=\ldots=i_{n-1}=i_{n}=0$. In this case $a=y_{1}^{i_{1}}$ and $s=i_{1}\geq 2$ since $\mathrm{ldeg}(a)\geq 2$. Case 2 gives that
+$\phi=(x_{1}+[x_{2},x_{3}]y_{1}^{s-1}y_{n},x_{2},\ldots,x_{n})$
+is tame. Set $\beta=(x_{1},,\ldots,x_{n-1},x_{n}+x_{1})$. Direct calculations give that
+$[\beta,\phi]=C(a)(x_{1},\ldots,x_{n-1},x_{n}-[x_{2},x_{3}]y_{1}^{s-1}(y_{1}+y_{n})$
+$=C(a)(x_{1}-[x_{2},x_{3}]y_{n}^{s},\ldots,x_{n})^{(1n)}(x_{1}-[x_{2},x_{3}]y_{1}y_{n}^{s-1},x_{2},\ldots,x_{n})^{(1n)}.$
+Since the second factor of this product is tame and the third factor is tame by Case 2 it follows that $C(a)$ is tame. $\Box$
+This theorem contradicts the results of [39].
 Corollary 1.
-$(a)$ The group $T$ is generated by all linear conjugates of matrices of the form (11) where $a$ does not depend on $y_{1}$.
-$(b)$ The group $S$ is generated by all linear conjugates of matrices of the form (11).
+The group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated, modulo all linear automorphisms, by the quadratic automorphism (2) and the cubic automorphism (4).
 Proof.
-By Lemma 1, every automorphism $\mathrm{TAut}(M_{n})\cap\mathrm{IAut}(M_{n})$ is a product of linear conjugates of automorphisms of the form (5).
-Notice that every automorphism of the form (5) is itself a product of linear conjugates of automorphisms of the form (7) with $a\in K[y_{2},\ldots,y_{n}]$.
-Then $J(\tau)$ is a matrix of the form (11).
-If $\alpha\in GL_{n}(K)$ then
-$J(\alpha\tau\alpha^{-1})=\alpha J(\tau)^{\alpha}\alpha^{-1}$
-by Lemma 4, that is, $J(\alpha\tau\alpha^{-1})$ is the linear conjugate of $J(\tau)$ by $\alpha$.
-If $\phi,\psi\in\mathrm{IAut}(M_{n})$ then $J(\phi\psi)=J(\phi)J(\psi)$ by the same lemma.
-This proves $(a)$.
-The statement $(b)$ can be checked similarly. $\Box$
-5. Tame automorphisms
-In this section we assume that either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$ or $n\geq 5$.
-Lemma 5.
-Every matrix of the form (11) belongs to $T$ if $\mathrm{ldeg}(a)\geq 2$.
-Proof. The proof practically repeats the proof of Theorem 1 from [39].
-Due to the additivity of (11) with respect to $a$, we can assume that $a=y_{1}^{i_{1}}y_{2}^{i_{2}}\ldots y_{n}^{i_{n}}$.
-If $i_{1}=0$ then $I+aF_{231}\in T$ by the definition of $T$.
-Assume that $i_{1}\geq 1$.
-Case 1. Suppose that $i_{j}\geq 1$ for some $2\leq j\leq n-1$. Set
-$A=I+y_{1}^{i_{1}}\ldots y_{j}^{i_{j}-1}F_{23n},\ \ B=I-y_{j+1}^{i_{j+1}}\ldots y_{n}^{i_{n}}F_{jn1}.$
-A direct calculation using the multiplication tables gives that
-$[A,B]=ABA^{-1}B^{-1}=I+aF_{231}\in T$
-since $A,B\in T$.
-Case 2. Let $i_{2}=i_{3}=\ldots=i_{n-1}=0$ and $i_{n}\neq 0$. Then $a=y_{1}^{i_{1}}y_{n}^{i_{n}}$ and we get
-$aF_{231}=y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1}F_{2n1}-y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1}F_{3n1}$
-by the Jacobi identity. If $i_{n}\geq 2$ then
-$I+y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1}F_{2n1},\ I-y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1}F_{3n1}\in T$
-by Case 1. Consequently,
-$I+aF_{231}=(I+y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1}F_{2n1})(I-y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1}F_{3n1})\in T.$
-If $i_{n}=1$, then set $\alpha=I+E_{42}$. Notice that $I+y_{1}^{i_{1}}y_{2}F_{231}\in T$ by Case 1. Hence
-$(I+E_{42})(I+y_{1}^{i_{1}}y_{2}F_{231})^{\alpha}(I-E_{42})=I+y_{1}^{i_{1}}y_{2}F_{231}+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341})-y_{1}^{i_{1}}y_{4}F_{341}$
-$=(I+y_{1}^{i_{1}}y_{2}F_{231})(I+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341}))(I-y_{1}^{i_{1}}y_{4}F_{341})\in T.$
-Since the first and the last factors of this product belong to $T$ it follows that
-$C=I+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341})\in T.$
-The Jacobi identity implies that
-$y_{4}F_{231}=y_{3}F_{241}-y_{2}F_{341}.$
-Consequently,
-$3aF_{231}=y_{1}^{i_{1}}[(y_{4}F_{231}-y_{2}F_{341})+(y_{4}F_{231}+y_{3}F_{241})]$
-and, by additivity,
-$I+3aF_{231}=[I+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341})][I+y_{1}^{i_{1}}(y_{4}F_{231}+y_{3}F_{241})].$
-The first factor is $C$ and belongs to $T$. The second factor also belongs to $T$ since
-$(23)(C^{-1})^{(23)}(23)=(23)((I+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341}))^{-1})^{(23)}(23)$
-$=(23)(I-y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341}))^{(23)}(23)=I+(y_{4}F_{231}+y_{3}F_{241}),$
-where $(23)=(x_{1},x_{3},x_{2},x_{4},\ldots,x_{n})$.
-Consequently, $I+3aF_{231}\in T$.
-If the characteristic of $K$ is not $3$, then its conjugation by $(x_{1},3x_{2},x_{3},\ldots,x_{n})$ gives that $I+aF_{231}\in T$.
-If $n\geq 5$, then conjugating (11) by $(4n)$, we can make $i_{4}>0$. Case 1 gives that $I+aF_{231}\in T$.
-Case 3.
-Let $i_{2}=i_{3}=\ldots=i_{n-1}=i_{n}=0$.
-In this case $a=y_{1}^{i_{1}}$ and $s=i_{1}\geq 2$ since $\mathrm{ldeg}(a)\geq 2$.
-Case 2 gives that $I+y_{4}y_{1}^{s-1}F_{231}\in T$.
-Set $\beta=I+E_{14}\in GL_{n}(K)$.
-Using the multiplication tables we get
-$\beta(I+y_{4}y_{1}^{s-1}F_{231})^{\beta}\beta^{-1}=I+(y_{4}+y_{1})y_{1}^{s-1}F_{231}-(y_{4}+y_{1})y_{1}^{s-1}F_{234}$
-$=(I+(y_{4}+y_{1})y_{1}^{s-1}F_{231})(I-(y_{4}+y_{1})y_{1}^{s-1}F_{234})$
-$=(I+y_{4}y_{1}^{s-1}F_{231})(I+y_{1}^{s}F_{231})(I-y_{4}y_{1}^{s-1}F_{234})(I-y_{1}^{s}F_{234})\in T.$
-Since all factors of this product except $I+y_{1}^{s}F_{231}$ belong to $T$, this implies that $I+y_{1}^{s}F_{231}$ also belongs to $T$. $\Box$
-Here is a reformulation of this lemma in the language of automorphisms.
-Corollary 3.
-Every Chein automorphism of the form (5) with $\mathrm{ldeg}(f)\geq 4$ is tame.
-Proof.
-By Corollary 1, any such Chein automorphism is a product of linear conjugates of Chein automorphisms of the form (7) with $\mathrm{ldeg}(a)\geq 2$.
-Then $J(\tau)$ is a matrix of the form (11) and belongs to $T$ by Lemma 5.
-$\Box$
-This corollary contradicts the results of [42].
-Corollary 4.
-Assume that either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$ or $n\geq 5$.
-Then the group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated by all linear automorphisms, as well as a quadratic automorphism (1) and a cubic automorphism (2).
-Proof.
-The group of all tame automorphisms $\mathrm{TAut}(M_{n})$ is generated by all linear automorphisms, as well as a quadratic automorphism (1) [39].
-All Chein automorphisms of the form (7) with $\mathrm{ldeg}(a)\geq 2$ belong to $\mathrm{TAut}(M_{n})$.
-By the additivity of automorphisms of the form (7) with respect to $a$, we may assume that $a$ is homogeneous of degree $\leq 1$.
-If $a$ does not depend on $y_{1}$ then (7).
-Again, by additivity, we may assume that $a=y_{1}$.
-In this case (7) is the cubic automorphism (2).
+The group of all tame automorphisms $\mathrm{TAut}(M_{n})$ is generated, modulo all linear automorphisms, by the quadratic automorphism (2) [35].
+Obviuosly, the group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated, modulo all linear automorphisms, by Chein automorphisms $C(a)$.
+We have
+$C(a)=C(a_{0})C(a_{1})\ldots C(a_{k}),$
+where $a_{i}\in U_{i}$ for all $0\leq i\leq k$. We have $C(a_{0})$ is elementary since $a_{0}\in K$ and $C(a_{2}),\ldots,C(a_{k})$ are tame by Theorem 1.
+Let
+$a_{1}=\alpha_{1}y_{1}+\alpha_{2}y_{2}+\ldots+\alpha_{n}y_{n}.$
+Then
+$C(a_{1})=C(\alpha_{1}y_{1})C(\alpha_{2}y_{2})\ldots C(\alpha_{n}y_{n}).$
+All $C(\alpha_{2}y_{2}),\ldots,C(\alpha_{n}y_{n})$ are again tame.
+Consider $C(\alpha_{1}y_{1})$. Conjugating this by $(x_{1},\alpha_{1}^{-1}x_{2},x_{3},\ldots,x_{n})$ we can make $\alpha_{1}=1$ if $\alpha_{1}\neq 0$.
+Consequently, the group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ modulo $\mathrm{TAut}(M_{n})$ is generated by $C(y_{1})$.
+Note that $C(y_{1})$ is the cubic automorphism (4).
+Therefore the group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated, modulo all linear automorphisms, by the quadratic automorphism (2) and the cubic automorphism (4).
 $\Box$
 Lemma 4.
-Every matrix of the form
-$A(a)=I+ay_{1}F_{121}+ay_{2}F_{122}$
-belongs to $T$ if $\mathrm{ldeg}(a)\geq 2$.
-Proof. By Lemma 5,
-$B=I-by_{1}F_{341}-by_{2}F_{342}=(I-by_{1}F_{341})(I-by_{2}F_{342})\in T$
-if $\mathrm{ldeg}(b)\geq 1$. Set $C=I+F_{123}\in T$. Then
-$[B,C]=BCB^{-1}C^{-1}=I+by_{1}y_{4}F_{121}+by_{2}y_{4}F_{122}=A(y_{4}b)\in T.$
-Consequently, the statement of the lemma is true if $y_{4}$ divides $a$.
-Obviously, the same is true if $y_{i}$ divides $a$ for all $3\leq i\leq n$.
-By the additivity of $A(a)$ with respect to $a$, we may assume that $a\in K[y_{1},y_{2}]$.
-Set $\alpha=I+E_{14}\in T$. If $b\in K[y_{1},y_{2}]$ then, using the multiplication tables, we get
-$\alpha A(y_{4}b)^{\alpha}\alpha^{-1}=I+y_{1}(y_{4}+y_{1})bF_{121}+y_{2}(y_{4}+y_{1})bF_{122}-y_{2}(y_{4}+y_{1})bF_{124}$
-$=A(y_{4}b)A(y_{1}b)(I-y_{2}(y_{4}+y_{1})bF_{124})\in T.$
-Since $A(y_{4}b)\in T$ and $I-y_{2}(y_{4}+y_{1})bF_{124}\in T$ by Lemma 5, this implies that $A(y_{1}b)\in T$.
-Therefore the statement of the lemma is true if $a\in K[y_{1},y_{2}]$ and $y_{1}$ divides $a$.
-Obviously, the same is true for $y_{2}$.
-$\Box$
-Corollary 5.
+Every automorphism of $M_{n}$ of the form
+$D(a)=(x_{1}+[x_{1},x_{2}]y_{1}a,x_{2}+[x_{1},x_{2}]y_{2}a,x_{3},\ldots,x_{n})$
+is tame if $\mathrm{ldeg}(a)\geq 2$.
+Proof. By Theorem 1, if $\mathrm{ldeg}(b)\geq 1$, then
+$\phi=(x_{1}-[x_{3},x_{4}]y_{1}b,x_{2},\ldots,x_{n})(x_{1},x_{2}-[x_{3},x_{4}]y_{2}b,x_{3},\ldots,x_{n})$
+$=(x_{1}-[x_{3},x_{4}]y_{1}b,x_{2}-[x_{3},x_{4}]y_{2}b,x_{3},\ldots,x_{n})$
+is tame. Let
+$\psi=(x_{1},x_{2},x_{3}+[x_{1},x_{2}],x_{4},\ldots,x_{n}).$
+Then
+$[\phi,\psi]=(x_{1}+[x_{1},x_{2}]y_{1}y_{4}b,x_{2}+[x_{1},x_{2}]y_{2}y_{4}b,x_{3},\ldots,x_{n})=D(y_{4}b)$
+is tame.
+Consequently, the statement of the lemma is true if $y_{4}$ divides $a$. Obviously, the same is true if $y_{i}$ divides $a$ for all $3\leq i\leq n$. Since
+$D(a_{1}+a_{2})=D(a_{1})+D(a_{2}),\ \ a_{1},a_{2}\in U,$
+we may assume that $a\in K[y_{1},y_{2}]$. Let $\alpha=(x_{1},x_{2},x_{3},x_{4}+x_{1},x_{5},\ldots,x_{n})$. If $b\in K[y_{1},y_{2}]$ then
+$D(y_{4}b)^{\alpha}=(x_{1}+[x_{1},x_{2}]y_{1}(y_{4}+y_{1})b,x_{2}+[x_{1},x_{2}]y_{2}(y_{4}+y_{1})b,$
+$x_{3},x_{4}-[x_{1},x_{2}]y_{2}(y_{4}+y_{1}),x_{5},\ldots,x_{n})$
+$=D(y_{4}b)D(y_{1}b)(x_{1},x_{2},x_{3},x_{4}-[x_{1},x_{2}]y_{2}(y_{4}+y_{1}),x_{5},\ldots,x_{n}).$
+Since $D(y_{4}b)$ is tame and
+$(x_{1},x_{2},x_{3},x_{4}-[x_{1},x_{2}]y_{2}(y_{4}+y_{1}),x_{5},\ldots,x_{n})$
+is tame by Theorem 1, this implies that $D(y_{1}b)$ is also tame. Therefore the statement of the lemma is true if $a\in K[y_{1},y_{2}]$ and $y_{1}$ divides $a$.
+Changing $\alpha$ to $\beta=(x_{1},x_{2},x_{3},x_{4}+x_{2},x_{5},\ldots,x_{n})$ and repeating the same discussions as above, we can show that $D(y_{2}b)$ is also tame. $\Box$
+Theorem 2.
+Every exponential automorphism of $M_{n}$ of lower degree $\geq 5$ is tame.
+Proof. Let
+$E(m)=\exp(\mathrm{ad}(m))=(x_{1}+[m,x_{1}],x_{2}+[m,x_{2}],\ldots,x_{n}+[m,x_{n}])$
+be an exponential automorphism with $\mathrm{ldeg}(m)\geq 4$. Since
+$E(m_{1}+m_{2})=E(m_{1})+E(m_{2}),$
+we may assume that $m=[x_{1},x_{2}]a$ with $\mathrm{ldeg}(a)\geq 2$. Then
+$E(m)=E([x_{1},x_{2}]a)=D(a)\phi_{3}\ldots\phi_{n},$
+where
+$\phi_{i}=(x_{1},\ldots,x_{i-1},x_{i}+[x_{1},x_{2}]y_{i}a,x_{i+1},\ldots,x_{n}),\ 3\leq i\leq n.$
+All $\phi_{i}$ are tame by Theorem 1 and $D(a)$ is tame by Lemma 4. Consequently, $E(m)$ is tame. $\Box$
+This theorem contradicts the main result of [5].
+5. Almost tame automorphisms
+In this section we prove that a huge class of automorphisms of $M_{n}$ that moves only two variables $x_{1}$ and $x_{2}$ are almost tame.
+We always assume that $n\geq 4$ and $K$ is an arbitrary field.
+Since all Chein automorphisms are automatically included into $\mathrm{ATAut}(M_{n})$ by the definition, we don’t need any restrictions to characteristic of the field from the preceeding section.
+First we give minor extensions of Lemma 4 and Theorem 2 for almost tame automorphisms.
+Lemma 5.
 Every automorphism of the form
-$(x_{1}+[x_{1},x_{2}]y_{1}a,x_{2}+[x_{1},x_{2}]y_{2}a,x_{3},\ldots,x_{n})$
+$D(a)=(x_{1}+[x_{1},x_{2}]y_{1}a,x_{2}+[x_{1},x_{2}]y_{2}a,x_{3},\ldots,x_{n})$
 is tame if $\mathrm{ldeg}(a)\geq 1$.
-Proof. By Lemma 6 the Jacobian matrices of these automorphisms belong to $T$. $\Box$
-Corollary 6.
-Every exponential automorphism of lower degree $\geq 5$ is tame.
-Proof. Let
-$\exp(\mathrm{ad}(m))=(x_{1}+[m,x_{1}],x_{2}+[m,x_{2}],\ldots,x_{n}+[m,x_{n}])$
-be an exponential automorphism with $\mathrm{ldeg}(m)\geq 4$.
-By the additivity with respect to $m$ we may assume that $m=[x_{1},x_{2}]a$ with $\mathrm{ldeg}(a)\geq 2$.
-Then the Jacobian matrix of $\exp(\mathrm{ad}(m))$ is
-$I+\sum_{i}y_{i}aF_{12i}=(I+y_{1}aF_{121}+y_{2}aF_{122})\prod_{i\geq 3}(I+y_{i}aF_{12i}).$
-This matrix belongs to $T$ by Lemmas 5 and 6. $\Box$
-This corollary contradicts the main result of [6].
-6. Almost tame automorphisms
-For the remainder of this article, we always assume that $n\geq 4$ and $K$ is an arbitrary field.
-By Corollary 2, every matrix of the form (11) and its conjugates belong to $S$.
-For this reason we do not need the proof of Lemma 5, where the condition that either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$ or $n\geq 5$ was used.
-The other proofs of Section 5 did not use this condition.
-Lemma 7.
-Every matrix of the form
-$C(a)=I+ay_{1}F_{121}+ay_{2}F_{122}$
-with $\mathrm{ldeg}(a)\geq 1$ belongs to $S$.
 Proof. For any $b\in U$ we have
-$B=I-by_{1}F_{341}-by_{2}F_{342}=(I-by_{1}F_{341})(I-by_{2}F_{342})\in S.$
-Repeating the proof of Lemma 6 with this data, we get that $C(a)\in S$ for all $a$ with $\mathrm{ldeg}(a)\geq 1$. $\Box$
-Notice that Lemma 7 does not include only the matrix
-(12) $I+y_{1}F_{121}+y_{2}F_{122}$
-of the same type, which is the Jacobian matrix of (3).
+$\phi=(x_{1}-[x_{3},x_{4}]y_{1}b,x_{2},\ldots,x_{n})(x_{1},x_{2}-[x_{3},x_{4}]y_{2}b,x_{3},\ldots,x_{n})$
+$=(x_{1}-[x_{3},x_{4}]y_{1}b,x_{2}-[x_{3},x_{4}]y_{2}b,x_{3},\ldots,x_{n})$
+is almost tame. Repeating the proof of Lemma 4 with this data, we get that $D(a)$ is almost tame for all $a$ with $\mathrm{ldeg}(a)\geq 1$. $\Box$
 Corollary 2.
@@
 Proof.
-Using the decomposition of the Jacobian matrix of an exponential automorphism from the proof of Corollary 6, we conclude that it belongs to $S$ by Lemma 7 if its lower degree $\geq 3$.
+Consider the decomposition of the exponential automorphism $E([x_{1},x_{2}]a)$ from the proof of Theorem 2.
+Note that $\mathrm{ldeg}(E([x_{1},x_{2}]a))\geq 4$ implies $\mathrm{ldeg}(a)\geq 1$.
+Consequently, $D(a)$ is almost tame by Lemma 5.
+Since all $\phi_{i}$ are Chein automorphisms it follows that $E([x_{1},x_{2}]a)$ is almost tame.
 Consequently, every exponential automorphism of lower degree $\geq 4$ belongs to $\mathrm{ATAut}(M_{n})$.
 $\Box$
-7. Special matrices
-We use matrices instead of matrix notation in many places because it makes the calculations easier to understand. To some extent this is due to the application of Suslin’s theorem [61].
-The main aim of this section is to prove that every matrix of the form
-(13) $I+\begin{bmatrix}y_{n}C&\begin{matrix}0\\ .\\ .\\ .\\ 0\end{matrix}\\ \begin{matrix}*&*&\ldots&*\end{matrix}&0\end{bmatrix}\in G$
-belongs to $S$. We will provide several lemmas to the proof of this result. A matrix of the form (13) will be called special.
 Lemma 6.
-Let $B$ be any matrix from $SL_{n-1}(U,y_{n}U)$. Then $B=I_{n-1}+y_{n}C$ for some matrix $C\in\mathrm{Mat}_{n-1}(U)$ and there exists a unique special matrix $A$ such that
-$A=I+\left[\begin{array}[]{cc}y_{n}C&0\\ &0\\ \end{array}\right].$
-Proof. Obviously, $B=I_{n-1}+y_{n}C$ for some matrix $C\in\mathrm{Mat}_{n-1}(U)$. If $C=[c_{ij}]$ then
-$A=I+\begin{bmatrix}y_{n}D&\begin{matrix}0\\ .\\ .\\ .\\ 0\end{matrix}\\ \begin{matrix}-\sum_{i=1}^{n-1}y_{i}d_{i1}&-\sum_{i=1}^{n-1}y_{i}d_{i2}&\ldots&-\sum_{i=1}^{n-1}y_{i}d_{i(n-1)}\end{matrix}&0\end{bmatrix}$
-belongs to $G$ by Lemma 2. If $C=0$ then $A=I$ by Lemma 2. This implies the uniqueness of $A$. $\Box$
-If $B\in SL_{n}(U)$ is a matrix of the form $B=\begin{bmatrix}C&0\\ 0&I_{n-k}\end{bmatrix}$, where $C\in SL_{k}(U)$ for some $1\leq k\leq n$, then we say that $B$ is $k$-reducible and $B_{k}=C$ will be called its $k$-reduction.
-Calculations with $k$-reducible matrices from $G$ can be done by using their $k$-reductions.
-Set $A_{n-1}=B$ for matrices $A,B$ satisfying the conditions of Lemma 8.
-We use this notation for special matrices despite the fact that $A$ is not n1n-1-reducible in general.
-But Lemma 16 allows us to use the matrix $B$ instead of $C$ in some calculations.
-Moreover, if $B$ is $k$-reducible for some $1\leq k\leq n-1$, then $B_{k}$ will also be called the $k$-reduction of the special matrix $A$ and the special matrix $A$ will be called special $k$-reducible.
-Let us specify how can we use $k$-reductions of special matrices.
-Using (6) it is easy to check the following statements:
-$(1)$ If $A$ and $B$ are two special $k$-reducible matrices, then $AB$ is a special $k$-reducible and $(AB)_{k}=A_{k}B_{k}$;
-$(2)$ If $A$ is special $k$-reducible and $B\in SL_{n}(U)$ (not necessarily $B\in G$) is $k$-reducible, then $A^{B}=BAB^{-1}$ is special $k$-reducible and $(A^{B})_{k}=B_{k}A_{k}B_{k}^{-1}$;
-$(3)$ If $A$ is special $k$-reducible and $\alpha\in SL_{n}(K)$ is $k$-reducible, then the linear conjugation $\alpha A^{\alpha}\alpha^{-1}$ of $A$ by $\alpha$ is special $k$-reducible and $(\alpha A^{\alpha}\alpha^{-1})_{k}=\alpha_{k}A_{k}^{\alpha_{k}}\alpha_{k}^{-1}$.
-Lemma 9.
-Every special matrix of the form
-$A(h,g)=I+hgF_{1n1}+hg^{2}F_{2n1}-hF_{1n2}-hgF_{2n2}$
-belongs to $S$.
-Proof. We have
-$A(h,g)_{2}=I_{2}+y_{n}h\begin{bmatrix}g&-1\\ g^{2}&-g\end{bmatrix}$.
-$(a)$ Let $\alpha=I-\lambda E_{21},\lambda\in K$. Then
-$A(h,g)_{2}^{\alpha}=I_{2}+y_{n}h^{\alpha}\begin{bmatrix}g^{\alpha}&-1\\ (g^{\alpha})^{2}&-g^{\alpha}\end{bmatrix}$
+Every automorphism of the form
+$A(h,g)=(x_{1}+[x_{1},x_{n}]hg+[x_{2},x_{n}]hg^{2},x_{2}-[x_{1},x_{n}]h-[x_{2},x_{n}]hg,x_{3},\ldots,x_{n}),$
+where $g,h\in U$, is almost tame.
+Proof. $(a)$ Let $\alpha=(x_{1}-\lambda x_{2},x_{2},\ldots,x_{n})$. Then
+$A(h,g)^{\alpha}=A(h^{\alpha},g^{\alpha}-\lambda).$
+This means that $A(h,g)$ is almost tame if and only if $A(h^{\alpha},g^{\alpha}-\lambda)$ is almost tame. Using this we may assume that $g(0,\ldots,0)=0$.
+$(b)$ Consider the almost tame automorphisms
+$\phi=(x_{1}-[x_{2},x_{3}]a,x_{2},\ldots,x_{n})$
 and
-$\alpha_{2}A(h,g)_{2}^{\alpha}\alpha_{2}^{-1}=\begin{bmatrix}1&0\\ -\lambda&1\end{bmatrix}\left(I_{2}+y_{n}h^{\alpha}\begin{bmatrix}g^{\alpha}&-1\\ (g^{\alpha})^{2}&-g^{\alpha}\end{bmatrix}\right)\begin{bmatrix}1&0\\ \lambda&1\end{bmatrix}$
-$=I_{2}+y_{n}h^{\alpha}\begin{bmatrix}(g^{\alpha}-\lambda)&-1\\ (g^{\alpha}-\lambda)^{2}&-(g^{\alpha}-\lambda)\end{bmatrix}=A(h^{\alpha},(g^{\alpha}-\lambda))_{2}.$
-Consequently, $\alpha A(h,g)^{\alpha}\alpha^{-1}=A(h^{\alpha},(g^{\alpha}-\lambda))$.
-This means that $A(h,g)\in S$ if and only if $A(h^{\alpha},(g^{\alpha}-\lambda))\in S$.
-Using this we may assume that $g(0,\ldots,0)=0$.
-$(b)$ Set $B=I-fF_{231}\in S$. Then
-$B_{3}=\begin{bmatrix}1&0&0\\ -y_{3}f&1&0\\ y_{2}f&0&1\end{bmatrix}$.
+$\psi=(x_{1}+[x_{3},x_{n}]y_{2}ah(g-y_{3}a),x_{2},\ldots,x_{n})(x_{1},x_{2}-[x_{3},x_{n}]y_{2}ah,x_{3},\ldots,x_{n})$
+$=(x_{1}+[x_{3},x_{n}]y_{2}ah(g-y_{3}a),x_{2}-[x_{3},x_{n}]y_{2}ah,x_{3},\ldots,x_{n})$
 Direct calculations give that
-$B_{3}A(h,g)_{3}B_{3}^{-1}=I_{3}+y_{n}h\begin{bmatrix}(g-y_{3}f)&-1&0\\ (g-y_{3}f)^{2}&-(g-y_{3}f)&0\\ y_{2}(g-y_{3}f)f&-y_{2}f&0\end{bmatrix}$
-[…]
-$=A(h,g-y_{3}f)_{3}C_{3},$
-where
-$C=I+y_{2}h(g-y_{3}f)fF_{3n1}-y_{2}hfF_{3n2}=(I+y_{2}h(g-y_{3}f)fF_{3n1})(I-y_{2}hfF_{3n2}).$
-Notice that $C\in S$.
-Consequently, $A(h,g)\in S$ if and only if $A(h,g-y_{3}f)\in S$.
+$A(h,g)^{\phi}=A(h,g-y_{3}a)\psi.$
+Consequently, $A(h,g)$ is almost tame if and only if $A(h,g-y_{3}f)$ is almost tame.
 Of course, the same trick can be done with any variable $y_{i}$ if $3\leq i\leq n-1$.
-Therefore, replacing $g$ by $g-y_{i}f$ for all $3\leq i\leq n-1$ we may assume that
-$g=G(y_{1},y_{2})-y_{n}a.$
-$(c)$ Now we show how to eliminate $y_{n}a$. First we replace $g$ to
-$g+y_{3}a=G(y_{1},y_{2})+y_{3}a-y_{n}a$
-using $(b)$. If $\gamma=I+E_{3n}$, then it is easy to check that
-$\gamma A(h,g+y_{3}a)^{\gamma}\gamma^{-1}=A(h^{\gamma},(g+y_{3}a)^{\gamma}).$
+Therefore, replacing $g$ by $g-y_{i}f$ for all $3\leq i\leq n-1$, we may assume that
+$g=G(y_{1},y_{2})-y_{n}b,\ \ b\in U.$
+$(c)$ Now we show how to eliminate $y_{n}b$. First we replace $g$ to
+$g+y_{3}b=G(y_{1},y_{2})+y_{3}b-y_{n}b$
+using $(b)$. If $\beta=(x_{1},x_{2},x_{3}+x_{n},x_{4},\ldots,x_{n})$, then it is easy to check that
+$A(h,g+y_{3}b)^{\beta}=A(h^{\beta},(g+y_{3}b)^{\beta}).$
 We have
-$(g+y_{3}a)^{\gamma}=g^{\gamma}+(y_{3}+y_{n})a^{\gamma}$
-$=G(y_{1},y_{2})-y_{n}a^{\gamma}+(y_{3}+y_{n})a^{\gamma}=G(y_{1},y_{2})+y_{3}a^{\gamma}.$
+$(g+y_{3}b)^{\beta}=g^{\beta}+(y_{3}+y_{n})b^{\beta}$
+$=G(y_{1},y_{2})-y_{n}b^{\beta}+(y_{3}+y_{n})b^{\beta}=G(y_{1},y_{2})+y_{3}b^{\beta}.$
 Using $(b)$ again, we may assume that
 $g=G(y_{1},y_{2}).$
-(d) Let $f$ be a polynomial in $y_{1},y_{2}$.
-Then $g$ can be replaced by $g+y_{i}f$, where $i=1,2$.
-For this we first replace $g$ by $g+y_{3}f$ using $(b)$, i.e., we can assume that $g=G(y_{1},y_{2})+y_{3}f$.
-If $\delta=I+E_{i3}$, then
-$\delta A(h,g)^{\delta}\delta^{-1}=(I-h^{\delta}g^{\delta}F_{1n1}-h^{\delta}(g^{\delta})^{2}F_{2n1})A(h^{\delta},g^{\delta}).$
-Since $I-h^{\delta}g^{\delta}F_{1n1}-h^{\delta}(g^{\delta})^{2}F_{2n1}\in S$ it follows that $A(h,g)\in S$ if and only if $A(h^{\delta},g^{\delta})\in S$.
-Notice that
-$g^{\gamma}=G(y_{1},y_{2})+y_{3}f^{\gamma}+y_{i}f^{\gamma}.$
-Using $(b)$, we can replace $g^{\gamma}$ by
-$G(y_{1},y_{2})+y_{i}f^{\gamma}.$
-This means that we can replace $g=G(y_{1},y_{2})$ by $G(y_{1},y_{2})+y_{i}f$ for any $f$. Then we can make $g=0$. Notice that $A(h,0)=I-hF_{1n2}\in S$. $\Box$
-Corollary 8.
+(d) Let $c$ be a polynomial in $y_{1},y_{2}$.
+Then $g$ can be replaced by $g+y_{i}c$, where $i=1,2$.
+For this, we first replace $g$ by $g+y_{3}c$ using $(b)$, i.e., we can assume that $g=G(y_{1},y_{2})+y_{3}c$.
+If $\gamma=(x_{1},x_{2},x_{3}+x_{1},x_{4},\ldots,x_{n})$, then
+$A(h,g)^{\gamma}=\sigma A(h^{\gamma},g^{\gamma}),$
+where
+$\sigma=(x_{1},x_{2},x_{3}-[x_{1},x_{n}]h^{\gamma}g^{\gamma}-[x_{2},x_{n}]h^{\gamma}(g^{\gamma})^{2},x_{4},\ldots,x_{n})$
+$=(x_{1},x_{2},x_{3}-[x_{1},x_{n}]h^{\gamma}g^{\gamma},x_{4},\ldots,x_{n})(x_{1},x_{2},x_{3}-[x_{2},x_{n}]h^{\gamma}(g^{\gamma})^{2},x_{4},\ldots,x_{n})$
+is an almost tame automorphism. Consequently, $A(h,g)$ is almost tame if and only if $A(h^{\gamma},g^{\gamma})$ is almost tame. Note that
+$g^{\gamma}=G(y_{1},y_{2})+y_{3}c+y_{1}c.$
+Using $(b)$, again we can replace $g^{\gamma}$ by
+$G(y_{1},y_{2})+y_{1}c.$
+This means that we can replace $g=G(y_{1},y_{2})$ by $G(y_{1},y_{2})+y_{1}c$ for any $c\in K[y_{1},y_{2}]$.
+If $\delta=(x_{1},x_{2},x_{3}+x_{2},x_{4},\ldots,x_{n})$ and $g=G(y_{1},y_{2})+y_{3}c$, then
+$A(h,g)^{\delta}=\tau A(h^{\delta},g^{\delta}),$
+where
+$\tau=(x_{1},x_{2},x_{3}+[x_{1},x_{n}]h^{\delta}+[x_{2},x_{n}]h^{\delta}g^{\delta},x_{4},\ldots,x_{n})$
+$=(x_{1},x_{2},x_{3}+[x_{1},x_{n}]h^{\delta},x_{4},\ldots,x_{n})(x_{1},x_{2},x_{3}+[x_{2},x_{n}]h^{\delta}g^{\delta},x_{4},\ldots,x_{n})$
+is again almost tame automorphism. In this case we have
+$g^{\delta}=G(y_{1},y_{2})+y_{3}c+y_{2}c.$
+Using $(b)$, we can replace $g^{\delta}$ by
+$G(y_{1},y_{2})+y_{2}c.$
+This means that we can replace $g=G(y_{1},y_{2})$ by $G(y_{1},y_{2})+y_{2}c$ for any $c\in K[y_{1},y_{2}]$.
+Thus we can make $g=0$. Observe that
+$A(h,0)=(x_{1},x_{2}-[x_{1},x_{n}]h,x_{3},\ldots,x_{n})$
+is a Chein automorphism. $\Box$
+Theorem 3.
 Every automorphism of the form
-$(x_{1}+[x_{1},x_{n}]hg+[x_{2},x_{n}]hg^{2},x_{2}+[x_{1},x_{n}]h+[x_{2},x_{n}]hg,x_{3},\ldots,x_{n})$
-belongs to $W$.
-Lemma 10.
-Every matrix of the form
-$A(h,f,g)=I+hfgF_{1n1}+hg^{2}F_{2n1}-hf^{2}F_{1n2}-hfgF_{2n2}$
-belongs to $S$.
-Proof. The proof of the lemma repeats the proof given on pages 90–91 of [5]. Notice that $A(h,f,g)$ is an induced $2$-reducible matrix and
-$A(h,f,g)_{2}=I_{2}+y_{n}h\begin{bmatrix}fg&-f^{2}\\ g^{2}&-fg\end{bmatrix}$
-We have
-$A(h,f,g)_{3}=\begin{bmatrix}1&0&0\\ 0&1&0\\ g&-f&1\end{bmatrix}\begin{bmatrix}1&0&-y_{n}hf\\ 0&1&-y_{n}hg\\ 0&0&1\end{bmatrix}\begin{bmatrix}1&0&0\\ 0&1&0\\ -g&f&1\end{bmatrix}\begin{bmatrix}1&0&y_{n}hf\\ 0&1&y_{n}hg\\ 0&0&1\end{bmatrix}$.
-In matrix notation, this can be written as
-$A(h,f,g)=[(I+gE_{31}-fE_{32}),(I-hfF_{1n3}-hgF_{2n3})]$
-$=(I-hfF_{1n3}-hgF_{2n3})^{(I+gE_{31}-fE_{32})}(I+hfF_{1n3}+hgF_{2n3})$
-The last factor belongs to $S$. Therefore it is sufficient to prove that
-$(I-hfF_{1n3}-hgF_{2n3})^{(I+gE_{31}-fE_{32})}=(I-hfF_{1n3})^{(I+gE_{31}-fE_{32})}(I-hgF_{2n3})^{(I+gE_{31}-fE_{32})}$
-belongs to $S$.
-Moreover, it is sufficient to prove that the first factor of the last product belongs to $S$ since its conjugation by the linear automorphism $(12)$ gives a matrix that has the form of the second factor.
-We have
-$\begin{bmatrix}1&0&0\\ 0&1&0\\ g&-f&1\end{bmatrix}\begin{bmatrix}1&0&-y_{n}hf\\ 0&1&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}1&0&0\\ 0&1&0\\ -g&f&1\end{bmatrix}=I_{3}+y_{n}h\begin{bmatrix}fg&-f^{2}&-f\\ 0&0&0\\ fg^{2}&-f^{2}g&-fg\end{bmatrix}$
-[…]
-which means that
-$(I-hfF_{1n3})^{(I+gE_{31}-fE_{32})}=(I-hf^{2}F_{1n2}-hf^{2}gF_{3n2})$
-$\times(I+(hf)gF_{1n1}+(hf)g^{2}F_{3n1}-(hf)F_{1n3}-(hf)gF_{3n3}).$
-The first factor of this product belongs to $S$ and the second factor belongs to $S$ by Lemma 9. $\Box$
-Corollary 9.
-Every automorphism of the form
-$(x_{1}+[x_{1},x_{n}]hfg+[x_{2},x_{n}]hg^{2},x_{2}+[x_{1},x_{n}]hf^{2}+[x_{2},x_{n}]hfg,x_{3},\ldots,x_{n})$
-belongs to $W$.
-First, let us recall some results of Cohn [13] and Suslin [61] on generators of special linear groups over polynomial algebras.
-In 1966 P. Cohn [13] proved that the matrix
-$\begin{bmatrix}1+x_{1}x_{2}&x_{2}^{2}\\ -x_{1}^{2}&1-x_{1}x_{2}\\ \end{bmatrix}$
-is not a product of elementary matrices over the polynomial algebra $K[x_{1},x_{2}]$.
-Denote by $E_{k}(U)$ the subgroup of $SL_{k}(U)$ generated by all elementary matrices, that is, by all matrices of the form
-$I_{k}+aE_{ij},\ a\in U.$
-The well-known Suslin’s Theorem [61] states that if $k\geq 3$ then $SL_{k}(U)=E_{k}(U)$, that is, every matrix from $SL_{k}(U)$ is a product of elementary matrices.
-If $I$ is an ideal of $U$ then the natural map $U\to U/I$ induces the natural maps $GL_{k}(U)\to GL_{k}(U/I)$ and $GL_{k}(U)\to GL_{k}(U/I)$.
-Denote by $GL_{k}(U,I)$ the kernel of $GL_{k}(U)\to GL_{k}(U/I)$ and by $SL_{k}(U,I)$ the kernel of $SL_{k}(U)\to SL_{k}(U/I)$.
-Denote by $E_{k}(U,I)$ the normal closure in $E_{k}(U)$ of the set of all elementary matrices of the form $I+aE_{ij}$, where $i\neq j$ and $a\in I$.
-The next corollary of Suslin’s results is formulated in [5].
-Proposition 1.
-Let $k\geq 3$ and $I=y_{n}U$. Then
-(1) $SL_{k}(U,I)=E_{k}(U,I)$;
-(2) $E_{k}(U,I)$ is generated by the set of all matrices of the form
-$I+h(fe_{i}+ge_{j})(ge_{i}-fe_{j})^{t}$
-where $i<j$, $h\in I$, and $f,g\in U$.
-Proof.
-We repeat a short proof given in [5].
-By [61], every matrix $\Phi\in SL_{k}(U,I)$ is a product of elementary matrices $\Phi=E_{1}\ldots E_{s}$.
-Any $a\in U=K[y_{1},\ldots,y_{n}]$ can be written in the form
-$a=y_{n}b+c,\ \ b\in U,c\in K[y_{1},\ldots,y_{n-1}],$
-and, consequently,
-$I+aE_{ij}=(I+y_{n}bE_{ij})(I+cE_{ij}).$
-Thus, $E_{i}=H_{i}G_{i}$ where $H_{i}\in E_{k}(U,I)$ and $G_{i}\in E_{k}(K[y_{1},\ldots,y_{n-1}])$. Then
-$\Phi=H_{1}(G_{1}H_{2}G_{1}^{-1})(G_{1}G_{2}H_{3}G_{2}^{-1}G_{1}^{-1})\ldots(G_{1}\ldots G_{s-1}H_{s}G_{s-1}^{-1}\ldots G_{1}^{-1})(G_{1}\ldots G_{s}).$
-Substituting $y_{n}=0$ we get $G_{1}\ldots G_{s}=1$. Since
-$G_{1}\ldots G_{i-1}H_{i}G_{i-1}^{-1}\ldots G_{1}^{-1}\in E_{k}(U,I)$
-it follows that $\Phi\in E_{k}(U,I)$.
-By [61, Corollary 1.4], the group $E_{k}(U,I)$ is generated by the set of all matrices of the form
-$I_{k}+hu(u_{j}e_{i}-u_{i}e_{j})^{t}$
-where $h\in I$, $u=(u_{1},\ldots,u_{k})$ is a unimodular column over $U$, and $1\leq i<j\leq k$.
-Recall that a column $u$ is called unimodular if the ideal generated by its components contains the identity element.
-We have
-$I_{k}+hu(u_{j}e_{i}-u_{i}e_{j})^{t}=(I_{k}+h(u_{i}e_{i}+u_{j}e_{j})(u_{j}e_{i}-u_{i}e_{j})^{t})$
-$\prod_{l\neq i,j}(I_{k}+hu_{l}e_{l}(u_{j}e_{i})^{t})\prod_{l\neq i,j}(I_{k}-h(u_{l}e_{l})^{t}u_{i}e_{j}).$
-This proves $(b)$. $\Box$
-Notice that if $i=1,j=2$, then the matrix from Proposition 1(2)(2) is $2$-reducible and its $2$-reduction is
-$I_{2}+h\left[\begin{array}[]{cc}fg&-f^{2}\\ g^{2}&-fg\\ \end{array}\right],$
-which is a generalization of the Cohn matrix mentioned above.
-Proposition 2.
-Every special matrix belongs to $S$.
-Proof. Since $n\geq 4$ it follows by Proposition 1 that $SL_{n-1}(U,y_{n}U)$ is generated by the set of all matrices
-$B=I_{n-1}+y_{n}h(fe_{i}+ge_{j})(ge_{i}-fe_{j})^{t},\ 1\leq i<j\leq n-1.$
-It is sufficient to prove that for any such matrix $B$ there exists a special matrix $A$ from $S$ such that its special (n1)(n-1)-reduction is $B$.
-Without loss of generality we may assume that $i=1$ and $j=2$, that is,
-$B=I_{n-1}+y_{n}h(fe_{1}+ge_{2})(ge_{1}-fe_{2})^{t}.$
-Then for
-$A=I+hfgF_{1n1}+hg^{2}F_{2n1}-hf^{2}F_{1n2}-hfgF_{2n2}$
-we get $A_{n-1}=B$ and $A$ belongs to $S$ by Lemma 10. $\Box$
-Corollary 10.
-Every automorphism of the form
-$(x_{1}+f_{1},x_{2}+f_{2},\ldots,x_{n-1}+f_{n-1},x_{n}),$
-where $f_{1},f_{2},\ldots,f_{n-1}$ belong to the ideal generated by $x_{1},x_{n}],\ldots,[x_{n-1},x_{n}]$, belongs to $W$.
-8. Inductive lemma
-The following lemma is part of the excellent proof given in [5, p.
-99–100], and we will repeat this for completeness.
-Since we have no exponential automorphisms of degree $3$, we first present this statement with a slight restriction.
-Using this lemma, in the next section we will deal with exponential automorphisms of degree $3$.
-Then we can prove this lemma in full form.
-Fortunately, the proof given here works in the general case.
-Lemma 11.
-Let $A=I+[a_{ij}]\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)\cap G$, $\mathrm{ldeg}(a_{ij})\geq 3$, and $a_{1n},a_{2n},\ldots,a_{(n-1)n}\in\sum_{p=s}^{n-1}y_{p}U+y_{n}^{2}U$.
-Then there exists $B\in S$ such that $B^{-1}A\in SL_{n}(U,\sum_{p=s}^{n-1}y_{p}U)$.
-Proof. For any $a\in U$ and any matrix $R$ over $U$ let $\overline{a}$ and $\overline{R}$ be the result of setting $y_{s}=y_{s+1}=\ldots=y_{n-1}=0$ in $a$ and $R$ respectively.
-Let
-$M=\begin{bmatrix}I_{n-1}&\begin{matrix}0\\ .\\ .\\ .\\ 0\end{matrix}\\ \begin{matrix}y_{1}&y_{2}&\ldots&y_{n-1}\end{matrix}&y_{n}\end{bmatrix}$.
+$B(h,f,g)=(x_{1}+[x_{1},x_{n}]hfg+[x_{2},x_{n}]hg^{2},x_{2}-[x_{1},x_{n}]hf^{2}-[x_{2},x_{n}]hfg,x_{3},\ldots,x_{n})$
+is almost tame.
+Proof. Consider the Chein automorphisms
+$\phi_{1}=(x_{1},x_{2}-[x_{1},x_{n}]hf^{2}-[x_{3},x_{n}]hf^{2}g,x_{3},\ldots,x_{n}),$
+$\psi_{1}=(x_{1}+[x_{2},x_{n}]hg^{2}-[x_{3},x_{n}]hfg^{2},x_{2},\ldots,x_{n})$
+and the automorphisms
+$\phi_{2}=(x_{1}+[x_{1},x_{n}]hfg+[x_{3},x_{n}]hfg^{2},x_{2},x_{3}-[x_{1},x_{n}]hf-[x_{3},x_{n}]hfg,x_{4},\ldots,x_{n}),$
+$\psi_{2}=(x_{1},x_{2}-[x_{2},x_{n}]hfg+[x_{3},x_{n}]hf^{2}g,x_{3}-[x_{2},x_{n}]hg+[x_{3},x_{n}]hfg,x_{4},\ldots,x_{n}).$
+Note that
+$(\phi_{2})^{(23)}=A((hf)^{(23)},g^{(23)})\ \hbox{and}\ (\psi)^{132)}=A((hg)^{(132)},-f^{(132)})$
+are almost tame by Lemma 6. Then $\phi_{2}$ and $\psi_{2}$ are also almost tame. Conseqently,
+$\phi=\phi_{1}\phi_{2}=(x_{1}+[x_{1},x_{n}]hfg+[x_{3},x_{n}]hfg^{2},x_{2}-[x_{1},x_{n}]hf^{2}-[x_{3},x_{n}]hf^{2}g,$
+$x_{3}-[x_{1},x_{n}]hf-[x_{3},x_{n}]hfg,x_{4},\ldots,x_{n})$
+and
+$\psi=(x_{1}+[x_{2},x_{n}]hg^{2}-[x_{3},x_{n}]hfg^{2},x_{2}-[x_{2},x_{n}]hfg+[x_{3},x_{n}]hf^{2}g,$
+$x_{3}-[x_{2},x_{n}]hg+[x_{3},x_{n}]hfg,x_{4},\ldots,x_{n})$
+are almost tame. Direct calculations give that
+$\phi\psi=(x_{1}+[x_{1},x_{n}]hfg+[x_{2},x_{n}]hg^{2},x_{2}-[x_{1},x_{n}]hf^{2}-[x_{2},x_{n}]hfg,$
+$x_{3}-[x_{1},x_{n}]hf-[x_{2},x_{n}]hg,x_{4},\ldots,x_{n}).$
 Then
-$\overline{M}=\begin{bmatrix}I_{n-1}&\begin{matrix}0\\ .\\ .\\ .\\ 0\end{matrix}\\ \begin{matrix}y_{1}&\ldots&y_{s-1}&0&\ldots&0\end{matrix}&y_{n}\end{bmatrix}$.
-We have
-$MAM^{-1}=I_{n}+\begin{bmatrix}[a_{ij}-y_{n}^{-1}y_{j}a_{in}]&\begin{matrix}y_{n}^{-1}a_{1n}\\ .\\ .\\ .\\ y_{n}^{-1}a_{(n-1)n}\end{matrix}\\ \begin{matrix}0&\ldots&0\end{matrix}&0\end{bmatrix}$
-[…]
-[…]
-and
-[…]
-where $\overline{d}_{ij}\in K[y_{1},\ldots,y_{s-1},y_{n}]$.
-The last equality is possible since $\overline{a}_{ij}\in y_{n}K[y_{1},\ldots,y_{s-1},y_{n}]$ for all $i,j$ and $\overline{a}_{in}\in y_{n}^{2}K[y_{1},\ldots,y_{s-1},y_{n}]$ for all $1\leq i\leq n-1$.
-Moreover, notice that $\mathrm{ldeg}(\overline{d}_{ij})\geq 2$ for all $1\leq i,j\leq n-1$ and $\mathrm{ldeg}(\overline{d}_{in})\geq 1$ for all $1\leq i\leq n-1$ since $\mathrm{ldeg}(\overline{a}_{ij})\geq 3$ for all $i,j$.
-For any $1\leq k\leq n-1$, let
-$\delta_{k}=I_{n}+\begin{bmatrix}0&0&\ldots&0\\ .&.&\ldots&.\\ .&.&\ldots&.\\ .&.&\ldots&.\\ 0&0&\ldots&0\\ y_{1}y_{n}\overline{d}_{kn}&y_{2}y_{n}\overline{d}_{kn}&\ldots&y_{n}^{2}\overline{d}_{kn}\\ 0&0&\ldots&0\\ .&.&\ldots&.\\ .&.&\ldots&.\\ .&.&\ldots&.\\ 0&0&\ldots&0\\ -y_{1}y_{k}\overline{d}_{kn}&-y_{2}y_{k}\overline{d}_{kn}&\ldots&-y_{n}y_{k}\overline{d}_{kn}\end{bmatrix}$
-$=\prod_{l\neq k,n}(I_{n}+y_{l}\overline{d}_{kn}F_{knl})\cdot(I_{n}+y_{k}\overline{d}_{kn}F_{knk}-y_{n}\overline{d}_{kn}F_{knn}).$
-Notice that all factors of the product $\prod_{l\neq k,n}(I_{n}+y_{l}\overline{d}_{kn}F_{knl})$ belong to $S$ and
-$I_{n}+y_{k}\overline{d}_{kn}F_{knk}-y_{n}\overline{d}_{kn}F_{knn}\in S$
-by Lemma 7 since $\mathrm{ldeg}(\overline{d}_{kn})\geq 1$ (this is the only place where the restriction $\mathrm{ldeg}(a_{ij})\geq 3$ was used). Consequently, $\delta_{k}\in S$.
-Notice that
-$\overline{M}\overline{\delta}_{1}\ldots\overline{\delta}_{n-1}\overline{M}^{-1}=\begin{bmatrix}I_{n-1}&\begin{matrix}y_{n}\overline{d}_{1n}\\ .\\ .\\ .\\ y_{n}\overline{d}_{(n-1)n}\end{matrix}\\ \begin{matrix}0&\ldots&0\end{matrix}&0\end{bmatrix}$
-and
-$\overline{M}(\overline{\delta}_{1}\ldots\overline{\delta}_{n-1})^{-1}\Phi\overline{M}^{-1}=\begin{bmatrix}[y_{n}\overline{d}_{ij}]&\begin{matrix}0\\ .\\ .\\ .\end{matrix}\\ \begin{matrix}0&\ldots\end{matrix}&0\end{bmatrix}$.
-Let
-$\epsilon=I_{n}+\begin{bmatrix}y_{n}\overline{d}_{11}&\ldots&y_{n}\overline{d}_{1(n-1)}&0\\ .&\ldots&.&.\\ .&\ldots&.&.\\ .&\ldots&.&.\\ y_{n}\overline{d}_{(n-1)1}&\ldots&y_{n}\overline{d}_{(n-1)(n-1)}&0\\ -\sum_{j=1}^{n-1}y_{j}\overline{d}_{j1}&\ldots&-\sum_{j=1}^{n-1}y_{j}\overline{d}_{j(n-1)}&0\end{bmatrix}$.
-By Proposition 2, $\epsilon\in S\cap SL_{n}(U,\sum_{j=s}^{n}y_{j}U)$, and
-$M\epsilon M^{-1}=\overline{M}\overline{\epsilon}\overline{M}^{-1}=\begin{bmatrix}[y_{n}\overline{d}_{ij}]&\begin{matrix}0\\ .\\ .\\ .\end{matrix}\\ \begin{matrix}0&\ldots\end{matrix}&0\end{bmatrix}$.
-Consequently,
-$\overline{M}\ \overline{\epsilon}^{-1}\ (\overline{\delta}_{1}\ldots\overline{\delta}_{n-1})^{-1}\overline{A}\ \overline{M}^{-1}=\overline{M}\ \overline{B}^{-1}\ \overline{A}\ \overline{M}^{-1}=I_{n},$
-where $B=\delta_{1}\ldots\delta_{n-1}\epsilon$. This means that $B^{-1}A\in S\cap SL_{n}(U,\sum_{p=s}^{n-1}y_{p}U)$. $\Box$
-9. Exceptional automorphism
-In 1981 Shafarevich [51] and in 1983 Anick [1] independently proved that every automorphism of the polynomial algebra $K[x_{1},x_{2},\ldots,x_{n}]$ over a field $K$ of characteristic zero can be approximated by tame automorphisms with respect to two different but similar topologies.
-This can be considered as a weaker generalization of the Jung-van der Kulk Theorem for polynomial algebras of rank $n\geq 3$.
-In particular, the Nagata automorphism is wild but can be approximated by tame automorphisms.
-An analogue of the Anick-Shafarevich Theorem [1, 51] for free metabelian Lie algebras was proven in 1993 by Bryant and Drensky [9].
-They proved that the group of tame automorphisms is dense in the group of all automorphisms of $M_{n}$ if $n\geq 4$.
-A sharpened version of this result can be found in [21].
-Now we prove that the automorphism (3) is almost tame using the ideas of approximation [9] and the results of the preceding sections. Recall that the Jacobian matrix of (3) is the matrix (12).
-Lemma 12.
-The matrix (12) belongs to $S$.
-Proof. We have
-$A=I+y_{2}F_{231}-\lambda y_{1}F_{231}\in S.$
-Let $\alpha=I+\lambda E_{12}$. Then $B=\alpha A^{\alpha}\alpha^{-1}\in S$. Since $A$ and $\alpha$ are $3$-reducible it follows that $B$ is $3$-reducible and
-$B_{3}=\alpha_{3}A_{3}^{\alpha}\alpha_{3}^{-1}=\begin{bmatrix}1&\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}1&0&0\\ (y_{2}-\lambda y_{1})y_{3}&1&0\\ -(y_{2}-\lambda y_{1})y_{2}&0&1\end{bmatrix}^{\alpha}\begin{bmatrix}1&-\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$
-$=\begin{bmatrix}1&\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}1&0&0\\ y_{2}y_{3}&1&0\\ -y_{2}(y_{2}+\lambda y_{1})&0&1\end{bmatrix}\begin{bmatrix}1&-\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$
-$=\begin{bmatrix}1+\lambda y_{2}y_{3}&-\lambda^{2}y_{2}y_{3}&0\\ y_{2}y_{3}&1-\lambda y_{2}y_{3}&0\\ -y_{2}(y_{2}+\lambda y_{1})&\lambda y_{2}(y_{2}+\lambda y_{1})&1\end{bmatrix}$
-Set $C=(I-y_{2}F_{231})(I+\lambda^{2}y_{2}F_{132})B$. Then
-$C_{3}=\begin{bmatrix}1&0&0\\ -y_{2}y_{3}&1&0\\ y_{2}^{2}&0&1\end{bmatrix}\begin{bmatrix}1&\lambda^{2}y_{2}y_{3}&0\\ 0&1&0\\ 0&-\lambda^{2}y_{2}y_{1}&1\end{bmatrix}\begin{bmatrix}1+\lambda y_{2}y_{3}&-\lambda^{2}y_{2}y_{3}&0\\ y_{2}y_{3}&1-\lambda y_{2}y_{3}&0\\ -y_{2}(y_{2}+\lambda y_{1})&\lambda y_{2}(y_{2}+\lambda y_{1})&1\end{bmatrix}$
-$=\begin{bmatrix}1+\lambda y_{2}y_{3}+\lambda^{2}y_{2}^{2}y_{3}^{2}&-\lambda^{3}y_{2}^{2}y_{3}^{2}&0\\ -\lambda y_{2}^{2}y_{3}^{2}(1+\lambda y_{2}y_{3})&1-\lambda y_{2}y_{3}+\lambda^{3}y_{2}^{3}y_{3}^{3}&0\\ \lambda(-y_{1}y_{2}+y_{2}^{3}y_{3})(1+\lambda y_{2}y_{3})&\lambda y_{2}^{2}+\lambda^{3}y_{1}y_{2}^{2}y_{3}-\lambda^{3}y_{2}^{4}y_{3}^{2}&1\end{bmatrix}$
-$=I_{3}+\lambda\begin{bmatrix}y_{2}y_{3}&0&0\\ 0&-y_{2}y_{3}&0\\ -y_{1}y_{2}&y_{2}^{2}&0\end{bmatrix}+y_{2}^{2}y_{3}C^{\prime},$
+$B(h,f,g)=\tau\phi\psi,$
 where
-$C^{\prime}=\begin{bmatrix}\lambda^{2}y_{3}&-\lambda^{3}y_{3}&0\\ -\lambda y_{3}(1+\lambda y_{2}y_{3})&\lambda^{3}y_{2}y_{3}^{2}&0\\ \lambda(y_{2}-\lambda y_{1}+\lambda y_{2}^{2}y_{3})&\lambda^{3}y_{1}-\lambda^{3}y_{2}^{2}y_{3}&0\end{bmatrix}$.
-Set also $D=(I+y_{2}F_{231})(I-\lambda^{2}y_{2}F_{132})B^{-1}$. Then
-$D_{3}=\begin{bmatrix}1-\lambda y_{2}y_{3}+\lambda^{2}y_{2}^{2}y_{3}^{2}&-\lambda^{3}y_{2}^{2}y_{3}^{2}&0\\ -\lambda y_{2}^{2}y_{3}^{2}(1-\lambda y_{2}y_{3})&1+\lambda y_{2}y_{3}-\lambda^{3}y_{2}^{3}y_{3}^{3}&0\\ \lambda(y_{1}y_{2}+y_{2}^{3}y_{3})(1-\lambda y_{2}y_{3})&-\lambda y_{2}^{2}+\lambda^{3}y_{1}y_{2}^{2}y_{3}+\lambda^{3}y_{2}^{4}y_{3}^{2}&1\end{bmatrix}$
-and
-$(23)_{3}D_{3}^{(23)}(23)_{3}=\begin{bmatrix}1-\lambda y_{2}y_{3}+\lambda^{2}y_{2}^{2}y_{3}^{2}&0&-\lambda^{3}y_{2}^{2}y_{3}^{2}\\ \lambda(y_{1}y_{3}+y_{2}y_{3}^{3})(1-\lambda y_{2}y_{3})&1&-\lambda y_{3}^{2}+\lambda^{3}y_{1}y_{2}y_{3}^{2}+\lambda^{3}y_{2}^{2}y_{3}^{4}\\ -\lambda y_{2}^{2}y_{3}^{2}(1-\lambda y_{2}y_{3})&0&1+\lambda y_{2}y_{3}-\lambda^{3}y_{2}^{3}y_{3}^{3}\end{bmatrix}$
-$=I_{3}+\lambda\begin{bmatrix}-y_{2}y_{3}&0&0\\ y_{1}y_{3}&0&-y_{3}^{2}\\ 0&0&y_{2}y_{3}\end{bmatrix}+y_{2}y_{3}^{2}D^{\prime},$
-where
-$D^{\prime}=\begin{bmatrix}\lambda^{2}y_{2}&0&-\lambda^{3}y_{2}\\ \lambda(y_{3}-\lambda y_{1}-\lambda y_{2}y_{3}^{2})&0&\lambda^{3}y_{1}+\lambda^{3}y_{2}y_{3}^{2}\\ -\lambda y_{2}(1-\lambda y_{2}y_{3})&0&-\lambda^{3}y_{2}^{2}y_{3}\end{bmatrix}$.
-Let $E=C(23)D^{(23)}(23)\in S$. A direct calculation gives that
-$E_{3}=C_{3}(23)_{3}D_{3}^{(23)}(23)_{3}=(I_{3}+\lambda\begin{bmatrix}y_{2}y_{3}&0&0\\ 0&-y_{2}y_{3}&0\\ -y_{1}y_{2}&y_{2}^{2}&0\end{bmatrix}+y_{2}^{2}y_{3}C^{\prime})$
-$\times(I_{3}+\lambda\begin{bmatrix}-y_{2}y_{3}&0&0\\ y_{1}y_{3}&0&-y_{3}^{2}\\ 0&0&y_{2}y_{3}\end{bmatrix}+y_{2}y_{3}^{2}D^{\prime})=I_{3}+\lambda\begin{bmatrix}0&0&0\\ y_{1}y_{3}&-y_{2}y_{3}&-y_{3}^{2}\\ -y_{1}y_{2}&y_{2}^{2}&y_{2}y_{3}\end{bmatrix}$
-$+\lambda^{2}y_{2}y_{3}\begin{bmatrix}-y_{2}y_{3}&0&0\\ -y_{1}y_{3}&0&y_{3}^{2}\\ 2y_{1}y_{2}&0&-y_{2}y_{3}\end{bmatrix}+(I_{3}+\lambda\begin{bmatrix}y_{2}y_{3}&0&0\\ 0&-y_{2}y_{3}&0\\ -y_{1}y_{2}&y_{2}^{2}&0\end{bmatrix}+y_{2}^{2}y_{3}C^{\prime})y_{2}y_{3}^{2}D^{\prime}$
-$+y_{2}^{2}y_{3}C^{\prime}+\lambda y_{2}^{2}y_{3}C^{\prime}\begin{bmatrix}-y_{2}y_{3}&0&0\\ y_{1}y_{3}&0&-y_{3}^{2}\\ 0&0&y_{2}y_{3}\end{bmatrix}=I_{3}+\lambda\begin{bmatrix}0&0&0\\ y_{1}y_{3}&-y_{2}y_{3}&-y_{3}^{2}\\ -y_{1}y_{2}&y_{2}^{2}&y_{2}y_{3}\end{bmatrix}+E^{\prime},$
-where all elements of $E^{\prime}$ are divisible by $y_{3}$ and all elements of the last column of $E^{\prime}$ are divisible by $y_{3}^{2}$. Let
-$P=(I+\lambda y_{1}F_{231})(I-\lambda y_{2}F_{232}-\lambda y_{3}F_{233})=I+\lambda y_{1}F_{231}-\lambda y_{2}F_{232}-\lambda y_{3}F_{233}.$
-Then
-$P_{3}=I_{3}+\lambda\begin{bmatrix}0&0&0\\ y_{1}y_{3}&-y_{2}y_{3}&-y_{3}^{2}\\ -y_{1}y_{2}&y_{2}^{2}&y_{2}y_{3}\end{bmatrix}$.
-Then
-$P_{3}^{-1}E_{3}=(I_{3}-\lambda\begin{bmatrix}0&0&0\\ y_{1}y_{3}&-y_{2}y_{3}&-y_{3}^{2}\\ -y_{1}y_{2}&y_{2}^{2}&y_{2}y_{3}\end{bmatrix})(I_{3}+\lambda\begin{bmatrix}0&0&0\\ y_{1}y_{3}&-y_{2}y_{3}&-y_{3}^{2}\\ -y_{1}y_{2}&y_{2}^{2}&y_{2}y_{3}\end{bmatrix}+E^{\prime})$
-$=I_{3}+(I_{3}-\lambda\begin{bmatrix}0&0&0\\ y_{1}y_{3}&-y_{2}y_{3}&-y_{3}^{2}\\ -y_{1}y_{2}&y_{2}^{2}&y_{2}y_{3}\end{bmatrix})E^{\prime}=I_{3}+P^{\prime}.$
-Notice that all elements of $P^{\prime}$ are of lower degree $\geq 4$, divisible by $y_{3}$, and all elements of the last column are divisible by $y_{3}^{2}$.
-Thus the matrix $(3n)(P^{-1}E)^{(3n)}(3n)\in SL_{n}(U,y_{n}U)$ and satisfies the conditions of Lemma 11 for $s=n$.
-By Lemma 11, there exists a matrix $Q$ such that $Q^{-1}(3n)(P^{-1}E)^{(3n)}(3n)\in SL_{n}(U,0)$, that is, $Q^{-1}(3n)(P^{-1}E)^{(3n)}(3n)=I$.
-Consequently, $(3n)(P^{-1}E)^{(3n)}(3n)=Q\in S$ and $P^{-1}E\in S$.
-Since $E\in S$ it follows that $P\in S$.
-Notice that the first factor $I+\lambda y_{1}F_{231}$ of $P$ belongs to $S$.
-Consequently,
-$(I-\lambda y_{1}F_{231})P=I-\lambda y_{2}F_{232}-\lambda y_{3}F_{233}\in S.$
-Conjugating the inverse of this matrix we can get the matrix (12). $\Box$
-Corollary 11.
-The automorphism (3) is almost tame.
-10. Elimination of a variable
-In this section, we prove the main proposition to reduce the number of variables in matrices from $G$ modulo $S$, repeating the proofs from [5, pages 99–100], with some minimal necessary changes.
-We present two lemmas to avoid yet another use of inner automorphisms in [5].
-Lemma 13.
-There exists a $3$-reducible matrix $C$ in $T$ such that $C_{3}$ after setting $y_{2}=y_{3}=0$ becomes equal to
-$I_{3}+\lambda y_{1}E_{31}=I_{3}+\begin{bmatrix}0&0&0\\ 0&0&0\\ \lambda y_{1}&0&0\end{bmatrix}$,
-where $\lambda\in K$.
-Proof. Set
-$A=I+F_{231},\ \alpha=I+\lambda E_{12},\ B=I-\lambda^{2}F_{132}.$
-Then
-$A_{3}\alpha_{3}(A_{3}^{-1})^{\alpha}\alpha_{3}^{-1}=\begin{bmatrix}1&0&0\\ y_{3}&1&0\\ -y_{2}&0&1\end{bmatrix}\begin{bmatrix}1&\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}1&0&0\\ -y_{3}&1&0\\ y_{2}&0&1\end{bmatrix}^{\alpha}\begin{bmatrix}1&-\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$
-$=\begin{bmatrix}1&0&0\\ y_{3}&1&0\\ -y_{2}&0&1\end{bmatrix}\begin{bmatrix}1&\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}1&0&0\\ -y_{3}&1&0\\ y_{2}+\lambda y_{1}&0&1\end{bmatrix}\begin{bmatrix}1&-\lambda&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$
-$=\begin{bmatrix}1-\lambda y_{3}&\lambda^{2}y_{3}&0\\ -\lambda y_{3}^{2}&1+\lambda y_{3}+\lambda^{2}y_{3}^{2}&0\\ \lambda y_{1}+\lambda y_{2}y_{3}&-\lambda y_{2}-\lambda^{2}y_{1}-\lambda^{2}y_{2}y_{3}&1\end{bmatrix}$
-and
-$B_{3}A_{3}\alpha_{3}(A_{3}^{-1})^{\alpha}\alpha_{3}^{-1}=\begin{bmatrix}1&-\lambda^{2}y_{3}&0\\ 0&1&0\\ 0&\lambda^{2}y_{1}&1\end{bmatrix}\begin{bmatrix}1-\lambda y_{3}&\lambda^{2}y_{3}&0\\ -\lambda y_{3}^{2}&1+\lambda y_{3}+\lambda^{2}y_{3}^{2}&0\\ \lambda y_{1}+\lambda y_{2}y_{3}&-\lambda y_{2}-\lambda^{2}y_{1}-\lambda^{2}y_{2}y_{3}&1\end{bmatrix}$
-$=\begin{bmatrix}1-\lambda y_{3}+\lambda^{3}y_{3}^{3}&-\lambda^{3}y_{3}^{2}-\lambda^{4}y_{3}^{3}&0\\ -\lambda y_{3}^{2}&1+\lambda y_{3}+\lambda^{2}y_{3}^{2}&0\\ \lambda y_{1}+\lambda y_{2}y_{3}-\lambda^{3}y_{1}y_{3}^{2}&-\lambda y_{2}-\lambda^{2}y_{2}y_{3}+\lambda^{3}y_{1}y_{3}+\lambda^{4}y_{1}y_{3}^{2}&1\end{bmatrix}$.
-If $C=BA\alpha(A^{-1})^{\alpha}\alpha^{-1}$ then $C\in T$ and $C_{3}=B_{3}A_{3}\alpha_{3}(A_{3}^{-1})^{\alpha}\alpha_{3}^{-1}$ after setting $y_{2}=y_{3}=0$ gives $I_{3}+\lambda y_{1}E_{31}$. $\Box$
-Lemma 14.
-Let
-$I_{n}+[a_{ij}]\in SL_{n}(U,y_{1}U+y_{2}U)\cap G.$
-Then $\mathrm{ldeg}(a_{ij})\geq 2$ for all $1\leq i,j\leq 2$.
-Proof. Let
-$A=[a_{ij}]=A_{1}+\ldots+A_{k},$
-where all elements of $A_{i}$ are homogeneous of degree $i$ for all $1\leq i\leq k$ and let $A_{1}=[b_{ij}]$.
-Since $a_{ij}\in y_{1}U+y_{2}U$ then $b_{ij}\in Ky_{1}+Ky_{2}$.
-Moreover, since $YA=0$ by Lemma 2, we get $YA_{1}=0$.
-By Lemma 2 there exist quadratic elements $f_{1},\ldots,f_{n}\in M_{n}$ such that
-$A_{1}=[b_{ij}]=[\partial(f_{1})\ldots\partial(f_{n})].$
-Therefore all components of $\partial(f_{i})$ belong to $Ky_{1}+Ky_{2}$.
-Let
-$f_{i}=\sum_{k=1}^{n-1}\alpha_{k}[x_{n},x_{k}]+f_{i}^{\prime},$
-where $f_{i}^{\prime}$ does not depend on $x_{n}$ and $\alpha_{1},\ldots,\alpha_{n-1}\in K$. If $\alpha_{j}\neq 0$ for some $1\leq j\leq n-1$ then
-$\frac{\partial f_{i}}{\partial x_{j}}=-\alpha_{j}x_{n}+\frac{\partial f_{i}^{\prime}}{\partial x_{j}}\notin Ky_{1}+Ky_{2}.$
-Consequently, $\alpha_{1}=\ldots=\alpha_{n-1}=0$ and $f_{i}$ does not depend on $x_{n}$.
-The same discussions give that $f_{i}$ does not depend on $x_{3},\ldots,x_{n}$.
-Hence $f_{i}=\lambda_{i}[x_{1},x_{2}]$ for all $i$.
-Then
-$A_{1}=\lambda_{1}F_{121}+\lambda_{2}F_{122}+\ldots+\lambda_{n}F_{12n}$
-and the trace $\mathrm{tr}(A_{1})$ of $A_{1}$ is $\lambda_{1}y_{2}-\lambda y_{1}$.
-Recall that the determinat of $I+A$ is $1$.
-The homogeneous part of degree $1$ of the determinant of $I+A_{1}+\ldots+A_{k}$ is $\mathrm{tr}(A_{1})$.
-Consequently, $\mathrm{tr}(A_{1})=0$.
-The equation $\lambda_{1}y_{2}-\lambda y_{1}=0$ gives that $\lambda_{1}=\lambda_{2}=0$.
-This means that $b_{ij}=0$ and $\mathrm{ldeg}(a_{ij})\geq 2$ for all $1\leq i,j\leq 2$.
-$\Box$
-Lemma 15.
-Let
-$A=I+[a_{ij}]\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)\cap G.$
-Then there exists $R\in S$ such that $R^{-1}A=I+[e_{ij}]\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)\cap G$ and $e_{1n},e_{2n},\ldots,e_{n-1n}\in\sum_{p=s}^{n-1}y_{p}U+y_{n}^{2}U$.
-Proof. For any $a\in U$ and any matrix $R$ over $U$ let $\overline{a}$ and $\overline{R}$ be the result of setting $y_{s}=y_{s+1}=\ldots=y_{n-1}=0$ in $a$ and $R$, respectively.
-By Lemma 2,
-$y_{1}a_{1j}+y_{2}a_{2j}+\ldots y_{n}a_{nj}=0,\ 1\leq j\leq n.$
-Consequently,
-$y_{1}\overline{a}_{1j}+\ldots y_{s-1}\overline{a}_{(s-1)j}+y_{n}\overline{a}_{nj}=0,\ 1\leq j\leq n.$
-Since $A=I+[a_{ij}]\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)$ it follows that $a_{ij}\in\sum_{p=s}^{n}y_{p}U$ and, consequently, $\overline{a}_{ij}\in y_{n}K[y_{1},\ldots,y_{s-1},y_{n}]$, that is, $\overline{a}_{ij}=y_{n}\overline{b}_{ij}$ for some $\overline{b}_{ij}\in K[y_{1},\ldots,y_{s-1},y_{n}]$.
-Therefore $\overline{A}=I+[\overline{a}_{ij}]=I+[y_{n}\overline{b}_{ij}]$ and the preceding equality gives that
-$y_{1}\overline{b}_{1j}+\ldots y_{s-1}\overline{b}_{(s-1)j}+y_{n}\overline{b}_{nj}=0,\ 1\leq j\leq n.$
-If $(\overline{c}_{1n},\ldots,\overline{c}_{(n-1)n})$ is the result of setting $y_{n}=0$ in $(\overline{b}_{1n},\ldots,\overline{b}_{(n-1)n})$ then from this we get
-(16) $y_{1}\overline{c}_{1n}+\ldots y_{s-1}\overline{c}_{(s-1)n}=0.$
-First we show that there exists a matrix $R\in S\cap SL_{n}(U,\sum_{p=s}^{n}y_{p}U)$ such that
-$\overline{R}=C=\begin{bmatrix}I_{n-1}&\begin{matrix}y_{n}\overline{c}_{1n}\\ .\\ .\\ .\\ y_{n}\overline{c}_{(n-1)n}\end{matrix}\\ \begin{matrix}0&\ldots&0\end{matrix}&1\end{bmatrix}$.
-belongs to $S$. In fact,
-$C=\begin{bmatrix}I_{n-1}&\begin{matrix}y_{n}\overline{c}_{1n}\\ .\\ .\\ .\\ y_{n}\overline{c}_{(s-1)n}\\ 0\\ .\\ .\\ .\\ 0\end{matrix}\\ \begin{matrix}0&\ldots&0\end{matrix}&1\end{bmatrix}\begin{bmatrix}I_{n-1}&\begin{matrix}0\\ .\\ .\\ .\\ 0\\ y_{n}\overline{c}_{sn}\\ .\\ .\\ .\\ y_{n}\overline{c}_{(n-1)n}\end{matrix}\\ \begin{matrix}0&\ldots&0\end{matrix}&1\end{bmatrix}$.
-The first factor belongs to $G$ by (16) and Lemma 2, and, consequently, belongs to $S$. The second factor is equal to the product
-$D=\prod_{k=s}^{n-1}(I+y_{n}\overline{c}_{kn}E_{kn})$
-Let
-$\overline{c}_{kn}=y_{1}f_{1}+\ldots+y_{s-1}f_{s-1}+\lambda,$
-where $f_{i}\in K[y_{1},\ldots,y_{s-1}]$ and $\lambda\in K$. Then
-$y_{n}\overline{c}_{kn}=y_{1}y_{n}f_{1}+\ldots+y_{s-1}y_{n}f_{s-1}+y_{n}\lambda.$
-Let
-$P_{k}=\begin{bmatrix}I_{n-1}&\begin{matrix}-y_{k}y_{n}f_{1}\\ .\\ .\\ .\\ -y_{k}y_{n}f_{s-1}\\ 0\\ y_{1}y_{n}f_{1}+\ldots+y_{s-1}y_{n}f_{s-1}\\ 0\end{matrix}\\ \begin{matrix}0&\ldots&0\end{matrix}&1\end{bmatrix}\in S.$
-Let s=n1s=n-1.
-Then k=n1k=n-1 and $a_{ij}\in y_{n-1}U+y_{n}U$.
-By Lemma 14, $\mathrm{ldeg}(a_{ij})\geq 2$ for all $n-1\leq i,j\leq n$.
-By construction the linear part of $y_{n}\overline{c}_{kn}$ is obtained from the linear part of $a_{(n-1)n}$.
-Consequently, $\lambda=0$ and $\overline{P}_{k}=I+y_{n}\overline{c}_{kn}E_{kn}$.
-Let s<n1s<n-1, i.e., the set $\{s,\ldots,n\}$ contains at least $3$ elements. Then, by Lemma 13, there exists a matrix $Q\in T$ such that
-$\overline{Q}=I_{n}+\lambda y_{n}E_{kn}.$
-Consequently, $QP\in S$ and $\overline{QP}=I+y_{n}\overline{c}_{kn}E_{kn}$.
-Therefore in both cases there exists a matrix $\gamma_{i}\in S$ such that $\overline{\gamma_{i}}=I+y_{n}\overline{c}_{kn}E_{kn}$.
-Then $\overline{\gamma_{s}\ldots\gamma_{n-1}}=D$.
-Consequently, there exists a matrix $R\in S\cap SL_{n}(U,\sum_{p=s}^{n}y_{p}U)$ such that $\overline{R}=C$.
-Then $R^{-1}A\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)\cap G$.
-Let $R^{-1}A=I_{n}+[e_{ij}]$.
-Recall that, by construction, $\overline{a}_{kn}=y_{n}\overline{c}_{kn}+y_{n}^{2}\overline{d}_{kn}$ for all $1\leq k\leq n-1$, where $\overline{d}_{kn}\in K[y_{1},\ldots,y_{s-1},y_{n}]$.
-Then
-[…]
-$=I+\begin{bmatrix}[\overline{a}_{ij}]_{i,j\leq n-1}&\begin{matrix}\overline{a}_{1n}-y_{n}\overline{c}_{1n}\\ .\\ .\\ .\\ \overline{a}_{(n-1)n}-y_{n}\overline{c}_{(n-1)n}\\ \overline{a}_{nn}\end{matrix}\end{bmatrix}=I+\begin{bmatrix}[\overline{a}_{ij}]_{i,j\leq n-1}&\begin{matrix}y_{n}^{2}\overline{d}_{1n}\\ .\\ .\\ .\\ y_{n}^{2}\overline{d}_{(n-1)n}\\ \overline{a}_{nn}\end{matrix}\end{bmatrix}$
-Then $\overline{e}_{kn}=y_{n}^{2}\overline{d}_{kn}$.
-This means that $e_{kn}\in\sum_{p=s}^{n-1}y_{p}U+y_{n}^{2}U$ for all $1\leq k\leq n-1$.
-The matrix $R^{-1}A$ satisfies the statement of the lemma.
-$\Box$
-For convenience, we formulate the following complete version of Lemma 11.
-Lemma 16.
-Let $A=I+[a_{ij}]\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)\cap G$ and $a_{1n},a_{2n},\ldots,a_{(n-1)n}\in\sum_{p=s}^{n-1}y_{p}U+y_{n}^{2}U$.
-Then there exists $B\in S$ such that $B^{-1}A\in SL_{n}(U,\sum_{p=s}^{n-1}y_{p}U)$.
-Proof. In the proof of Lemma 11, the restriction $\mathrm{ldeg}(a_{ij})\geq 3$ was used only once to apply that the matrix
-$I_{n}+y_{k}\overline{d}_{kn}F_{knk}-y_{n}\overline{d}_{kn}F_{knn}$
-belong to $S$ by Lemma 7 since $\mathrm{ldeg}(\overline{d}_{kn})\geq 1$. By Lemma 12, it belongs to $S$ anyway. Repeating the proof of Lemma 11, we get the statement of the lemma. $\Box$
-Proposition 3.
-Let
-$A=I+[a_{ij}]\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)\cap G.$
-Then there exists $B\in S$ such that $B^{-1}A\in SL_{n}(U,\sum_{p=s}^{n-1}y_{p}U)$.
-Proof.
-By Lemma 15, there exists a matrix $C\in S$ such that $C^{-1}A=I+[c_{ij}]\in SL_{n}(U,\sum_{p=s}^{n}y_{p}U)$ and $c_{1n},c_{2n},\ldots,c_{n-1n}\in\sum_{p=s}^{n-1}y_{p}U+y_{n}^{2}U$.
-By Lemma 16, there exists a matrix $D\in S$ such that $D^{-1}C^{-1}A\in SL_{n}(U,\sum_{p=s}^{n-1}y_{p}U)$.
-$\Box$
-11. Main results
-Theorem 1.
-Every automorphism of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ over an arbitrary field $K$ is almost tame.
-Proof. First we show that $S=G$. Let $A\in G$. Then
-$A\in SL_{n}(U,\sum_{j=1}^{n}y_{j}U).$
-By Proposition 3, there exist matrices $B_{1},\ldots,B_{n}\in S$ such that
-$B_{s}^{-1}\ldots B_{1}^{-1}A\in SL_{n}(U,\sum_{j=s+1}^{n}y_{j}U)$
-for all $s<n$ and
-$B_{n}^{-1}\ldots B_{1}^{-1}A\in SL_{n}(U,0)=\{I\}.$
-Consequently,
-$B_{n}^{-1}\ldots B_{1}^{-1}A=I$
-and
-$A=B_{1}\ldots B_{n}\in S.$
-This proves that $S=G$.
-Therefore $\mathrm{ATAut}(M_{n})\cap\mathrm{IAut}(M_{n})=\mathrm{IAut}(M_{n})$.
-It follows that $\mathrm{ATAut}(M_{n})=\mathrm{Aut}(M_{n})$ since both groups contain all linear automorphisms.
-$\Box$
-Theorem 2.
-The group of all automorphisms $\mathrm{Aut}(M_{n})$ of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ over a field $K$ of characteristic $\neq 3$ is generated by all linear automorphisms, as well as a quadratic automorphism (1) and a cubic automorphism (2).
-Corollary 12.
-The group of all automorphisms $\mathrm{Aut}(M_{n})$ of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 5$ over any field $K$ is generated by all linear automorphisms, as well as a quadratic automorphism (1) and a cubic automorphism (2).
-Theorem 2 and Corollary 12 are direct corollaries of Corollary 4 and Theorem 1.
+$\tau=(x_{1},x_{2},x_{3}+[x_{1},x_{n}]hf+[x_{2},x_{n}]hg,x_{4},\ldots,x_{n}).$
+Note that $\tau$ is a Chein automorphism. Hence $B(h,f,g)$ is almost tame. $\Box$
 Acknowledgments